### v1

# On Modular Invariants of Truncated Polynomial Rings in low ranks

Abstract.
We verify the conjectures due to Lewis, Reiner and Stanton about the Hilbert series of the invariant ring of the truncated polynomial ring for all parabolic subgroups up to rank $3$. This is done by constructing an explicit set of generators for each invariant ring in question. We also propose a conjecture concerning the action of the Steenrod algebra and the Dickson algebra on a certain naturally occurring filtration of the invariant ring under the general linear group.
Key words and phrases:
Truncated Polynomial Ring, Modular invariant, Dickson invariants, Steenrod algebra, q-binomial, q-multinomial, finite general linear group, Frobenius power
1991 Mathematics Subject Classification:
Primary 54C40, 14E20; Secondary 46E25, 20C20
1. Introduction
Let $\mathbb{F}_{q}$ denote the finite field of $q$ elements, where $q$ is a power of a fixed prime number $p$. Let $I$ denote the ideal generated by the $m$th Frobenius power $x_{1}^{q^{m}},x_{2}^{q^{m}},\ldots,x_{n}^{q^{m}}$, of the polynomial algebra $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$. We denote by ${\mathcal{Q}}_{m}(n)$ or simply ${\mathcal{Q}}_{m}$, if $n$ is fixed, the quotient ring $S/I$. The general linear group $G=\operatorname{GL}_{n}(\mathbb{F}_{q})$ acts on $S$ by linear substitutions of variables. Since $I$ is a $G$-invariant ideal, ${\mathcal{Q}}_{m}(n)$ inherits a $G$-action from $S$. This paper discusses the problem of computing invariants of the truncated ring ${\mathcal{Q}}_{m}(n)$ under the action of various parabolic subgroups of $G$.
This problem arises in various contexts, in connection with questions in algebraic combinatorics ([14], [15], [10], [4]), algebraic topology ([1], [7], [9]) and number theory [2]. In particular, taking the limit $m$ to infinity yields results about the cofixed space $S_{G}$ of the polynomial ring under the action of the general linear group. This problem is surprisingly still open in modular invariant theory.
Let $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ be a composition of $n$ and denote by $P(\alpha)$ the corresponding parabolic subgroup. In [10], Lewis, Reiner, and Stanton proposed a marvelous series of conjectures about the Hilbert-Poincaré series of the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ for arbitrary $m$ and any composition $\alpha$ of $n$ in terms of a new combinatorial object called the $(q,t)$-binomial coefficient introduced in [15]:
Conjecture 1.1.
([10, Conjecture 1.5]) The Hilbert-Poincaré series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the parabolic subgroup $P(\alpha)$ is:
$C_{\alpha,m}(t)=\sum_{\beta,\beta\leq\alpha,|\beta|\leq m}t^{e(m,\alpha,\beta)}\;\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t},$
where $e(m,\alpha,\beta)=\sum(\alpha_{i}-\beta_{i})(q^{m}-q^{B_{i}})$, $B_{i}=\beta_{1}+\ldots+\beta_{i}$, and for a composition $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ of $d$ with partial sum $\alpha_{1}+\ldots+\alpha_{i}=A_{i}$, the $(q,t)$ binomial coefficient $\begin{bmatrix}d\\ \alpha\end{bmatrix}_{q,t}$ is defined as the quotient
$\begin{bmatrix}d\\ \alpha\end{bmatrix}_{q,t}=\frac{\prod_{j=0}^{d-1}(1-t^{q^{d}-q^{j}})}{\prod_{i=1}^{\ell}\prod_{j=0}^{\alpha_{i}}(1-t^{q^{A_{i}}-q^{A_{i-1}+j}})}.$
In particular, when $\alpha=(n)$, they conjectured that
Conjecture 1.2.
([10, Conjecture 1.2]) The Hilbert-Poincaré series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the general linear group $G=\operatorname{GL}_{n}(\mathbb{F}_{q})$ is:
$C_{n,m}(t)=\sum_{k=0}^{\min(n,m)}t^{(n-k)(q^{m}-q^{k})}\;\begin{bmatrix}m\\ k\end{bmatrix}_{q,t}.$
These conjectures are inspired by properties of the $q$-Catalan and $q$-Fuss Catalan numbers, connecting Hilbert-Poincaré series for certain invariant subspaces with the representation theory of rational Cherednik algebras for Coxeter and complex reflection groups. Some partial results about these two conjectures have been obtained (see Drescher-Shepler [4], Goyal [5] and the original paper [10]). Still, how these combinatorial objects occur in the series remains quite mysterious.
In our previous work [6], we have proved their parabolic conjecture for the minimal parabolic subgroup by constructing an explicit basis. We also refined their conjectures by proposing an explicit $\mathbb{F}_{q}$-basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ as follows:
Conjecture 1.3.
A basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is given inductively by the following formula:
$\mathcal{B}_{m}(\alpha)=\{\delta_{s+1;m}^{\alpha_{1}-s}(f\cdot\Phi^{s}g)\colon 0\leq s\leq\min(\alpha_{1},m),f\in\Delta^{m}_{s},g\in\mathcal{B}_{m-s}(\alpha_{2},\ldots,\alpha_{\ell})\}.$
Here $\Delta^{m}_{s}$ is a certain subspace of the Dickson algebra ([3]) of all $\operatorname{GL}_{s}$-invariant polynomials and $\delta_{a;b}$ is the operator constructed in [6]. The Frobenius-like operator $\Phi$ is given by the rules:
  1. (i)
    $\Phi(\delta_{a;b})=\delta_{a+1;b+1}$;
  2. (ii)
    $\Phi(Q_{r,i})=Q_{r+1,i+1}$ and expand multiplicatively.
The reader of [6] may notice that in this conjecture, we have used the core subspace $\Delta^{m}_{s}$ of the Dickson algebra instead of $\nabla^{m}_{s}$ in the original conjecture. The two candidates can be used interchangeably, but $\Delta^{m}_{s}$ is more convenient for explicit calculation, as we will see later.
In the case $\alpha=(n)$ so that $P(\alpha)=G$ is the full general linear group, we have
Conjecture 1.4.
The set $\mathcal{B}_{m}(n)$ consisting of the following family of elements:
$\delta_{s+1;m}^{n-s}(f),f\in\Delta^{m}_{s},0\leq s\leq\min(m,n),$
forms a basis for the $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}^{G}(n)$.
It is a simple counting problem that our proposed bases have the correct series as predicted. Hence our conjectures refine the two conjectures proposed by Lewis, Reiner, and Stanton. Furthermore, they also explain where the summands of the Hilbert-Poincaré series occur in the original conjectures: The summands of $C_{n,m}(t)$ are just the Hilbert-Poincaré series for $\Delta^{m}_{s}$, $0\leq s\leq\min(m,n)$, shifted by an appropriate degree because of the iterated delta operator.
In [6], we have verified the refined parabolic conjecture 1.3 for the Borel subgroup, that is when $\alpha=1^{n}$. The aim of this note is to verify our conjectures for other parabolic subgroups, including the full general linear group, for $n\leq 3$. As an immediate corollary, the parabolic conjecture of Lewis, Reiner and Stanton is true in these cases.
Theorem 1.5.
The refined parabolic conjectures 1.3 and 1.4 are true for $n\leq 3$ and all $m\geq 1$. Hence, the parabolic conjecture of Lewis, Reiner, and Stanton is true in these cases.
In this paper, we choose to work as explicitly as possible to demonstrate the underlying mechanism of our construction and the additional structure of our basis. The computation reveals an interesting property of the invariant rings considered as a module over the Dickson algebra. We also describe the action of the Steenrod algebra. In subsequent work, we hope to return to this problem in greater generality.
When $n=2$, the only parabolic subgroup other than the Borel subgroup is $\operatorname{GL}_{2}$ itself, and the parabolic conjecture has already been verified in the original paper [10]. However, the authors worked with the dual situation for cofixed spaces and used a particular property of the bivariate case to avoid the truncated ring. Thus, it is unclear whether the invariant subspace ${\mathcal{Q}}_{m}^{G}$, even in rank $2$, can be described explicitly using their method.
In [5, Proposition 4.1], Goyal constructed a family of $G$-invariant polynomials in ${\mathcal{Q}}_{m}(2)$, which he denoted as $y_{k^{\prime}}$ ($0\leq k^{\prime}\leq\frac{q^{m}-q}{q-1}$). It turns out that this family coincides with the $G$-invariant family $\delta_{2}(\Delta^{m}_{1})$ in our notation.
To simplify the notation, we will often drop the index $n$ in ${\mathcal{Q}}_{m}(n)$ when the number of variables is well-understood from the context. Moreover, since the Frobenius power $m$ is fixed, $\delta_{s}$ will be always used in place of $\delta_{s;m}$ throughout the paper. We will also write $(i_{1},i_{2})\in\Delta^{m}_{2}$ instead of $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}\in\Delta^{m}_{2}$. Finally we denote by $[a]_{q}$ the $q$-integer $(q^{a}-1)/(q-1)$.
Here is our first result in rank $2$.
Proposition 1.6.
The set $\mathcal{B}_{m}(2)$ consisting of 3 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}^{G}(2)$ of the full general linear group.
  1. (1)
    $\delta_{1}(\delta_{1}(1))=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$.
  2. (2)
    $\delta_{2}(\Delta^{m}_{1})$.
  3. (3)
    $\Delta^{m}_{2}$.
In rank 3, other than the Borel subgroups, there are 2 proper parabolic subgroups. Their corresponding rings of invariants are described in the next 2 propositions.
Proposition 1.7.
The set $\mathcal{B}_{m}(2,1)$ consisting of 6 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}^{P(2,1)}$ of the parabolic subgroup corresponding to the composition $(2,1)$.
  1. (1)
    $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}Q_{3,2}^{i}$, $(i_{1},i_{2})\in\Delta^{m}_{2}$, $i<[m-2]_{q}$.
  2. (2)
    $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}\delta_{3}(1)$, $(i_{1},i_{2})\in\Delta^{m}_{2}$.
  3. (3)
    $\delta_{2}(Q_{1,0}^{i_{1}}Q_{2,1}^{i})$, $i_{1}<[m]_{q}$, $i<[m-1]_{q}$.
  4. (4)
    $\delta_{2}(Q_{1,0}^{i_{1}}\delta_{2}(1))$, $i_{1}<[m]_{q}$.
  5. (5)
    $\delta_{1}(\delta_{1}(Q_{1,0}^{i}))$, $i<[m]_{q}$.
  6. (6)
    $\delta_{1}(\delta_{1}(\delta_{1}(1)))$.
Proposition 1.8.
The set $\mathcal{B}_{m}(1,2)$ consisting of 6 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}^{P(1,2)}$ of the parabolic subgroup corresponding to the composition $(1,2)$.
  1. (1)
    $Q_{1,0}^{j_{1}}Q_{3,2}^{i_{1}}Q_{3,1}^{i_{2}}$, $j_{1}<[m]_{q},(i_{1},i_{2})\in\Delta^{m-1}_{2}$.
  2. (2)
    $Q_{1,0}^{j_{1}}\delta_{3}(Q_{2,1}^{j_{2}})$, $j_{1}<[m]_{q},j_{2}<[m-1]_{q}$.
  3. (3)
    $Q_{1,0}^{j_{1}}\delta_{2}(\delta_{2}(1))$, $j_{1}<[m]_{q}$.
  4. (4)
    $\delta_{1}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$, $(i_{1},i_{2})\in\Delta^{m}_{2}$.
  5. (5)
    $\delta_{1}(\delta_{2}(Q_{1,0}^{j_{1}}))$, $j_{1}<[m]_{q}$.
  6. (6)
    $\delta_{1}(\delta_{1}(\delta_{1}(1)))$.
Our final result, which is also the most technical, is for the full general linear group in rank 3:
Proposition 1.9.
The set $\mathcal{B}_{m}(3)$ consisting of 4 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}(3)^{G}$ of the full general linear group.
  1. (1)
    $\delta_{1}(\delta_{1}(\delta_{1}(1)))=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$.
  2. (2)
    $\delta_{2}(\delta_{2}(\Delta^{m}_{1}))$.
  3. (3)
    $\delta_{3}(\Delta^{m}_{2})$.
  4. (4)
    $\Delta^{m}_{3}$.
Since the case of the Borel subgroup, which corresponds to the composition $(1,1,1)$ of $n=3$, has been discussed in full generality in [6], the Propositions above combined imply the main Theorem 1.5.
Note that when $m$ is small compared to $n$, all results remain valid with proper interpretation. For example if $n=3$ and $m=2$, then $\Delta^{2}_{3}$ is empty, and $\delta_{3;2}(\Delta^{2}_{2})$ just consists of a single element, which is $1$ in degree 00. To avoid too many special cases (whose proof are not complicated), we will generally assume that $m$ is sufficiently large, comparing to $n$.
In the course of proving these results, we have also computed the structure of these invariant rings, considered as a module over the Dickson algebra, ${\mathcal{D}}_{n}$, and a module over the Steenrod algebra $\mathcal{A}$. At least in the case we have computed, our basis yields a remarkable filtration, namely the subspaces $\mathcal{F}_{n,k}$, for $0\leq k\leq n$, of ${\mathcal{Q}}_{m}^{G}(n)$ spanned by $\delta_{s+1}^{n-s}(\Delta^{m}_{s})$ for all $s\leq k$. We have the following conjecture about this filtration:
Conjecture 1.10.
For each $1\leq k\leq\min(m,n)$, the subspace $\mathcal{F}_{n,k}$ given by
$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq k\},$
is an $\mathcal{A}$-submodule as well as a ${\mathcal{D}}_{n}$-submodule of ${\mathcal{Q}}_{m}^{G}(n)$. Moreover, $\mathcal{F}_{n,k}$ is annihilated by the Dickson invariants $Q_{n,0},Q_{n,1}\ldots Q_{n,n-k-1}$.
This conjecture suggests that the filtration $\mathcal{F}_{n,*}$ may have another description of topological nature. We can prove the following:
Theorem 1.11.
The conjecture 1.10 is true for $n\leq 3$.
The layout of this paper is as follows: Section 2 recalls basic results about modular invariant theory and the key construction from our previous work.
In Section 3, we give a lower bound on the total dimension of the invariant subspaces. The following four Sections are devoted to demonstrating the four Propositions from 1.6 to 1.9 above. The last section discusses the action of the Dickson algebra and the Steenrod algebra and the proof of Theorem 1.11.
Acknowledgement
This work was partially completed while the second and third authors visited the Vietnam Institute for Advanced Study in Mathematics (VIASM). We want to thank Vo Thanh Tung for some computer calculations in the initial stage of this project.
2. Recollection
In this section, we recall some background materials about modular invariant theory and the main construction in [6].
2.1. Upper triangular invariants and Dickson invariants
The canonical projection $S\to{\mathcal{Q}}_{m}$ is clearly a $\operatorname{GL}$-map, hence for each composition $\alpha$ of $n$, there is an induced $\mathbb{F}_{q}$-algebra homomorphism
$S^{G}\subset S^{P(\alpha)}\to{\mathcal{Q}}_{m}^{P(\alpha)},$
making the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}$ into a module over $S^{G}$. Thanks to the work of Dickson [3] since 1911, the structure of the algebra $S^{G}$, which we now call the Dickson algebra, is well-known, and plays a significant role in our investigation.
For each positive integer $k$, denote by $V_{k}$ the product
$V_{k}(x_{1},\ldots,x_{k})=\prod_{\lambda_{i}\in\mathbb{F}_{q}}(x_{k}+\lambda_{1}x_{1}+\ldots+\lambda_{k-1}x_{k-1}).$
It is well-known that the invariant ring $S^{B}$ under the Borel subgroup of upper triangular matrices is a polynomial algebra generated by $V_{i}^{q-1}$, $1\leq i\leq n$. Let $X$ be an indeterminate and recall the fundamental equation
$V_{n+1}(x_{1},\ldots,x_{n},X)=X^{q^{n}}+\sum_{i=0}^{n-1}(-1)^{n-i}Q_{n,i}X^{q^{i}}.$
The polynomials $Q_{n,i}=Q_{n,i}(x_{1},\ldots,x_{n})$ are evidently $\operatorname{GL}_{n}$-invariant. They are called the Dickson invariants, because of the following celebrated theorem due to Dickson:
${\mathcal{D}}_{n}=S^{G}=\mathbb{F}_{q}[Q_{n,0},\ldots,Q_{n,n-1}].$
The Dickson invariants $Q_{n,i}$ can be described in a more explicit form as follows. For any $n$-tuple of non-negative integers $(r_{1},\dots,r_{n})$, put
$[r_{1},\dots,r_{n}]:=\det(x_{i}^{q^{r_{j}}})_{1\leq i,j\leq n}$
and let $L_{n}=L_{n}(x_{1},\ldots,x_{n})$ denote $[0,1\ldots,n-1]$. We then have for $1\leq i\leq n$
$Q_{n,i}=[0,\ldots,\hat{i},\ldots,n]/L_{n},$
In particular, $Q_{n,n}=1$ and $Q_{n,0}=L_{n}^{q-1}$. By convention, we set $Q_{n,i}=0$ if $i<0$ or $i>n$. The following relations are well-known (see for example [13]):
$L_{n}=V_{1}\cdots V_{n},\quad Q_{n,i}=Q_{n-1,i-1}^{q}+Q_{n-1,i}V_{n}^{q-1}.$
Note that $L_{n}$ is, up to sign, the product of all lines in the $\mathbb{F}_{q}$-vector space spanned by $x_{1},\ldots,x_{n}$.
2.2. The $\Delta$ subspace of the Dickson algebra
Given a positive integer $s$ and a partition $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{s}$ with at most $s$ nonzero parts, we say that a Dickson monomial $Q_{s,s-1}^{e_{1}}\ldots Q_{s,0}^{e_{s}}$ in ${\mathcal{D}}_{s}$ is of type $(\lambda_{1},\ldots,\lambda_{s})$ (or q-compatible with $(\lambda_{1},\ldots,\lambda_{s})$, in the language of [15, Definition 5.4]) if for each $1\leq i\leq s$,
$e_{i}\in\bigg{[}\frac{q^{\lambda_{i}}-q^{\lambda_{i+1}}}{q-1},\frac{q^{\lambda_{i}+1}-q^{\lambda_{i+1}}}{q-1}\bigg{)}.$
It is clear that each Dickson monomial is $q$-compatible with a unique partition $(\lambda_{1},\ldots,\lambda_{s})$ of length $\leq s$.
Now if $s\leq m$ then we restrict our attention to partitions which fit inside a rectangular $s\times(m-s)$, that is partitions with at most $s$ nonzero parts and $m-s\geq\lambda_{1}$. Define $\Delta_{(\lambda_{1},\ldots,\lambda_{s})}$ as the subset of the Dickson algebra ${\mathcal{D}}_{s}$ consisting of Dickson monomials of type $(\lambda_{1},\ldots,\lambda_{s})$, and let $\Delta^{m}_{s}$ denote the disjoint union of all subsets over all such partitions fitting inside $s\times(m-s)$. By convention, we put $\Delta^{m}_{s}=\emptyset$ if $s>m$, $\Delta^{m}_{0}=\{1\}$.
The Dickson monomials in $\Delta^{m}_{s}$ for $s\leq\min(m,n)$ together with the delta operators recalled below are the building blocks for our basis. For this reason, we will say that a Dickson monomial in $\Delta^{m}_{s}$ an essential monomial. We will also work with the edge monomials, which is a Dickson monomial not in $\Delta^{m}_{s}$ but has the form $Q_{s,i}f$ where $f\in\Delta^{m}_{s}$, for some $10\leq i\leq s-$ 1. Clearly, the edge monomials are those such that $e_{i}\in[\frac{q^{\lambda_{i}}-q^{\lambda_{i+1}}}{q-1},\frac{q^{\lambda_{i}+1}-q^{\lambda_{i+1}}}{q-1})$ for all $i\neq j$ and $e_{j}=\frac{q^{\lambda_{j}+1}-q^{\lambda_{j+1}}}{q-1}$ for some $1\leq j\leq s$.
It is easy to check that $\Phi\Delta^{m}_{s}\subset\Delta^{m}_{s+1}$. In fact, $\Phi$ sends a Dickson monomial in ${\mathcal{D}}_{s}$ of type $(\lambda_{1},\ldots,\lambda_{s})$ to a Dickson monomial in ${\mathcal{D}}_{s+1}$ of type $(\lambda_{1},\ldots,\lambda_{s},0)$.
2.3. The delta operator
The operator $\delta_{a;b}$ introduced in [6] is fundamental to our construction. We now recall its definition and main properties.
Definition 2.1.
Let $a,b,c$ be positive integers such that $11\leq a\leq c+$ 1. If $f$ is a rational function in $c$ variables, then $\delta_{a;b}(f)$ is a function in $(c+1)$ variables, defined as the quotient
$\delta_{a;b}(f)=\frac{\begin{vmatrix}x_{1}&\cdots&x_{a}\\ x_{1}^{q}&\cdots&x_{a}^{q}\\ \vdots&\ddots&\vdots\\ x_{1}^{q^{a-2}}&\cdots&x_{a}^{q^{a-2}}\\ x_{1}^{q^{b}}f(\widehat{x_{1}},x_{2},\ldots,x_{c+1})&\cdots&x_{a}^{q^{b}}f(x_{1},\ldots,\widehat{x_{a}},\ldots,x_{c+1})\end{vmatrix}}{\begin{vmatrix}x_{1}&\cdots&x_{a}\\ x_{1}^{q}&\cdots&x_{a}^{q}\\ \vdots&\ddots&\vdots\\ x_{1}^{q^{a-1}}&\cdots&x_{a}^{q^{a-1}}\end{vmatrix}}.$
Thus if $b\geq a-1$, $\delta_{a;b}$ increases degree by $q^{b}-q^{a-1}$. By definition, if we separate the first $(a-1)$ variables and write $f$ in the form $f=\sum g(x_{1},\ldots,x_{a-1})h(x_{a},\ldots,x_{c})$, then
$\delta_{a;b}(f)=\sum\delta_{a;b}(g)h(x_{a+1},\ldots,x_{c+1}).$
In this paper, with the number $m$ fixed from the beginning, we will always write $\delta_{s}$ for $\delta_{s;m}$ and occasionally use the full notation $\delta_{a;b}$ when necessary.
It was shown in greater generality in [6] that $\delta_{s}(f)$ is generally not a polynomial, but it is very often a polynomial and is a $\operatorname{GL}_{s}$-invariant. We give a precise statement and provide an elementary proof in the case for $\delta_{3}$ to make the paper self-contained.
Proposition 2.2.
If $f$ is a $\operatorname{GL}_{2}$-invariant polynomial, then $\delta_{3}(f)$ is a polynomial and is $\operatorname{GL}_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$.
Proof.
We first show that $\delta_{3}(f)$ is a polynomial. The denominator $L_{3}$ is a product of linear forms, so it suffices to prove that the numerator becomes zero whenever there exists a nontrivial linear relation $a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}=0$. This is reduced to the proof that
$a_{1}x_{1}^{q^{m}}f(x_{2},x_{3})+a_{2}x_{2}^{q^{m}}f(x_{1},x_{3})+a_{3}x_{3}^{q^{m}}f(x_{1},x_{2})=0,$
provided $a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}=0$ and not all scalars $a_{i}$ are zero. Without loss of generality, assume that $a_{3}=-1$. If both $a_{1},a_{2}\neq 0$, then since $f$ is a $\operatorname{GL}_{2}$-invariant, we have
$f(x_{2},a_{1}x_{1}+a_{2}x_{2})=f(x_{1},a_{1}x_{1}+a_{2}x_{2})=f(x_{1},x_{2}),$
and the above expression becomes
$[a_{1}x_{1}^{q^{m}}+a_{2}x_{2}^{q^{m}}+a_{3}x_{3}^{q^{m}}]f(x_{1},x_{2})=[a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}]^{q^{m}}f(x_{1},x_{2})=0.$
If one of the $a_{i}$ is zero, say $a_{1}=0$, then $f(x_{2},x_{3})=f(x_{2},a_{3}x_{2})=0$ and $f(x_{1},x_{3})=f(x_{1},x_{2})$, so that the sum is reduced to
$a_{2}x_{2}^{q^{m}}f(x_{1},x_{3})+a_{3}x_{3}^{q^{m}}f(x_{1},x_{2})=(a_{2}x_{2}+a_{3}x_{3})^{q^{m}}f(x_{1},x_{2})=0.$
If $a_{1}=a_{2}=0$ or $x_{3}=0$, this is also clear. We have proved that $\delta_{3}(f)$ is a polynomial.
Next we show that $\delta_{3}(f)$ is $\operatorname{GL}_{3}$-invariant modulo $I_{m}$. Clearly, it is symmetric, so it suffices to prove that $\delta_{3}(f)$ does not change under the operation that sends $x_{1}\mapsto x_{1}+x_{2}$ and fixes $x_{2},x_{3}$. The denominator $L_{3}$ certainly satisfies this condition. By direct inspection, $\delta_{3}(f)(x_{1}+x_{2},x_{2},x_{3})-\delta_{3}(f)(x_{1},x_{2},x_{3})$ equals
$x_{2}^{q^{m}}\times\frac{\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{q}&x_{2}^{q}&x_{3}^{q}\\ f(x_{2},x_{3})-f(x_{1}+x_{2},x_{3})&f(x_{1}+x_{2},x_{3})-f(x_{1},x_{3})&0\end{vmatrix}}{L_{3}}.$
For this expression to be zero in ${\mathcal{Q}}_{m}(3)$, we must prove that the second factor is a polynomial. The rest of the argument is similar to the first part of the proof. ∎
Example 2.3.
The $\operatorname{GL}_{1}$-polynomials are of the form $f(x)=x^{s(q-1)}=Q_{1,0}^{s}$, $s\geq 0$. Then $\delta_{2}(x^{s(q-1)})$ is a $\operatorname{GL}_{2}$-invariant polynomial in ${\mathcal{Q}}_{m}(2)$. For $s<[m]_{q}$, this is Goyal’s $y_{s}$ family in [5] where it was defined as
$y_{s}=x_{1}^{q^{m}-q}x_{2}^{s(q-1)}+x_{1}^{q^{m}-q-(q-1)}x_{2}^{(s+1)(q-1)}+\ldots+x_{1}^{s(q-1)}x_{2}^{q^{m}-q}.$
Our construction is also valid for $s\geq[m]_{q}$. However, it is trivial in ${\mathcal{Q}}_{m}$ except when $s=[m]_{q}+1$, in which case
$\delta_{2}(Q_{1,0}^{[m]_{q}+1})=-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}=-\delta_{1}^{2}(1).$
It is easy to check that $y_{s}$ is a genuine $\operatorname{GL}_{2}$-invariant (that is, a Dickson polynomial) whenever $s=\frac{q^{i}-1}{q-1}$ for some non-negative integer $i$.
Example 2.4.
For $s\leq[m]_{q}$, $\delta_{2}^{2}(Q_{1,0}^{s})$ is a $\operatorname{GL}_{3}$-invariant of degree $2(q^{m}-q)+s(q-1)$ in ${\mathcal{Q}}_{m}$. This family of elements coincides with Goyal’s polynomials $a_{m,3,s}$ in [5, Corollary 4.3] where it was written in a different form:
$a_{m,3,s}=\sum_{(i_{1},i_{2},i_{3})}x_{1}^{i_{1}(q-1)}x_{2}^{i_{2}(q-1)}x_{3}^{i_{3}(q-1)},$
where the sum runs over all triple $(i_{1},i_{2},i_{3})$ of non-negative integers such that $i_{s}<[m]_{q}$ and $i_{1}+i_{2}+i_{3}=2[m]_{q}-2+s$. Again, our definition works also for $s>[m]_{q}$:
$\delta_{2}^{2}(Q_{1,0}^{[m]_{q}+2})=\delta_{1}^{3}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$
and is trivial in all other cases.
2.4. The delta operator and the Dickson algebra
For each composition $\alpha$ of $n$, the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is a module over the Dickson algebra ${\mathcal{D}}_{n}$. The next result describes how the delta operator interacts with the Dickson algebra in low rank.
Proposition 2.5.
We have the following identities in ${\mathcal{Q}}_{m}$.
  1. (1)
    $Q_{s,0}\delta_{s}(f)=0$ for all $f$.
  2. (2)
    $Q_{2,1}\delta_{2}(f)=\delta_{2}(Q_{1,0}^{q}f)$ for all $f\in{\mathcal{D}}_{1}$.
  3. (3)
    $Q_{3,i}\delta_{3}(f)=\delta_{3}(Q_{2,i-1}^{q}f)$ for $i=1$ or $2$ and for all $f\in{\mathcal{D}}_{2}$.
  4. (4)
    $Q_{3,2}\delta_{2}^{2}(f)=\delta_{2}^{2}(Q_{1,0}^{q^{2}}f)$ for all $f\in{\mathcal{D}}_{1}$.
  5. (5)
    $Q_{3,1}\delta_{2}^{2}(f)=0$ for all $f\in{\mathcal{D}}_{1}$.
Proof.
We will prove only statements (3) for $i=1$ and the last two parts, as the proof of other statements use similar arguments and are even more straightforward.
Proof of (3). Consider the difference $Q_{3,1}\delta_{3}(f)-\delta_{3}(Q_{2,0}^{q}f)$. Using the identity $Q_{3,1}=V_{3}^{q-1}Q_{2,1}+Q_{2,0}^{q}$, the (3,1)-entry in the last row of the determinant in the numerator can be simplified as
$x_{1}^{q^{m}}V_{3}^{q-1}(x_{2},x_{3},x_{1})Q_{2,1}(x_{2},x_{3}).$
Since $V_{3}(x_{2},x_{3},x_{1})L_{2}(x_{2},x_{3})=L_{3}(x_{2},x_{3},x_{1})=L_{3}(x_{1},x_{2},x_{3})$, the Laplace expansion along the last row of the determinant in the numerator shows that the resulting polynomial belongs to $I_{m}$, which is zero in ${\mathcal{Q}}_{m}$.
Proof of (4). Write $Q_{3,2}=V_{3}^{q-1}+Q_{2,1}^{q}$. From (1) and (2) we have in ${\mathcal{Q}}_{m}$
$Q_{3,2}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,1}^{q}\delta_{2}f).$
Using the formula $Q_{2,1}=V_{2}^{(q-1)}+Q_{1,0}^{q}$ and unravelling the definition, we see that the difference $\delta_{2}(Q_{2,1}^{q}\delta_{2}f)-\delta_{2}^{2}(Q_{1,0}^{q^{2}}f)$ can be written as the sum
$x_{3}^{q^{m}}\frac{x_{1}^{q^{m}}x_{2}f(x_{2})V_{2}^{q^{2}-q-1}(x_{2},x_{3})-x_{2}^{q^{m}}x_{1}f(x_{1})V_{2}^{q^{2}-q-1}(x_{1},x_{3})}{L_{2}(x_{1},x_{2})}+\\ x_{1}^{q^{m}}x_{2}^{q^{m}}f(x_{3})\frac{x_{2}V_{2}^{q^{2}-q-1}(x_{3},x_{2})-x_{1}V_{2}^{q^{2}-q-1}(x_{3},x_{1})}{L_{2}(x_{1},x_{2})}.$
It is clear that the first summand is a polynomial, and is a multiple of $x_{3}^{q^{m}}$. The second summand belongs to the ideal $(x_{1}^{q^{m}},x_{2}^{q^{m}})$.
Proof of (5). We have in ${\mathcal{Q}}_{m}$
$Q_{3,1}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,0}^{q}\delta_{2}f)$
and
$Q_{2,0}^{q}\delta_{2}(f)=L_{2}^{q^{2}-q-1}(x_{1},x_{2})x_{1}f(x_{1})x_{2}^{q^{m}}-L_{2}^{q^{2}-q-1}(x_{1},x_{2})x_{2}f(x_{2})x_{1}^{q^{m}}.$
The rest of the argument is similar to (4) above. ∎
2.5. The Borel subgroup
In [6], we have constructed a basis for the invariant ring of ${\mathcal{Q}}_{m}^{B}$ under the action of the Borel subgroup for arbirary $m$ and $n$. In the case $n=3$, the $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{B}$ consists 4 families:
  1. (1)
    $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{j_{3}(q-1)}$, $j_{3}\leq[m]_{q}$.
  2. (2)
    $x_{1}^{q^{m}-1}x_{2}^{j_{2}(q-1)}Q_{2,1}(x_{2},x_{3})^{j_{3}}$, $j_{2}<[m]_{q},j_{3}\leq[m-1]_{q}$.
  3. (3)
    $x_{1}^{j_{1}(q-1)}\delta_{2}(Q_{2,1}^{j_{2}})$, $j_{1}<[m]_{q}$, $j_{2}\leq[m-1]_{q}$.
  4. (4)
    $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$, $j_{1}<[m]_{q}$, $j_{2}<[m-1]_{q}$, $j_{3}\leq[m-2]_{q}$.
3. A lower bound on the total dimension of the invariant subspaces
The original parabolic conjecture predicted that the Hilbert-Poincaré series for the space of invariants ${\mathcal{Q}}_{m}^{\alpha}(n)$, where $\alpha$ is a composition of $n$, is the (finite) polynomial $C_{\alpha,m}(t)$, whose explicit form is given in 1.1. In particular, the value of $C_{\alpha,m}(t)$ when $t=1$ should be the total dimension of the graded $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}^{\alpha}(n)$.
In this section, we will show that:
Proposition 3.1.
For each $m,n\geq 1$ and any composition $\alpha$ of $n$, the total dimension of the graded vector space ${\mathcal{Q}}_{m}^{\alpha}(n)$ is at least $C_{\alpha,m}(1)$.
This result is already implicit in the discussion in Section 6 of [10], and is valid for any $n$. This relatively simple remark has the following important consequence in practice: suppose we can construct a generating set for the invariant subspace ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ such that, by dimension counting, the Hilbert-Poincare series $C^{\prime}_{m,\alpha}(t)$ for ${\mathcal{Q}}_{m}^{\alpha}(n)$ is not greater than $C_{\alpha,m}(t)$ ($f(t)\leq g(t)$ iff $g-f$ is a polynomial with non-negative coefficients). Then since there is a reverse inequality $C^{\prime}_{\alpha,m}(1)\geq C_{\alpha,m}(1)$ when evaluated at $t=1$, we can conclude that these two series are identical since both are finite degree polynomials with non-negative coefficients. Hence our generating set is actually a basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$.
Proof of Proposition 3.1: Consider the following ungraded quotient of the polynomial algebra
$\mathcal{R}=\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{q^{m}}-x_{1},\ldots,x_{n}^{q^{m}}-x_{n}).$
The action of $\operatorname{GL}_{n}$ on $\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ descends to this quotient. There is a natural filtration:
$\{0\}=F_{-1}\subset F_{0}\subset F_{1}\subset\cdots\subset F_{n(q^{m}-1)}={\mathcal{R}},$
where $F_{i}$ is the image of the set of all polynomials of degree at most $i$ under the canonical projection $\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]\twoheadrightarrow{\mathcal{R}}$. Let $\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}}$ denote the associated graded vector space $\bigoplus_{i\geq 0}F_{i}/F_{i-1}$ and equip $\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}}$ with a ring structure where the multiplication is induced by the product $F_{i}F_{j}\to F_{i+j}$. The canonical $\operatorname{\mathbb{F}}_{q}$-algebra map
$\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]\to{\mathcal{R}}\to\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}},\quad x_{i}\mapsto\bar{x}_{i}\in F_{1}/F_{0},$
then induces an isomorphism of $\operatorname{GL}_{n}$-equivariant $\operatorname{\mathbb{F}}_{q}$-algebras:
${\mathcal{Q}}_{m}(n)\cong\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}}.$
On the other hand, if we apply the functor taking $\operatorname{GL}_{n}$-invariant on each short exact sequence of $\operatorname{GL}_{n}$-modules
$0\to F_{i-1}\to F_{i}\to F_{i}/F_{i-1}\to 0,$
and computing the dimension, we obtain an inequality
$\dim(F_{i})^{\operatorname{GL}_{n}}\leq\dim(F_{i-1})^{\operatorname{GL}_{n}}+\dim(F_{i}/F_{i-1})^{\operatorname{GL}_{n}}$
and so
$\dim{\mathcal{R}}^{\operatorname{GL}_{n}}\leq\dim(\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}})^{\operatorname{GL}_{n}}\cong\dim{\mathcal{Q}}_{m}^{\operatorname{GL}_{n}}(n).$
To compute the $\operatorname{\mathbb{F}}_{q}$-dimension of ${\mathcal{R}}^{\operatorname{GL}_{n}}$, we extend the field $\operatorname{\mathbb{F}}_{q}$ to a field $\operatorname{\mathbb{F}}$ which contains $\operatorname{\mathbb{F}}_{q^{m}}$. Then the evaluation map
$f(x_{1},\ldots,x_{n})\mapsto[(v_{1},\ldots,v_{n})\in\operatorname{\mathbb{F}}_{q^{m}}^{n}\mapsto f(v_{1},\ldots,v_{n})\in\operatorname{\mathbb{F}}],$
induces an isomorphism of $\operatorname{\mathbb{F}}$-algebras
$\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}\xrightarrow{\cong}\operatorname{Map}\big{(}\operatorname{\mathbb{F}}_{q^{m}}^{n},\operatorname{\mathbb{F}}\big{)}.$
This isomorphism is also $\operatorname{GL}_{n}$-equivariant where $\operatorname{GL}_{n}$ acts on the right-hand side via the embedding $\operatorname{GL}_{n}=\operatorname{GL}_{n}(\operatorname{\mathbb{F}}_{q})\subset\operatorname{GL}_{n}(\operatorname{\mathbb{F}}_{q^{m}}).$ It follows that there is an isomorphism
$(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{\operatorname{GL}_{n}}\cong\operatorname{Map}\big{(}\operatorname{\mathbb{F}}_{q^{m}}^{n}/\operatorname{GL}_{n},\operatorname{\mathbb{F}}\big{)}.$
The $\operatorname{\mathbb{F}}$-dimension of $(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{\operatorname{GL}_{n}}$ is thus equal to the cardinality of the orbit set $(\operatorname{\mathbb{F}}_{q^{m}})^{n}/\operatorname{GL}_{n}$. Finally this cardinality is equal to the sum $\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ where $\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ is the $q$-binomial coefficient which counts the number of $s$-dimensional subspaces of the $\operatorname{\mathbb{F}}_{q}$-vector space $\operatorname{\mathbb{F}}_{q^{m}}$.
In summary, we have shown that
$\dim_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{\operatorname{GL}_{n}}=\dim_{F}(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{\operatorname{GL}_{n}})=\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}=C_{m,n}(1)\leq\dim{\mathcal{Q}}_{m}^{\operatorname{GL}_{n}}(n).$
The argument above clearly is not affected if one uses a parabolic subgroup $P_{\alpha}$ instead of $\operatorname{GL}_{n}$. The proposition is proved. ∎
We end this section with the following observation about our proposed basis $\mathcal{B}_{m}(\alpha)$.
Lemma 3.2.
For each composition $\alpha$ of $n$, the Hilbert-Poincaré series for the $\mathbb{F}_{q}$-space spanned by $\mathcal{B}_{m}(\alpha)$ is not greater than $C_{\alpha,m}(t)$.
Proof.
Recall that
$C_{\alpha,m}(t)=\sum_{\beta\leq\alpha,|\beta|\leq m}t^{e(m,\alpha,\beta)}\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t},$
where
$e(m,\alpha,\beta)=\sum(\alpha_{i}-\beta_{i})(q^{m}-q^{B_{i}})=(\alpha_{1}-\beta_{1})(q^{m}-q^{\beta_{1}})+q^{\beta_{1}}\sum_{i\geq 2}(\alpha_{i}-\beta_{i})(q^{m-\beta_{1}}-q^{\beta_{2}+\ldots+\beta_{i}}).$
Also, from [15, Section 7], we have the following identity for the $(q,t)$-multinomial coefficient:
$\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t}=\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\;\Biggl{(}\begin{bmatrix}m-\beta_{1}\\ \beta_{2},\ldots,\beta_{\ell},m-|\beta|\end{bmatrix}_{q,t}\biggl{)}^{q^{\beta_{1}}}.$
Thus if we write $\beta^{\prime}=(\beta_{2},\ldots,\beta_{\ell})$ and similarly for $\alpha^{\prime}$, then $C_{\alpha,m}(t)$ can be written as
$\sum_{\beta_{1}\leq\min(\alpha_{1},m)}t^{(\alpha_{1}-\beta_{1})(q^{m}-q^{\beta_{1}})}\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\biggl{(}\sum_{\beta^{\prime}\leq\alpha^{\prime},|\beta^{\prime}|\leq m-\beta_{1}}t^{e(m-\beta_{1},\beta^{\prime},\alpha^{\prime})}\begin{bmatrix}m-\beta_{1}\\ \beta^{\prime},m-\beta_{1}-|\beta^{\prime}|\end{bmatrix}_{q,t}\biggl{)}^{q^{\beta_{1}}}.$
Comparing this last formula with the recursive definition of $\mathcal{B}_{m}(\alpha)$, noting that $\delta_{s+1;m}^{\alpha_{1}-s}$ raises degree by $(\alpha_{1}-s)(q^{m}-q^{s})$, the Frobenius operator multiply the degree by $q$, and the Hilbert-Poincaré series for $\Delta^{m}_{s}$ is exactly $\begin{bmatrix}m\\ s\end{bmatrix}_{q,t}$, we see that the result follow easily by induction. ∎
The results in this section simplify our tasks considerably. Indeed, in order to show that $\mathcal{B}_{m}(\alpha)$ is a basis for ${\mathcal{Q}}_{m}^{P_{\alpha}}(n)$, it suffices to verify that it is a generating set.
4. Proof of Proposition 1.6
As remarked in the previous section, it suffices to show that $\mathcal{B}_{m}(2)$ is a generating set for ${\mathcal{Q}}_{m}^{G}(2)$. The proof of the Proposition 1.6 is done in two steps. Firstly, we use the relative transfer from the known invariant ring ${\mathcal{Q}}^{B}$ of the Borel subgroup $B$ to show that a bigger set $\mathcal{B}^{\prime}$, consisting of 3 families:
  1. (1)
    $\delta^{2}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$,
  2. (2)
    $\delta_{2}(Q_{1,0}^{s})=y_{s}$, $0\leq s<[m]_{q}$,
  3. (3)
    ${\mathcal{D}}_{2}$,
is a generating set for ${\mathcal{Q}}_{m}^{G}(2)$. Then we show that to generate the invariant ring, it is possible to replace the Dickson algebra ${\mathcal{D}}_{2}$ in the third family by the smaller subspace $\Delta^{m}_{2}$.
Lemma 4.1.
The set $\mathcal{B}^{\prime}$ spans ${\mathcal{Q}}_{m}^{G}(2)$.
Proof.
In [6], we showed that an $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(1,1)$ for the ring of invariants under the Borel subgroup consisting of the following 2 families
  1. (i)
    $x_{1}^{q^{m}-1}x_{2}^{i(q-1)}$, $i\leq[m]_{q}$,
  2. (ii)
    $Q_{1,0}^{i_{1}}Q_{2,1}^{i_{2}}$, $i_{1}<[m]_{q}$, $i_{2}\leq[m-1]_{q}$.
It is a standard fact about the relative transfer $\operatorname{Tr}\colon{\mathcal{Q}}_{m}^{B}(2)\to{\mathcal{Q}}_{m}^{G}(2)$ that it is onto, so that a generating set for ${\mathcal{Q}}^{G}(2)$ can be obtained from the image of $\operatorname{Tr}$.
By direct inspection, we have the following identity in ${\mathcal{Q}}_{m}$ for any $0\leq s<[m]_{q}$:
$x_{1}^{q^{m}-1}x_{2}^{s(q-1)}=y_{0}x_{1}^{(s+1)(q-1)}-y_{s+1}.$
Since $y_{s}$ is already a $\operatorname{GL}_{2}$-invariant, this formula allows us to compute the transfer of the first family:
$\operatorname{Tr}(x_{1}^{q^{m}-1}x_{2}^{s(q-1)})=y_{0}\operatorname{Tr}(x_{1}^{(s+1)(q-1)})-y_{s+1}.$
Clearly, $\operatorname{Tr}(x_{1}^{(s+1)(q-1)})$ is again a Dickson polynomial in ${\mathcal{D}}_{2}$. By Proposition 2.5, the product of $y_{0}$ with a Dickson polynomial is either zero or another element in the $y$ family. Thus $\operatorname{Tr}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})$ is either $-y_{s+1}$ or zero.
The transfer of the second family $Q_{1,0}^{i_{1}}Q_{2,1}^{i_{2}}$ is clearly a Dickson polynomial. The lemma follows. ∎
The next step is to show that one only needs part of the Dickson algebra. The following observation about the Dickson monomials $Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}$ at the "edge" of $\Delta^{m}_{2}$ is crucial.
Proposition 4.2.
For each $10\leq i\leq m-$ 1, we have the following decomposition in ${\mathcal{Q}}_{m}(2)$:
$Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}=\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})+\text{essential monomials divisible by $Q_{2,0}^{\frac{q^{i+1}-1}{q-1}}$}.$
Proof.
It is easy to see, by determinant manipulation, that $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})=y_{\frac{q^{i}-1}{q-1}}$ is a genuine Dickson polynomial. When $i=0$, the Dickson monomial $y_{0}$ must contains $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ as a montrivial summand since both are reduced to $x_{1}^{q^{m}-q}$ when setting $x_{2}=0$. Furthermore, $y_{0}-Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ is divisible by $Q_{2,0}$. Note that in degree $q^{m}-q$, $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ is the only non-essential Dickson monomial. We have proved that
$y_{0}=Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}+\text{essential monomials divisible by $Q_{2,0}$}.$
For the general case, we make use of the identity $y_{m,\frac{q^{i}-1}{q-1}}=Q_{2,0}^{\frac{q^{i}-1}{q-1}}y_{m-i,0}^{q^{i}}$ where $y_{m,s}$ stands for $\delta_{2;m}(Q_{1,0}^{s})$. It follows that
$y_{\frac{q^{i}-1}{q-1}}=Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}+(fQ_{2,0})^{q^{i}}Q_{2,0}^{\frac{q^{i}-1}{q-1}},$
for some Dickson polynomial $f$. A Dickson monomial appearing in the second term of the sum above has the form $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$ where $i_{2}\geq\frac{q^{i+1}-1}{q-1}$. Comparing the degree, we have $i_{1}\leq\frac{q^{m-1}-q^{i+1}}{q-1}$, hence $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$ is essential. ∎
Remark 4.3.
It is easy to check that if $i>[m-1]_{q}$, then $Q_{2,0}^{i}=L_{2}^{i(q-1)}=0$ in ${\mathcal{Q}}_{m}(2)$. Also, by induction on $m$, we have another amusing formula:
$Q_{2,1}^{\frac{2(q^{m-1}-1)}{q-1}}=y_{\frac{q^{m}-q}{q-1}}=x_{1}^{q^{m}-q}x_{2}^{q^{m}-q}\quad\text{in ${\mathcal{Q}}_{m}(2)$}.$
We are now ready to finish the proof of Proposition 1.6.
Consider a Dickson monomial $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$ and suppose, by way of contradiction, that it is not in the span of $\mathcal{B}_{m}(2)$. Further, assume that it is the smallest monomial in grevlex order with this property. Since $Q_{2,0}^{k}=0$ if $k>[m-1]_{q}$, we must have $i_{2}\leq\frac{q^{m-1}-1}{q-1}=[m-1]_{q}$. Let $i\leq m-1$ be the unique integer such that $[i]_{q}\leq i_{2}<[i+1]_{q}$. Then $i_{1}\geq\frac{q^{m-1}-q^{i}}{q-1}$. It follows from Proposition 4.2 that
$Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}-Q_{2,1}^{i_{1}-\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{i_{2}-\frac{q^{i}-1}{q-1}}y_{\frac{q^{i}-1}{q-1}}$
is a sum of Dickson monomials in strictly smaller grevlex order than $(i_{1},i_{2})$, and at least one of them is again not in the span of $\mathcal{B}_{m}(2)$. We have a contradiction.
5. Proof of Proposition 1.7
It is straightforward to check that all elements in the six families listed in Proposition 1.7 are $P(2,1)$-invariants. It suffices to show that $\mathcal{B}_{m}(2,1)$ is a generating set for ${\mathcal{Q}}_{m}^{P(2,1)}(3)$.
Suppose $f$ is a nonzero polynomials in ${\mathcal{Q}}_{m}^{P(2,1)}(3)$. If $f$ is considered as a polynomial in $x_{3}$, then its coefficients are $\operatorname{GL}_{2}$-invariant in the $2$ variables $x_{1},x_{2}$, which are classified in Section 4. If one of these coefficients has the form $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$, then the corresponding term belongs to the last 2 families in the list. By subtracting it if necessary, let us assume that this term does not occur in $f$.
Now suppose that the $x_{1}$ (and hence $x_{2}$ by symmetry) degree of $f$ is strictly less than $q^{m}-1$. Then $f$ is a linear combination of polynomials of the last two types in the list of Borel invariants in 2.5. In particular, the $x_{3}$-degree of $f$ must be divisible by $q^{2}-q$. Write $f$ as
$f=x_{3}^{aq(q-1)}f_{aq}+x_{3}^{(aq-1)(q-1)}f_{aq-1}+\ldots,$
where $f_{i}=f_{i}(x_{1},x_{2})\in{\mathcal{Q}}_{m}^{G}(2)$. If $f_{aq}$ contains $y_{t}$ for some $t\geq 0$, then we can subtract from $f$ an appropriate polynomial in the third or fourth family in the list 1.7. So we can assume that $f_{aq}$ belongs to $\Delta^{m}_{2}$. But this implies that this top degree monomial is a non-trivial summand of a $B$-invariant polynomial of the form $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$. In particular, this shows that $a$ is divisible by $q$. Write $a=bq$. Then by subtracting from $f$ an appropriate linear combination of first two families of $\mathcal{B}_{m}(2,1)$, we will obtain a $G(2,1)$-invariant polynomial of strictly smaller $x_{3}$-degree. The result follows by induction on $x_{3}$-degree. ∎
Remark 5.1.
We can also prove that $\mathcal{B}_{m}(2,1)$ is linearly independent as follows. Consider these elements of $\mathcal{B}_{m}(2,1)$ as polynomials in $x_{3}$. The table below records the coefficient of the highest $x_{3}$-exponent.
$\mathcal{B}_{m}(2,1)$ $x_{3}$-degree Coefficient
(1) $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}Q_{3,2}^{i}$ $i(q^{3}-q^{2})<q^{m}-q^{2}$ $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$, $(i_{1},i_{2})\in\Delta^{m}_{2}$
(2) $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}\delta_{3}(1)$ $q^{m}-q^{2}$ $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$, $(i_{1},i_{2})\in\Delta^{m}_{2}$
(3) $\delta_{2}(Q_{1,0}^{i_{1}}Q_{2,1}^{i})$ $i(q^{2}-q)<q^{m}-q$ $\delta_{2}(Q_{1,0}^{i_{1}})$, $i_{1}<[m]_{q}$
(4) $\delta_{2}(Q_{1,0}^{i_{1}}\delta_{2}(1))$ $q^{m}-q$ $\delta_{2}(Q_{1,0}^{i_{1}})$, $i_{1}<[m]_{q}$
(5) $\delta_{1}(\delta_{1}(Q_{1,0}^{i}))$ $i(q-1)<q^{m}-1$ $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$
(6) $\delta_{1}(\delta_{1}(\delta_{1}(1)))$ $q^{m}-1$ $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$
From the table, it is easy to see that the polynomials in $\mathcal{B}_{m}(2,1)$ are linearly independent.
6. Proof of Proposition 1.8
It suffices to show that $\mathcal{B}_{m}(1,2)$ is a generating set, and the proof is by downward induction on the smallest $x_{1}$-degree. Write
$f=x_{1}^{i(q-1)}g_{i}(x_{2},x_{3})+x_{1}^{(i+1)(q-1)}g_{i+1}(x_{2},x_{3})+\ldots$
Note that all coefficient polynomials $g_{j}$ are $\operatorname{GL}_{2}$-invariant in $x_{2}$ and $x_{3}$. If $i(q-1)=q^{m}-1$, then $x_{1}^{q^{m}-1}g(x_{2},x_{3})$ is a linear combination of polynomials from the last 3 families of $\mathcal{B}_{m}(1,2)$. We can then subtract from $f$ this polynomial and can now assume that $i(q-1)\leq q^{m}-q$.
Since $f$ is a $B$-invariant, it is a linear combination of polynomials of the form $x_{1}^{j_{1}(q-1)}\delta_{2}(Q_{2,1}^{j_{2}})$ or $Q_{1,0}^{j_{1}}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$. By direct inspection, the lowest $x_{1}$-degree of these polynomials are $x_{1}^{j_{1}(q-1)}x_{2}^{q^{m}-q}x_{3}^{j_{2}q(q-1)}$ and $x_{1}^{j_{1}(q-1)}x_{2}^{j_{2}q(q-1)}Q_{2,1}^{qj_{3}}(x_{2},x_{3})$ respectively. In both cases, the exponent of $x_{2}$ and $x_{3}$ in the coefficient of the lowest $x_{1}$-degree is divisible by $q(q-1)$.
We conclude that $g_{i}(x_{2},x_{3})$ is a polynomial in which all exponents of $x_{2}$ and $x_{3}$ are divisible by $q(q-1)$. Since $g_{i}\in{\mathcal{Q}}_{m}^{G}(x_{2},x_{3})$, it must be a $q$-th power of a polynomial in ${\mathcal{Q}}_{m-1}(2)$ in the variables $x_{2},x_{3}$.
On the other hand, the lowest $x_{1}$-degree of the first 3 families of $\mathcal{B}_{m}(1,2)$ are
  1. (1)
    $x_{1}^{j_{1}(q-1)}[Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})]^{q}$, $(i_{1},i_{2})\in\Delta^{m-1}_{2}$;
  2. (2)
    $x_{1}^{j_{1}(q-1)}[\delta_{2;m-1}(Q_{1,0}^{j_{2}})]^{q}$, $j_{2}<[m-1]_{q}$;
  3. (3)
    $x_{1}^{j_{1}(q-1)}[x_{2}^{q^{m-1}-1}x_{3}^{q^{m-1}-1}]^{q}$.
Hence $x_{1}^{i(q-1)}g_{i}(x_{2},x_{3})$ is the lowest $x_{1}$-degree of a polynomial in the span of $\mathcal{B}_{m}(1,2)$. So by subtracting from $f$ an appropriate linear combination of polynomials in $\mathcal{B}_{m}(1,2)$, we obtain another $P(1,2)$-invariant polynomial whose smallest $x_{1}$-degree is strictly greater than that of $f$. By induction, the required result follows. ∎
Remark 6.1.
We can also prove directly $\mathcal{B}_{m}(1,2)$ is linearly independent. In fact, the generators can be distinguished by looking at the lowest $x_{1}$-degrees and the corresponding coefficients as given in the table below.
$\mathcal{B}_{m}(1,2)$ $x_{1}$-degree Coefficient
(1) $Q_{1,0}^{j_{1}}Q_{3,2}^{i_{1}}Q_{3,1}^{i_{2}}$ $j_{1}(q-1)<q^{m}-1$ $Q_{2,1}^{qi_{1}}(x_{2},x_{3})Q_{2,0}^{qi_{2}}(x_{2},x_{3})$, $(i_{1},i_{2})\in\Delta^{m-1}_{2}$
(2) $Q_{1,0}^{j_{1}}\delta_{3}(Q_{2,1}^{j_{2}})$ $j_{1}(q-1)<q^{m}-1$ $y_{m-1,j_{2}}^{q}(x_{2},x_{3})$, $j_{2}<[m-1]_{q}$
(3) $Q_{1,0}^{j_{1}}\delta_{2}(\delta_{2}(1))$ $j_{1}(q-1)<q^{m}-1$ $x_{2}^{q^{m}-q}x_{3}^{q^{m}-q}$
(4) $\delta_{1}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$ $q^{m}-1$ $Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})$, $(i_{1},i_{2})\in\Delta^{m}_{2}$
(5) $\delta_{1}(\delta_{2}(Q_{1,0}^{j_{1}}))$ $q^{m}-1$ $y_{j_{1}}(x_{2},x_{3}),j_{1}<[m]_{q}$
(6) $\delta_{1}(\delta_{1}(\delta_{1}(1)))$ $q^{m}-1$ $x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$
7. Proof of Proposition 1.9
We need to prove that $\mathcal{B}_{m}(3)$ is a generating set for ${\mathcal{Q}}_{m}^{G}(3)$. The proof’s strategy is the same as in the rank $2$ case, but the details are much more involved. The proof is divided into three steps. Using the transfer, we first construct a slightly bigger generating set $\mathcal{B}^{\prime}$ of the required form without restricting the Dickson polynomials:
$\mathcal{B}^{\prime}=\delta_{1}^{3}(\Delta^{m}_{0})\coprod\delta_{2}^{2}(\Delta^{m}_{1})\coprod\delta_{3}({\mathcal{D}}_{2})\coprod{\mathcal{D}}_{3}.$
Once this is done, we then show that $\delta_{3}({\mathcal{D}}_{2})$ is actually contained in the span of
$\delta_{1}^{3}(\Delta^{m}_{0})\coprod\delta_{2}^{2}(\Delta^{m}_{1})\coprod\delta_{3}(\Delta^{m}_{2}),$
and as a result, ${\mathcal{D}}_{2}$ can be replaced by the much smaller subset $\Delta^{m}_{2}$. Finally, we prove that when restricted to ${\mathcal{Q}}_{m}(3)$, the rank $3$ Dickson algebra belongs to the subspace spanned by $\mathcal{B}_{m}(3)$. This is proved by first showing that all the edge monomials of $\Delta^{m}_{3}$ are in this span. Thus the subspace spanned by $\mathcal{B}_{m}(3)$ is a ${\mathcal{D}}_{3}$-submodule of ${\mathcal{Q}}_{m}^{G}(3)$. Since it contains $\Delta^{m}_{3}$, it will contain the entire ${\mathcal{D}}_{3}$.
Lemma 7.1.
The set $\mathcal{B}^{\prime}$ consisting of 4 families below is a generating set for ${\mathcal{Q}}_{m}^{G}(3)$.
  1. (1)
    $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$.
  2. (2)
    $\delta_{2}^{2}(\Delta^{m}_{1})=\{a_{m,3,s}$, $0\leq s<[m]_{q}\}$.
  3. (3)
    $\delta_{3}({\mathcal{D}}_{2})$.
  4. (4)
    ${\mathcal{D}}_{3}$.
Proof.
We use the relative transfer from $P(1,2)$ to $G$. Among the six families in $\mathcal{B}_{m}(1,2)$, the last family is already a $G$-invariant. So are $\delta_{3}(Q_{2,1}^{j_{2}})$ and $\delta_{2}^{2}(1)$. On the other hand, $\operatorname{Tr}(x_{1}^{j_{1}(q-1)})$ is clearly an element of ${\mathcal{D}}_{3}$ and is a multiple of $Q_{3,0}$. It is known that
$Q_{3,0}\delta_{3}(Q_{2,1}^{j_{2}})=Q_{3,0}\delta_{2}^{2}(1)=0\quad\text{in ${\mathcal{Q}}_{m}$}.$
It follows that the transfer of the second and third family in $\mathcal{B}_{m}(1,2)$ contribute $G$-invariants of the form $\delta_{3}(Q_{2,1}^{j_{2}})$, $j_{2}<[m-1]_{q}$ and $\delta_{2}^{2}(1)$. The transfer of the 4th family clearly belongs to ${\mathcal{D}}_{3}$, so it remains to consider the 5th family.
We claim that if $j<[m]_{q}$, then
$\operatorname{Tr}_{P(1,2)}^{G}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=a_{m,3,j+1}.$
Indeed, consider $x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3})$ as a $B_{3}$-invariant and taking the transfer from $B_{3}$ to $P(2,1)$. In Section 4, we have already proved that
$\operatorname{Tr}_{B_{2}}^{\operatorname{GL}_{2}}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})=y_{a+1}(x_{1},x_{2}).$
Hence
$\operatorname{Tr}_{B_{3}}^{P(2,1)}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=\sum_{a=j}^{(q^{m}-q)/(q-1)}y_{a+1}(x_{1},x_{2})x_{3}^{q^{m}-q+j(a-1)-a(q-1)}.$
The last sum is exactly $a_{m,3,j+1}=\delta_{2}^{2}(Q_{1,0}^{j+1})$, which is already a $\operatorname{GL}_{3}$-invariant. By the transitivity property of the transfer, we have the required claim. The proof is finished. ∎
The next step is to show that in the third family, one need not use the entire ${\mathcal{D}}_{2}$. The following technical lemma will be used repeatedly.
Lemma 7.2.
Let $s,t$ be non-negative integers and $0\leq i<[m]_{q}$. Then the following equality holds in ${\mathcal{Q}}_{m}(3)$:
$\delta_{2}\big{(}V_{1}^{s(q-1)}V_{2}^{t(q-1)}\delta_{2}(Q_{1,0}^{i})\big{)}=\begin{cases}0&\text{if $t\geq 1$ and $s>1$};\\ x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-1)(q-1)}&\text{if $t\geq 1$ and $s=1$};\\ (t+1)x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-2)(q-1)}&\text{if $t\geq 1$ and $s=0$};\\ \delta_{2}^{2}(Q_{1,0}^{i+s})&\text{if $t=0,s\leq 1$};\\ \delta_{2}^{2}(Q_{1,0}^{i+s})-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(s+i-2)(q-1)}&\text{if $t=0,s\geq 2$}.\end{cases}$
Proof.
Recall that $\delta_{s+1}(f)$ is a polynomial if $f$ is a polynomial (possibly with higher than $s$ variables), which is $\operatorname{GL}_{s}$-invariant. Hence the expression under consideration is a polynomial. If $t\geq 1$, then by definition, $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{t(q-1)}\delta_{2;m}(Q_{1,0}^{i}))$ equals
$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{2}^{q^{m}}x_{1}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{1},x_{3})\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{1}^{i(q-1)}\end{vmatrix}-\\ x_{1}^{q^{m}}x_{2}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{2},x_{3})\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{2}^{i(q-1)}\end{vmatrix}\bigg{]}\\ =x_{3}^{q^{m}}\;\frac{x_{2}^{q^{m}}x_{1}^{(s+i)(q-1)+1}V_{2}^{t(q-1)-1}(x_{1},x_{3})-x_{1}^{q^{m}}x_{2}^{(s+i)(q-1)+1}V_{2}^{t(q-1)-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}-\\ x_{1}^{q^{m}}x_{2}^{q^{m}}x_{3}^{i(q-1)+1}\;\frac{x_{1}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{1},x_{3})-x_{2}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}.$
The first summand is clearly trivial in ${\mathcal{Q}}_{m}$ because the quotient involved is a polynomial. For the second summand, it is also trivial if $s>1$, and is $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-1)(q-1)}$ if $s=1$. If $s=0$, the result is
$-(t(q-1)-1)x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-2)(q-1)}.$
Now suppose $t=0$. Then we have
$\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{2}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{1}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\big{]}\\ =\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{2}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{1}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\big{]}+\\ \frac{x_{1}^{q^{m}}x_{2}^{q^{m}}x_{3}^{i(q-1)}}{L_{2}(x_{1},x_{2})}\Bigg{[}\frac{x_{2}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{2}^{q-1}-x_{3}^{q-1}}-\frac{x_{1}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{1}^{q-1}-x_{3}^{q-1}}\Bigg{]}$
The first summand is $\delta_{2}^{2}(Q_{1,0}^{s+i})$. In ${\mathcal{Q}}_{m}(3)$, the second summand is trivial when $s\leq 1$ and equals $-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(s-2+i)(q-1)}$ if $s\geq 2$. ∎
We are now ready to show that
Lemma 7.3.
The subspace of ${\mathcal{Q}}_{m}^{G}(3)$ spanned by the following three families of elements
  1. (1)
    $\delta_{1}^{3}(\Delta^{m}_{0})$,
  2. (2)
    $\delta_{2}^{2}(\Delta^{m}_{1})$,
  3. (3)
    $\delta_{3}(\Delta^{m}_{2})$,
contains $\delta_{3}({\mathcal{D}}_{2})$.
Proof.
Suppose $f$ is a Dickson monomial in ${\mathcal{D}}_{2}$ but is not essential. Then it can be written in the form
$Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}Q_{2,1}^{a}Q_{2,0}^{b},$
for some $10\leq i\leq m-$ 1, $0\leq a,b$. Recall from 4.2 that
$Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}=\delta_{2}(x^{q^{i}-1})+\text{essential monomials divisible by $Q_{2,0}^{\frac{q^{i+1}-1}{q-1}}$}.$
By repeating this process if necessary, we can write $f$ as a sum of essential monomials and Dickson polynomials of the form
$\delta_{2}(x^{q^{j}-1})Q_{2,1}^{a_{j}}Q_{2,0}^{b_{j}},\quad j\geq i.$
Thus it suffices to prove the lemma for $f=g\delta_{2}(x^{q^{i}-1})$, where $g\in{\mathcal{D}}_{2}$ and $10\leq i\leq m-$ 1.
To evaluate $\delta_{3}(g\delta_{2}(x^{q^{i}-1}))$, we first observe that by direct inspection, it can be simplified as a determinant:
$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))=\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{2}^{q^{m}}&x_{3}^{q^{m}}\\ x_{1}^{q^{m}}g(x_{2},x_{3})&x_{2}^{q^{m}}g(x_{1},x_{3})&x_{3}^{q^{m}}g(x_{1},x_{2})\end{vmatrix}}{L_{3}(x_{1},x_{2},x_{3})}.$
Since
$L_{3}(x_{1},x_{2},x_{3})=$ $V_{3}(x_{1},x_{2},x_{3})L_{2}(x_{1},x_{2})$
$=$ $[V_{2}(x_{2},x_{3})^{q-1}-V_{2}(x_{2},x_{1})^{q-1}]V_{2}(x_{2},x_{3})L_{2}(x_{1},x_{2})$
$=$ $[V_{2}(x_{1},x_{3})^{q-1}-V_{2}(x_{1},x_{2})^{q-1}]V_{2}(x_{1},x_{3})L_{2}(x_{1},x_{2}),$
this determinant can be rewritten in the form:
$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}h(x_{1},x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}h(x_{2},x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\bigg{]}.$
where
$h(x_{1},x_{2},x_{3})=\frac{g(x_{2},x_{3})-g(x_{2},x_{1})}{V_{2}(x_{2},x_{3})^{q-1}-V_{2}(x_{2},x_{1})^{q-1}}.$
This quotient is a symmetric polynomial in $V_{2}^{q-1}(x_{2},x_{3})$ and $V_{2}^{q-1}(x_{2},x_{1})$ with coefficient in $\mathbb{F}_{q}[x_{2}^{q-1}]$. Moreover, if one summand of this quotient is divisible by $V_{2}(x_{2},x_{1})^{q-1}$, then the common denominator $L_{2}(x_{1},x_{2})$ will be canceled out, and the result is zero in ${\mathcal{Q}}_{m}$.
Thus if we write $g$ in terms of the upper triangular invariants, $g=\sum_{s,t}V_{1}^{s(q-1)}V_{2}^{t(q-1)}$, then the determinant above equals in ${\mathcal{Q}}_{m}$ to a sum over all such pair $(s,t)$, $t\geq 1$ of
$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}V_{2}^{(t-1)(q-1)}(x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}(x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\bigg{]},$
which is $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$. We can now use the computation in Lemma 7.2 to see that
$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))\in\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{2}(Q_{1,0}^{a}),a\geq 0)\}.$
On the other hand, $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})$ is a genuine Dickon polynomial, hence $\delta_{3}(g\delta_{2}(x^{q^{i}-1}))$ is a $\operatorname{GL}_{3}$-invariant in ${\mathcal{Q}}_{m}$. This shows that
$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))\in\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{3}(1)\}.$
The third, and final step is to replace the full Dickson algebra ${\mathcal{D}}_{3}$ by $\Delta^{m}_{3}$. We first need some preliminary calculations. In the following, we write $\delta_{3;a}(i_{1},i_{2})$ for $\delta_{3;a}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$ to simplify the notation. Again $\delta_{3}$ will stand for $\delta_{3;m}$.
Lemma 7.4.
The following equalities hold:
  1. (1)
    $\delta_{3}(0,\frac{q^{i}-1}{q-1})=\delta_{3;m-i}^{q^{i}}(1)Q_{3,0}^{\frac{q^{i}-1}{q-1}}$ for all $i\geq 0$. In particular,
    $\delta_{3}(0,{\frac{q^{m-2}-1}{q-1}})=Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}\text{ and }\delta_{3}(0,\frac{q^{m-3}-1}{q-1})=Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}.$
  2. (2)
    $Q_{3,0}^{k}=0$ in ${\mathcal{Q}}_{m}$ for all $k>\frac{q^{m-2}-1}{q-1}$.
  3. (3)
    $\delta_{3}(q^{m-3},\frac{q^{m-3}-1}{q-1})=Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$.
  4. (4)
    In general, for $\lambda_{2}\geq\lambda_{3}\geq 0$, then
    $\delta_{3}\bigg{(}\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1}\bigg{)}=\delta_{3;m-\lambda_{3}}^{q^{\lambda_{3}}}\bigg{(}\frac{q^{\lambda_{2}-\lambda_{3}}-1}{q-1},0\bigg{)}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}.$
Proof.
The proof of these statements are quite similar, and essentially makes use of the Laplace expansion. We will prove the last statement only. Apply the Laplace expansion along the last row of the determinant in the numerator for $\delta_{3}(\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1})$, we obtain
$x_{1}^{q^{m}}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}(x_{2},x_{3})L_{2}^{q^{\lambda_{3}}}(x_{2},x_{3})-x_{2}^{q^{m}}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}(x_{1},x_{3})L_{2}^{q^{\lambda_{3}}}(x_{1},x_{3})+x_{3}^{q^{m}}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}(x_{1},x_{2})L_{2}^{q^{\lambda_{3}}}(x_{1},x_{3}).$
The result follows easily after writing this expression as a $q^{\lambda_{3}}$ power. ∎
Our second technical result plays a similar role as Proposition 4.2 in the case of 3 variables.
Proposition 7.5.
Suppose $20\leq\lambda\leq m-$ 2, then in ${\mathcal{Q}}_{m}$, $\delta_{3}(\frac{q^{\lambda}-1}{q-1},0)$ equals a Dickson polynomial in which the only non-essential summand which is not divisible by $Q_{3,0}$ is $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,1}^{\frac{q^{\lambda}-1}{q-1}}$.
Proof.
If $\lambda=0$, then $\delta_{3}(1)$ is a genuine Dickson polynomial in degree $q^{m}-q^{2}$. In this degree, there is only one Dickson monomial not in $\Delta^{m}_{3}$, namely $Q_{3,2}^{\frac{q^{m-2}-1}{q-1}}$. If we set $x_{2}=x_{3}=0$, then both $Q_{3,1}$ and $Q_{3,0}$ becomes zero, $\delta_{3}(1)$ becomes $x_{1}^{q^{m}-q^{2}}$. Thus in the expansion of $\delta_{3}(1)$ into a sum of Dickson monomials, $Q_{3,2}^{\frac{q^{m-2}-1}{q-1}}$ must occur with non-zero coefficient. The lemma is proved for $\lambda=0$.
We now assume that $\lambda\geq 1$. Note that $\delta_{3}(Q_{2,1})=\frac{[0,2,m]}{[0,1,2]}$ is a genuine Dickson polynomial, hence in ${\mathcal{Q}}_{m}$,
$\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})=Q_{3,2}^{\frac{q^{\lambda-1}-1}{q-1}}\delta_{3}(Q_{2,1})=Q_{3,2}^{\frac{q^{\lambda-1}-1}{q-1}}\frac{[0,2,m]}{[0,1,2]}.$
So $\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ equals in ${\mathcal{Q}}_{m}$ to a Dickson polynomial and we are going to use this representation.
Setting $x_{3}=0$, $\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ is reduced to
$\frac{[\lambda+1,m]}{[1,2]}=y^{q^{\lambda+1}}_{m-\lambda-1,0}Q_{2,0}^{\frac{q^{\lambda+1}-q}{q-1}}=Q_{2,1}^{\frac{q^{m-1}-q^{\lambda+1}}{q-1}}Q_{2,0}^{\frac{q^{\lambda+1}-q}{q-1}}+\text{others}.$
It follows that in this representation in ${\mathcal{Q}}_{m}$ of $\delta_{3;m}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ as a Dickson polynomial, $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,1}^{\frac{q^{\lambda-1}-1}{q-1}}$ must occur with nonzero coefficient. Moreover, any other monomials which involve $Q_{3,1}$ and $Q_{3,2}$ only must be divisible by $Q_{3,1}^{\frac{q^{\lambda}-1}{q-1}}$, and by degree reason, these monomials must be essential. ∎
The following corollary will be useful for induction.
Corollary 7.6.
Suppose $m-3\leq\lambda_{2}\leq\lambda_{3}\leq 0$ and $\lambda\leq m-2$.
  1. (1)
    $\delta_{3}(\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1})=Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}+\text{other terms},$
    where the other terms are either essential or non-essential and divisible by $Q_{3,0}^{\frac{q^{\nu}-1}{q-1}}$, where $\nu>\lambda_{3}$.
  2. (2)
    $\delta_{3}(\frac{q^{m-2}-q^{\lambda}}{q-1},\frac{q^{\lambda}-1}{q-1})=Q_{3,1}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,0}^{\frac{q^{\lambda}-1}{q-1}}+\text{other terms},$
    where the other terms are either essential or non-essential and divisible by $Q_{3,0}^{\frac{q^{\nu}-1}{q-1}}$ where $\nu>\lambda$.
Proof.
This is just combining Lemma 7.4 and Proposition 7.5. ∎
Corollary 7.7.
The Dickson monomials at the edge of $\Delta^{m}_{3}$ listed below are also in the span of $\mathcal{B}_{m}(3)$:
  1. (1)
    $Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}$,
  2. (2)
    $Q_{3,1}^{\frac{q^{m-2}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$, where $m-3\geq\lambda_{3}$,
  3. (3)
    $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$, where $m-3\geq\lambda_{2}\geq\lambda_{3}$.
Proof.
We have already seen from part (1) of Lemma 7.4 that $Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}$ belongs to $\Delta^{m}_{3}$. For the other 2 families, we will prove by downward induction on $\lambda_{3}$.
If $\lambda_{3}=m-3$, then $\lambda_{2}=\lambda_{3}=m-3$. The edge monomials $Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$ and $Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$ belong to the span of $\mathcal{B}_{m}(3)$ by the last 2 parts of Lemma 7.4. We now suppose that the results have been verified for all edge elements corresponding to $\lambda_{3}+1$ and above. Consider the edge elements corresponding to $\lambda_{3}$:
  1. (1)
    $Q_{3,1}^{\frac{q^{m-2}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$,
  2. (2)
    $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$.
Corollary 7.6 shows that these are in the image of $\delta_{3}(\Delta^{m}_{2})$ modulo those non-essential elements with strictly higher $\lambda_{3}$. By induction, these elements are already in the span of $\mathcal{B}_{m}(3)$. ∎
Now we are ready to finish the proof. Corollary 7.7 shows that the subspace of ${\mathcal{Q}}_{m}^{G}(3)$ spanned by $\mathcal{B}_{m}(3)$ is closed under the action of the Dickson algebra ${\mathcal{D}}_{3}$. Indeed, we have already computed the action of ${\mathcal{D}}_{3}$ on $\delta_{s+1}^{3-s}(\Delta^{m}_{s})$ for $0\leq s\leq 2$ in 2.5, and the product of $Q_{3,i}$ with a monomial in $\Delta^{m}_{3}$ is clearly an edge monomial, which is in the span of $\mathcal{B}_{m}(3)$ by the above demonstration. The proof is finished.
8. A filtration of submodules
For each positive integer $n$, $0\leq k\leq n$, let $\mathcal{F}_{n,k}$ denote the $\mathbb{F}_{q}$-subspace of ${\mathcal{Q}}^{G}_{m}(n)$:
$\mathcal{F}_{n,k}=\operatorname{Span}(\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq\min(m,k)).$
For $n=3$, we have shown that this increasing filtration of subspaces is exhaustive. The calculation in previous sections shows that our basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{G}(3)$ is natural in the sense that the resulting filtration is a filtration of ${\mathcal{D}}_{3}$-submodules. Moreover, the action of the Dickson invariants is very similar to the pattern discussed in [10] about the structure of the cofixed space $S_{G}$ as a $S^{G}$-module. Indeed, we can easily check that $Q_{3,0}$ annihilates $\mathcal{F}_{3,2}$; $Q_{3,0}$ and $Q_{3,1}$ annihilate $\mathcal{F}_{3,1}$; and $Q_{3,0},Q_{3,1},Q_{3,2}$ annihilate $\mathcal{F}_{3,0}$. A similar pattern also occurs in the case $n=2$.
The aim of this final section is to demonstrate Theorem 1.11. We will show the filtration $\mathcal{F}_{n,*}$ also behaves well with respect to the action of the mod $q$ Steenrod algebra $\mathcal{A}$, at least in the case $n\leq 3$.
For background on the mod $q$ Steenrod algebra and its role in modular invariant theory, we refer to Larry Smith’s article [16], which treats the Steenrod algebra completely algebraically over a general finite field.
Recall that for each $i\geq 0$, the $i$th Steenrod reduced power operation $\mathcal{P}^{i}$ is a $\mathbb{F}_{q}$-linear map satisfying the following conditions:
  1. (1)
    The unstable condition: $\mathcal{P}^{i}(f)=f^{q}$ if $i=\deg(f)$ and $\mathcal{P}^{i}(f)=0$ if $i>\deg(f)$.
  2. (2)
    The Cartan formulae: $\mathcal{P}^{i}(fg)=\sum_{a+b=i}\mathcal{P}^{a}(f)\mathcal{P}^{b}(g)$.
Moreover, $\mathcal{P}^{0}$ acts as the identity. Over the polynomial ring $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, in which $x_{i}$ are given degree $1$, the action of the Steenrod algebra is completely determined by the unstable condition and the Cartan formulae. More concretely, we have
$\mathcal{P}^{i}(v^{j})=\binom{j}{i}v^{j+i(q-1)},$
for any linear polynomial $v$. Here the binomial coefficient is taken modulo $p$. Thus, if $j$ is a $q$-power, then $\mathcal{P}^{i}(v^{j})\neq 0$ iff $i=0$ (in which case $\mathcal{P}^{0}$ is the identity operator) or $i=j$ (in which case $\mathcal{P}^{i}(v^{i})=v^{qi}$ is the Frobenius operator).
The Steenrod algebra action and the $G$ action on $S$ commutes. Furthermore, it is clear that $I_{m}$ is stable under the action of the Steenrod operations, so is ${\mathcal{Q}}_{m}=S/I_{m}$. It follows that a Steenrod reduced power yields a map from ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ to itself. In other words, ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is an $\mathcal{A}$-module.
The proof of Theorem 1.11 uses several lemmas. We begin with a general formula describing how the Steenrod operation interacts with the delta operator.
Lemma 8.1.
If $f$ is a polynomial which is $\operatorname{GL}_{1}$-invariant in the first variable and $1\leq k\leq\deg(\delta_{2}(f))=q^{m}-q+\deg(f)$, then
$\mathcal{P}^{k}(\delta_{2}f)+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}f)+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}f)=\\ \delta_{2}(Q_{1,0}\mathcal{P}^{k-1}f)+\delta_{2}(\mathcal{P}^{k}f)+\delta_{2;m+1}(Q_{1,0}\mathcal{P}^{k-1-q^{m}}f)+\delta_{2;m+1}(\mathcal{P}^{k-q^{m}}f).$
Moreover, the last two summands on the right-hand side vanish in ${\mathcal{Q}}_{m}$.
Proof.
The condition imposed on $f$ assures that $\delta_{2}(f)$ as well as the other terms in this lemma are polynomial. The equation above is then obtained by applying the Cartan formula for the product $L_{2}\delta_{2}(f)$ and then dividing both sides by $L_{2}$. Note that $\mathcal{P}^{q}L_{2}=Q_{2,1}L_{2}$, $\mathcal{P}^{q+1}L_{2}=Q_{2,0}L_{2}$ and $\mathcal{P}^{k}L_{2}=0$ for all $k\geq 1$. ∎
Corollary 8.2.
For each $0\leq k\leq 2$, $\mathcal{F}_{2,k}$ is a submodule of ${\mathcal{Q}}_{m}^{G}(2)$ over the Steenrod algebra.
Proof.
We need to compute the action of the Steenrod operation on a typical element $\delta_{2}(f)$ where $f=Q_{1,0}^{a}$, $a<[m]_{q}$. In the rank $2$ case, the last two summands on the right-hand side of the formula given in Lemma 8.1 vanish in ${\mathcal{Q}}_{m}$. Indeed, one check by direct inspection that the degree of $\delta_{2;m+1}(Q_{1,0}\mathcal{P}^{k-1}f)$ is strictly greater than $2(q^{m}-1)$ for all $q$ and $m$. For $\delta_{2;m+1}(\mathcal{P}^{k-q^{m}}f)$, its degree is not greater than $2(q^{m}-1)$ only when $q=2$, $k\geq q^{m}$ and $\deg f=0$, but in this case $k>q^{m}-q+\deg(f)$.
Now we can proceed by induction on $k$ using the above formula and the calculation in Proposition 2.5. ∎
Corollary 8.3.
$\mathcal{F}_{3,1}$ is a submodule of ${\mathcal{Q}}_{m}^{G}(3)$ over the Steenrod algebra.
Proof.
We need to show that for each $k\geq 1$ and $a<[m]_{q}$, $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ is in the span of $\delta_{1}^{3}(1)$ and $\delta_{2}^{2}(\Delta^{m}_{1})$.
Apply the formula provided in Lemma 8.1 for $f=\delta_{2}Q_{1,0}^{a}$. Again, the last two summands are trivial in ${\mathcal{Q}}_{m}$ because of the instability condition and degree calculation. We then obtain a recursive equation in ${\mathcal{Q}}_{m}$:
$\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}Q_{1,0}^{a})+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}Q_{1,0}^{a})=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}\delta_{2}Q_{1,0}^{a})+\delta_{2}(\mathcal{P}^{k}\delta_{2}Q_{1,0}^{a}).$
We now proceed by double induction on $k$ that $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ belongs to $\mathcal{F}_{3,1}$. We already know from Corolarry 8.2 that $\mathcal{P}^{k-1}(\delta_{2}Q_{1,0}^{a})$ and $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{a})$ belongs to the span of $\delta_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}(1)$. By Lemma 7.2, the two summands on the right-hand side in the equation above belong to the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$.
Also by induction, we see that $Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}Q_{1,0}^{a})=0$ in ${\mathcal{Q}}_{m}$ and $Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}Q_{1,0}^{a})$ is also in the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$. We conclude that $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ is also in the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$, but since it is a $\operatorname{GL}_{3}$-invariant, it is spanned by $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{3}(1)$. ∎
We now consider the interaction between the Steenrod powers and the operator $\delta_{3}$.
Lemma 8.4.
If $f\in{\mathcal{D}}_{2}$ and $k\geq 1$, then
$\mathcal{P}^{k}(\delta_{3}f)+\sum_{i<k}\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)=\frac{\mathcal{P}^{k}\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{q}&x_{2}^{q}&x_{3}^{q}\\ x_{1}^{q^{m}}f(x_{2},x_{3})&x_{2}^{q^{m}}f(x_{1},x_{3})&x_{3}^{q^{m}}f(x_{1},x_{2})\end{vmatrix}}{L_{3}}.$
Moreover, the right-hand side equals $\delta_{3;m}(F)$ or $\delta_{3;m+1}(F)$ where $F$ is a certain linear combination of polynomials of the form $g\mathcal{P}^{i}f$ where $g\in{\mathcal{D}}_{2}$ and $0\leq i\leq k$.
Proof.
The above equality is a simple application of the Cartan formula for $\mathcal{P}^{k}(L_{2}\delta_{2;m}(f))$ (or $\mathcal{P}^{k}(L_{3}\delta_{3}f)$). For the right-hand side, note that because of the instability condition, we have $\mathcal{P}^{i}(x_{1})$ is either $x_{1}$ or $x_{1}^{q}$, depending on whether $i=0$ or $i=1$, and there is a similar formula for other entries in the determinant.
In the first case, $\delta_{2;m+1}(g\mathcal{P}^{i}f)$ is trivial in ${\mathcal{Q}}_{m}$ unless $q=2$ and $g\mathcal{P}^{i}f=1$. In this case, $\delta_{2;m+1}(1)=x_{1}^{2^{m}-1}x_{2}^{2^{m}-1}$ because of degree reason. We also have $g=f=1$, and $k=2^{m}$, but then $\delta_{2}f$ has degree $2^{m}-2<2^{m}$.
Finally, observe that any determinant of the form
$\begin{vmatrix}x_{1}^{q^{a}}&x_{2}^{q^{a}}&x_{3}^{q^{a}}\\ x_{1}^{q^{b}}&x_{2}^{q^{b}}&x_{3}^{q^{b}}\\ x_{1}^{q^{c}}\mathcal{P}^{d}f(x_{2},x_{3})&x_{2}^{q^{c}}\mathcal{P}^{d}f(x_{1},x_{3})&x_{3}^{q^{c}}\mathcal{P}^{d}f(x_{1},x_{2})\end{vmatrix}$
can be written as
$L_{3}\delta_{3;c}\bigg{(}\frac{[a,b]}{[0,1]}\mathcal{P}^{d}f\bigg{)}.$
where $\frac{[a,b]}{[0,1]}\in\mathcal{D}_{2}$. Note that $c$ is either $m$ or $m+1$. ∎
It is also clear that $\mathcal{P}^{i}L_{3}$ is again a matrix whose entries are $q$-powers, so it is always a multiple of $L_{3}$ and the quotient $\frac{\mathcal{P}^{i}L_{3}}{L_{3}}$ is a Dickson polynomial in ${\mathcal{D}}_{3}$. The next lemma discusses how to deal with the case when $c=m+1$.
Lemma 8.5.
If $h\in{\mathcal{D}}_{2}$, then the following holds in ${\mathcal{Q}}_{m}$,
$\delta_{3;m+1}(h)=\begin{cases}0&\text{if $q>3$ or $q=3$ and $\deg(h)>0$;}\\ \delta_{3;m}(Q_{2,1}^{q^{m-1}})&\text{if $q=3$ and $h=1$;}\\ \in\operatorname{Span}\{\delta_{2;m}^{2}(\Delta^{m}_{1}),\delta_{1;m}^{3}(1)\}&\text{if $q=2$.}\\ \end{cases}$
Proof.
The degree of $\delta_{3;m+1}(h)$ equals $q^{m+1}-q^{2}+\deg(h)$. If $q>3$, then this is strictly greater than $3(q^{m}-1)$, and we immediately have $\delta_{3;m+1}(h)\equiv 0$ in this case.
From the fundamental equation, we know that $X^{q^{3}}-Q_{3,2}X^{q^{2}}+Q_{3,1}X^{q}-Q_{3,0}X=0$ if $X$ is in the span of $x_{1},x_{2},x_{3}$. It follows that
$x_{i}^{q^{m+1}}-Q_{3,2}^{q^{m-2}}x_{i}^{q^{m}}+Q_{3,1}^{q^{m-2}}x_{i}^{q^{m-1}}-Q_{3,0}^{q^{m-2}}x_{i}^{q^{m-2}}=0,$
and so
$\delta_{3;m+1}(h)-Q_{3,2}^{q^{m-2}}\delta_{3;m}(h)+Q_{3,1}^{q^{m-2}}\delta_{3;m-1}(h)-Q_{3,0}^{q^{m-2}}\delta_{3;m-2}(h)=0.$
If $q=3$, then $q^{m+1}-q^{2}+\deg(h)\leq 3(q^{m}-1)$ implies $\deg(h)\leq q^{2}-3=6$. Since $h\in{\mathcal{D}}_{2}$, it follows that up to a scalar, $h=1$ or $h=Q_{2,1}$.
In the case where $h=Q_{2,1}$, $\delta_{3;m+1}(Q_{2,1})$ has degree $33^{m+1}-$ 3 so that it is either zero or the top element $x_{1}^{3^{m}-1}x_{2}^{3^{m}-1}x_{3}^{3^{m}-1}$. We claim that $\delta_{3;m+1}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m}$. Indeed, when $q=3$, then by direct computation, we have $Q_{3,0}=Q_{3,1}=0$ in ${\mathcal{Q}}_{2}$ so that $Q_{3,1}^{q^{m-2}}=Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}$. It follows that in ${\mathcal{Q}}_{m}$, we have
$\delta_{3;m+1}(Q_{2,1})=Q_{3,2}^{q^{m-2}}\delta_{3;m}(Q_{2,1}).$
On the other hand, $Q_{3,2}=x_{1}^{6}x_{2}^{6}x_{3}^{6}$ in ${\mathcal{Q}}_{2}$ so that $Q_{3,2}^{q^{m-2}}=x_{1}^{2.3^{m-1}}x_{2}^{2.3^{m-1}}x_{3}^{2.3^{m-1}}$ in ${\mathcal{Q}}_{m}$. So if we can show that $\delta_{3;m}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m-1}$, then it follows immediately that $\delta_{3;m+1}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m}$. The claim can now be easily verified by checking that $\delta_{3;3}(Q_{2,1})=Q_{3,1}=0$ in ${\mathcal{Q}}_{2}$.
Suppose now $h=1$. In this case, we have
$\delta_{3;m+1}(1)=Q_{3,2}^{3^{m-2}}\delta_{3;m}(1)=\delta_{3;m}(Q_{2,1}^{3^{m-1}}),$
as required. Here the second equality follows from 2.5.
Now consider the case $q=2$. We still have $Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}(3)$, but now
$Q_{3,1}^{2^{m-2}}=x_{1}^{2^{m-1}}x_{1}^{2^{m-1}}x_{1}^{2^{m-1}}.$
By direct inspection, we have $\delta_{3;m+1}(h)=Q_{3,2}^{2^{m-2}}\delta_{3;m}(h)-Q_{3,1}^{2^{m-2}}\delta_{3;m-1}(h)$ equals in ${\mathcal{Q}}_{m}$ to
$\frac{1}{L_{3}}\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{2}&x_{2}^{2}&x_{3}^{2}\\ x_{1}^{2^{m}}h(x_{2},x_{3})(x_{2}^{2^{m}}+x_{3}^{2^{m}})&x_{2}^{2^{m}}h(x_{1},x_{3})(x_{1}^{2^{m}}+x_{3}^{2^{m}})&x_{3}^{2^{m}}h(x_{1},x_{2})(x_{1}^{2^{m}}+x_{2}^{2^{m}})\end{vmatrix}.$
It is straightforward to verify that
$x_{1}^{2^{m}}+x_{2}^{2^{m}}=Q_{2,1}\delta_{2;m}(1)-\delta_{2;m}(Q_{1,0}^{2}),$
so the expression above becomes
$\delta_{3;m}(Q_{2,1}h\delta_{2;m}(1))-\delta_{3;m}(h\delta_{2;m}(Q_{1,0}^{2})).$
Now we can use the argument as in the proof of Lemma 7.3 to conclude. ∎
Remark 8.6.
With a little more work, using 7.2, it can be showed that when $q=2$, then for $h\in{\mathcal{D}}_{2}$, we have an explicit formula:
$\delta_{3;m+1}(h)=\begin{cases}0&\text{if $h$ is a multiple of $Q_{2,0}$,}\\ \delta_{2;m}^{2}(Q_{1,0}^{2s})&\text{if $h=Q_{2,1}^{s}$}.\end{cases}$
We leave the details for the interested reader.
Now we can finish the proof of Theorem 1.11. It remains to show that $\mathcal{F}_{3,1}$ is an $\mathcal{A}$-submodule of ${\mathcal{Q}}_{m}^{G}(3)$. From Lemma 8.5 we see that if $f\in{\mathcal{D}}_{2}$, then
$\mathcal{P}^{k}(\delta_{3}f)+\sum_{i<k}\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)$
belongs to $\mathcal{F}_{3,1}$. By induction, so is $\mathcal{P}^{i}(\delta_{3}f)$ for each $i<k$ and from Proposition 2.5, $\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)$ also belongs to $\mathcal{F}_{3,1}$.

### v2

@@
-On Modular Invariants of Truncated Polynomial Rings in low ranks
+On Modular Invariants of the Truncated Polynomial Rings in low ranks
 Abstract.
-We verify the conjectures due to Lewis, Reiner and Stanton about the Hilbert series of the invariant ring of the truncated polynomial ring for all parabolic subgroups up to rank $3$.
+We verify the conjectures due to Lewis, Reiner, and Stanton about the Hilbert series of the invariant ring of the truncated polynomial ring for all parabolic subgroups up to rank $3$.
 This is done by constructing an explicit set of generators for each invariant ring in question.
-We also propose a conjecture concerning the action of the Steenrod algebra and the Dickson algebra on a certain naturally occurring filtration of the invariant ring under the general linear group.
+We also propose a conjecture concerning the action of the Steenrod algebra and the Dickson algebra on a certain naturally occurring filtration of the invariant ring under the action of the general linear group.
 Key words and phrases:
-Truncated Polynomial Ring, Modular invariant, Dickson invariants, Steenrod algebra, q-binomial, q-multinomial, finite general linear group, Frobenius power
+Truncated Polynomial Ring, Modular invariants, Dickson invariants, Steenrod algebra, q-binomial, q-multinomial, finite general linear group, Frobenius power
 1991 Mathematics Subject Classification:
@@
 Let $\mathbb{F}_{q}$ denote the finite field of $q$ elements, where $q$ is a power of a fixed prime number $p$.
-Let $I$ denote the ideal generated by the $m$th Frobenius power $x_{1}^{q^{m}},x_{2}^{q^{m}},\ldots,x_{n}^{q^{m}}$, of the polynomial algebra $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$.
-We denote by ${\mathcal{Q}}_{m}(n)$ or simply ${\mathcal{Q}}_{m}$, if $n$ is fixed, the quotient ring $S/I$.
-The general linear group $G=\operatorname{GL}_{n}(\mathbb{F}_{q})$ acts on $S$ by linear substitutions of variables.
-Since $I$ is a $G$-invariant ideal, ${\mathcal{Q}}_{m}(n)$ inherits a $G$-action from $S$.
-This paper discusses the problem of computing invariants of the truncated ring ${\mathcal{Q}}_{m}(n)$ under the action of various parabolic subgroups of $G$.
-This problem arises in various contexts, in connection with questions in algebraic combinatorics ([14], [15], [10], [4]), algebraic topology ([1], [7], [9]) and number theory [2].
-In particular, taking the limit $m$ to infinity yields results about the cofixed space $S_{G}$ of the polynomial ring under the action of the general linear group.
-This problem is surprisingly still open in modular invariant theory.
-Let $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ be a composition of $n$ and denote by $P(\alpha)$ the corresponding parabolic subgroup.
-In [10], Lewis, Reiner, and Stanton proposed a marvelous series of conjectures about the Hilbert-Poincaré series of the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ for arbitrary $m$ and any composition $\alpha$ of $n$ in terms of a new combinatorial object called the $(q,t)$-binomial coefficient introduced in [15]:
+Let $I_{m}(n)=(x_{1}^{q^{m}},\ldots,x_{n}^{q^{m}})$ denote the $m$-th Frobenius power ideal in the polynomial algebra $S(n)=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$.
+Denote by ${\mathcal{Q}}_{m}(n)$ the truncated polynomial ring $S(n)/I_{m}(n)$.
+The general linear group $G_{n}=GL_{n}(\mathbb{F}_{q})$ acts on $S(n)$ by linear substitutions of variables.
+Since $I_{m}(n)$ is a $G_{n}$-invariant ideal, ${\mathcal{Q}}_{m}(n)$ inherits a $G_{n}$-action from $S(n)$.
+As we will be working with a fixed number of generators $n$, we will usually omit $n$ from the notation.
+This paper addresses the problem of computing invariants of the truncated ring ${\mathcal{Q}}_{m}(n)$ under the action of various parabolic subgroups of $G_{n}$.
+This problem arises in several contexts, including algebraic combinatorics ([11], [12], [9], [4]), algebraic topology ([1], [7], [8]) and number theory [2].
+In particular, as shown in [9, Conjecture 1.3], knowing the Hilbert series of ${\mathcal{Q}}_{m}(n)^{G}$ provides the Hilbert series of the cofixed space $S_{G}$ of the polynomial ring under the action of the general linear group.
+Surprisingly, while the invariant ring $S^{G}$ has been completely determined since the beginning of the 20th century, by the fundamental work of Dickson [3], the problem of determining $S_{G}$ is still an open problem in modular invariant theory.
+For a positive integer $n$, a composition of $n$ of length $\ell$ is an ordered set $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ of positive integers $\alpha_{i}$ such that $|\alpha|=\alpha_{1}+\ldots+\alpha_{\ell}=n$.
+We will also consider weak compositions, which are ordered sets of non-negative integers.
+The set of weak compositions of a given length is partially ordered by declaring that $\beta\leq\alpha$ iff $\beta_{i}\leq\alpha_{i}$ for all $i$.
+If $\alpha$ is a composition of $n$, we denote by $P(\alpha)$ the corresponding parabolic subgroup of $G_{n}$.
+In [9], Lewis, Reiner, and Stanton proposed a marvelous set of conjectures about the Hilbert series of the invariant ring ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$ for arbitrary $m$ and any composition $\alpha$ of $n$, expressed in terms of a new combinatorial object called the $(q,t)$-binomial coefficient, introduced in [12]:
 Conjecture 1.1.
-([10, Conjecture 1.5]) The Hilbert-Poincaré series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the parabolic subgroup $P(\alpha)$ is:
+([9, Parabolic Conjecture 1.5]) The Hilbert series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the parabolic subgroup $P(\alpha)$ is:
 $C_{\alpha,m}(t)=\sum_{\beta,\beta\leq\alpha,|\beta|\leq m}t^{e(m,\alpha,\beta)}\;\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t},$
@@
 $\begin{bmatrix}d\\ \alpha\end{bmatrix}_{q,t}=\frac{\prod_{j=0}^{d-1}(1-t^{q^{d}-q^{j}})}{\prod_{i=1}^{\ell}\prod_{j=0}^{\alpha_{i}}(1-t^{q^{A_{i}}-q^{A_{i-1}+j}})}.$
-In particular, when $\alpha=(n)$, they conjectured that
+In particular, in the most interesting case where $\alpha=(n)$, they conjectured that
 Conjecture 1.2.
-([10, Conjecture 1.2]) The Hilbert-Poincaré series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the general linear group $G=\operatorname{GL}_{n}(\mathbb{F}_{q})$ is:
+([9, Conjecture 1.2]) The Hilbert series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the general linear group $G_{n}$ is:
 $C_{n,m}(t)=\sum_{k=0}^{\min(n,m)}t^{(n-k)(q^{m}-q^{k})}\;\begin{bmatrix}m\\ k\end{bmatrix}_{q,t}.$
-These conjectures are inspired by properties of the $q$-Catalan and $q$-Fuss Catalan numbers, connecting Hilbert-Poincaré series for certain invariant subspaces with the representation theory of rational Cherednik algebras for Coxeter and complex reflection groups.
-Some partial results about these two conjectures have been obtained (see Drescher-Shepler [4], Goyal [5] and the original paper [10]).
-Still, how these combinatorial objects occur in the series remains quite mysterious.
-In our previous work [6], we have proved their parabolic conjecture for the minimal parabolic subgroup by constructing an explicit basis.
-We also refined their conjectures by proposing an explicit $\mathbb{F}_{q}$-basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ as follows:
+These conjectures are inspired by properties of the $q$-Catalan and $q$-Fuss-Catalan numbers, connecting Hilbert series of certain invariant subspaces with the representation theory of rational Cherednik algebras for Coxeter and complex reflection groups.
+Some partial results about these two conjectures have been obtained (see Drescher-Shepler [4], Goyal [5], and the original paper [9]).
+Nevertheless, the role of these combinatorial objects in the series remains quite mysterious.
+In our previous work [6], we proved their parabolic conjecture for the minimal parabolic subgroup $B$ by constructing an explicit basis for the $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}(n)^{B}$ for all $m$ and $n$.
+We also refined their conjectures by proposing an explicit $\mathbb{F}_{q}$-basis for ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$.
 Conjecture 1.3.
-A basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is given inductively by the following formula:
-$\mathcal{B}_{m}(\alpha)=\{\delta_{s+1;m}^{\alpha_{1}-s}(f\cdot\Phi^{s}g)\colon 0\leq s\leq\min(\alpha_{1},m),f\in\Delta^{m}_{s},g\in\mathcal{B}_{m-s}(\alpha_{2},\ldots,\alpha_{\ell})\}.$
-Here $\Delta^{m}_{s}$ is a certain subspace of the Dickson algebra ([3]) of all $\operatorname{GL}_{s}$-invariant polynomials and $\delta_{a;b}$ is the operator constructed in [6].
-The Frobenius-like operator $\Phi$ is given by the rules:
-  1. (i)
-    $\Phi(\delta_{a;b})=\delta_{a+1;b+1}$;
-  2. (ii)
-    $\Phi(Q_{r,i})=Q_{r+1,i+1}$ and expand multiplicatively.
-The reader of [6] may notice that in this conjecture, we have used the core subspace $\Delta^{m}_{s}$ of the Dickson algebra instead of $\nabla^{m}_{s}$ in the original conjecture.
-The two candidates can be used interchangeably, but $\Delta^{m}_{s}$ is more convenient for explicit calculation, as we will see later.
-In the case $\alpha=(n)$ so that $P(\alpha)=G$ is the full general linear group, we have
+A basis for the space of $P(\alpha)$-invariants of ${\mathcal{Q}}_{m}(n)$ is given by the set $\mathcal{B}_{m}(\alpha)$ consisting of elements of the form
+$\delta_{B_{1}+1;m}^{\alpha_{1}-\beta_{1}}\bigg{(}f_{1}\delta_{B_{2}+1;m}^{\alpha_{2}-\beta_{2}}\big{(}\cdots f_{\ell-1}\delta_{B_{\ell}+1;m}^{\alpha_{\ell}-\beta_{\ell}}(f_{\ell})\big{)}\bigg{)}$
+where $\beta\leq\alpha,|\beta|\leq m$, $B_{i}=\beta_{1}+\ldots+\beta_{i}$ (by convention, $B_{0}=0$) and $f_{i}\in\Phi^{B_{i-1}}\Delta^{m-B_{i-1}}_{\beta_{i}}\subseteq\Delta^{m-B_{i}}_{B_{i}}$.
+Here $\delta_{a;b}$ is the operator constructed in [6] and $\Delta^{m}_{s}$ is a certain subspace of the Dickson algebra $\mathcal{D}_{s}$ ([3]) of all $GL_{s}$-invariant polynomials:
+$\Delta^{m}_{s}\subset\mathcal{D}_{s}=\mathbb{F}_{q}[Q_{s,s-1},\ldots,Q_{s,1},Q_{s,0}].$
+The Dickson invariants $Q_{s,i}$ are crucial to our investigation and will be reviewed in detail in Section 2.
+The Frobenius-like operator $\Phi$ is a ring map from $\mathcal{D}_{s}$ to $\mathcal{D}_{s+1}$ which sends $Q_{s,i}$ to $Q_{s+1,i+1}$.
+The reader of [6] may notice that in this conjecture, we describe a member of the conjectural generating set $\mathcal{B}_{m}(\alpha)$ explicitly rather than by induction.
+In addition, we make use of the core subspace $\Delta^{m}_{s}$ (see 2.2 below) of the Dickson algebra instead of $\nabla^{m}_{s}$ in the original conjecture.
+The two candidates for core spaces can be used interchangeably, but $\Delta^{m}_{s}$ is more convenient for explicit calculation, as we will see later.
+In the case $\alpha=(n)$ so that $P(\alpha)$ is the full general linear group, we have
 Conjecture 1.4.
@@
 $\delta_{s+1;m}^{n-s}(f),f\in\Delta^{m}_{s},0\leq s\leq\min(m,n),$
-forms a basis for the $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}^{G}(n)$.
-It is a simple counting problem that our proposed bases have the correct series as predicted.
+forms a basis for the $\mathbb{F}_{q}$-vector space of $G_{n}$-invariants of ${\mathcal{Q}}_{m}(n)$.
+It is a simple counting problem that our proposed bases have the correct Hilbert series as predicted.
 Hence our conjectures refine the two conjectures proposed by Lewis, Reiner, and Stanton.
-Furthermore, they also explain where the summands of the Hilbert-Poincaré series occur in the original conjectures: The summands of $C_{n,m}(t)$ are just the Hilbert-Poincaré series for $\Delta^{m}_{s}$, $0\leq s\leq\min(m,n)$, shifted by an appropriate degree because of the iterated delta operator.
+Furthermore, they also explain where the summands of the Hilbert series occur in the original conjectures: The summands of $C_{n,m}(t)$ are just the Hilbert series of $\Delta^{m}_{s}$, $0\leq s\leq\min(m,n)$, shifted by an appropriate degree due to the iterated delta operator.
 In [6], we have verified the refined parabolic conjecture 1.3 for the Borel subgroup, that is when $\alpha=1^{n}$.
-The aim of this note is to verify our conjectures for other parabolic subgroups, including the full general linear group, for $n\leq 3$.
+The aim of this paper is to verify our conjectures for other parabolic subgroups, including the full general linear group, for $n\leq 3$.
 As an immediate corollary, the parabolic conjecture of Lewis, Reiner and Stanton is true in these cases.
@@
 The refined parabolic conjectures 1.3 and 1.4 are true for $n\leq 3$ and all $m\geq 1$. Hence, the parabolic conjecture of Lewis, Reiner, and Stanton is true in these cases.
-In this paper, we choose to work as explicitly as possible to demonstrate the underlying mechanism of our construction and the additional structure of our basis.
-The computation reveals an interesting property of the invariant rings considered as a module over the Dickson algebra.
+In this paper, we choose to work as explicitly as possible to illuminate the underlying mechanism of our construction and to highlight the additional structure present in our basis.
+Our computations reveal an interesting property of the invariant rings when considered as modules over the Dickson algebra.
 We also describe the action of the Steenrod algebra.
-In subsequent work, we hope to return to this problem in greater generality.
-When $n=2$, the only parabolic subgroup other than the Borel subgroup is $\operatorname{GL}_{2}$ itself, and the parabolic conjecture has already been verified in the original paper [10].
-However, the authors worked with the dual situation for cofixed spaces and used a particular property of the bivariate case to avoid the truncated ring.
-Thus, it is unclear whether the invariant subspace ${\mathcal{Q}}_{m}^{G}$, even in rank $2$, can be described explicitly using their method.
-In [5, Proposition 4.1], Goyal constructed a family of $G$-invariant polynomials in ${\mathcal{Q}}_{m}(2)$, which he denoted as $y_{k^{\prime}}$ ($0\leq k^{\prime}\leq\frac{q^{m}-q}{q-1}$).
-It turns out that this family coincides with the $G$-invariant family $\delta_{2}(\Delta^{m}_{1})$ in our notation.
-To simplify the notation, we will often drop the index $n$ in ${\mathcal{Q}}_{m}(n)$ when the number of variables is well-understood from the context.
-Moreover, since the Frobenius power $m$ is fixed, $\delta_{s}$ will be always used in place of $\delta_{s;m}$ throughout the paper.
-We will also write $(i_{1},i_{2})\in\Delta^{m}_{2}$ instead of $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}\in\Delta^{m}_{2}$.
-Finally we denote by $[a]_{q}$ the $q$-integer $(q^{a}-1)/(q-1)$.
+In future work, we hope to revisit this problem in greater generality.
+When $n=2$, the only parabolic subgroup other than the Borel subgroup is $G_{2}$ itself, and the parabolic conjecture has already been verified in the original paper [9].
+However, the authors worked with the dual situation involving cofixed spaces and used a specific property of the bivariate case to avoid the truncated ring.
+As a result, it remains unclear whether the invariant subspace of ${\mathcal{Q}}_{m}(n)$, even in rank $n=2$, can be described explicitly using their method.
+In [5, Corollary 4.3], Goyal constructed a family of $G_{n}$-invariant polynomials in ${\mathcal{Q}}_{m}(n)$ which corresponds to the family $\delta_{2}^{n-1}(\Delta^{m}_{1})$ in our notation.
+To simplify the notation, we will often drop the index $n$ in ${\mathcal{Q}}_{m}(n),S(n)$, and $G_{n}$ when the number of variables is well-understood from the context.
+If the Frobenius power $m$ is fixed, we write $\delta_{s}$ instead of $\delta_{s;m}$.
+It is also convenient to write $(e_{1},\ldots,e_{s})\in\Delta^{m}_{s}$ to indicate a Dickson monomial $Q_{s,s-1}^{e_{1}}\ldots Q_{s,0}^{e_{s}}$ in $\Delta^{m}_{s}$.
+For any nonnegative integer $a$, we denote by $[a]_{q}$ the $q$-integer $(q^{a}-1)/(q-1)$ with the convention that $[0]_{q}=1$.
 Here is our first result in rank $2$.
 Proposition 1.6.
-The set $\mathcal{B}_{m}(2)$ consisting of 3 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}^{G}(2)$ of the full general linear group.
+The set $\mathcal{B}_{m}(2)$ consisting of 3 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}(2)^{G}$ of the full general linear group.
   1. (1)
@@
     $\Delta^{m}_{2}$.
-In rank 3, other than the Borel subgroups, there are 2 proper parabolic subgroups. Their corresponding rings of invariants are described in the next 2 propositions.
+In rank 3, other than the Borel subgroups, there are two proper parabolic subgroups. Their corresponding rings of invariants are described in the next two propositions.
 Proposition 1.7.
@@
     $\Delta^{m}_{3}$.
-Since the case of the Borel subgroup, which corresponds to the composition $(1,1,1)$ of $n=3$, has been discussed in full generality in [6], the Propositions above combined imply the main Theorem 1.5.
-Note that when $m$ is small compared to $n$, all results remain valid with proper interpretation.
-For example if $n=3$ and $m=2$, then $\Delta^{2}_{3}$ is empty, and $\delta_{3;2}(\Delta^{2}_{2})$ just consists of a single element, which is $1$ in degree 00.
-To avoid too many special cases (whose proof are not complicated), we will generally assume that $m$ is sufficiently large, comparing to $n$.
-In the course of proving these results, we have also computed the structure of these invariant rings, considered as a module over the Dickson algebra, ${\mathcal{D}}_{n}$, and a module over the Steenrod algebra $\mathcal{A}$.
-At least in the case we have computed, our basis yields a remarkable filtration, namely the subspaces $\mathcal{F}_{n,k}$, for $0\leq k\leq n$, of ${\mathcal{Q}}_{m}^{G}(n)$ spanned by $\delta_{s+1}^{n-s}(\Delta^{m}_{s})$ for all $s\leq k$.
-We have the following conjecture about this filtration:
+Since the case of the Borel subgroup, which corresponds to the composition $(1,1,1)$ of $n=3$, was treated in [6], the propositions stated above together imply the main Theorem 1.5.
+It is worth noting that when $m$ is small relative to $n$, all results still hold, provided they are interpreted appropriately.
+For example, when $n=3$ and $m=2$, the set $\Delta^{2}_{3}$ is empty, and $\delta_{3;2}(\Delta^{2}_{2})$ consists of just a single element, namely $\delta_{3;2}(1)=1\in\mathbb{F}_{q}$ in degree 00.
+To avoid excessive attention to these special cases (whose proofs are straightforward), we will generally assume that $m$ is sufficiently large compared to $n$ (in fact, $m\geq n$ is enough!).
+In the process of establishing these results, we have also computed the structure of these invariant spaces considered as modules over the Dickson algebra and the mod $q$ Steenrod algebra.
+Definition 1.10.
+For each $0\leq k\leq\min(m,n)$, let $\mathcal{F}_{n,k}$ denote the subspace of ${\mathcal{Q}}_{m}(n)^{G}$ spanned by the subsets $\delta_{s+1}^{n-s}(\Delta^{m}_{s})$ where $0\leq s\leq k$:
+$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq k\}.$
+These subspaces form an increasing filtration of ${\mathcal{Q}}_{m}(n)^{G}$, starting with the one-dimensional subspace $\mathcal{F}_{n,0}$, which is spanned by the top-degree class $(x_{1}\ldots x_{n})^{q^{m}-1}$.
+At least in the cases we have computed, this filtration is exhaustive and exhibits remarkable properties.
+We propose the following:
 Conjecture 1.11.
-For each $1\leq k\leq\min(m,n)$, the subspace $\mathcal{F}_{n,k}$ given by
-$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq k\},$
-is an $\mathcal{A}$-submodule as well as a ${\mathcal{D}}_{n}$-submodule of ${\mathcal{Q}}_{m}^{G}(n)$.
+For each $1\leq k<\min(m,n)$, $\mathcal{F}_{n,k}$ is an $\mathcal{A}$-submodule as well as a ${\mathcal{D}}_{n}$-submodule of ${\mathcal{Q}}_{m}(n)^{G}$.
 Moreover, $\mathcal{F}_{n,k}$ is annihilated by the Dickson invariants $Q_{n,0},Q_{n,1}\ldots Q_{n,n-k-1}$.
-This conjecture suggests that the filtration $\mathcal{F}_{n,*}$ may have another description of topological nature. We can prove the following:
+This conjecture suggests that the filtration $\mathcal{F}_{n,*}$ may admit an alternative description of a topological nature. We are able to verify this conjecture up to rank $3$:
 Theorem 1.12.
 The conjecture 1.11 is true for $n\leq 3$.
-The layout of this paper is as follows: Section 2 recalls basic results about modular invariant theory and the key construction from our previous work.
-In Section 3, we give a lower bound on the total dimension of the invariant subspaces.
-The following four Sections are devoted to demonstrating the four Propositions from 1.6 to 1.9 above.
-The last section discusses the action of the Dickson algebra and the Steenrod algebra and the proof of Theorem 1.11.
+The layout of this paper is as follows: Section 2 recalls basic results in modular invariant theory and the key constructions from our previous work.
+In Section 3, we establish a lower bound on the total dimension of the invariant subspaces.
+The subsequent four sections are devoted to proving Propositions 1.6 through 1.9.
+The final section discusses the actions of the Dickson algebra and the Steenrod algebra, and presents the proof of Theorem 1.12.
 Acknowledgement
-This work was partially completed while the second and third authors visited the Vietnam Institute for Advanced Study in Mathematics (VIASM).
-We want to thank Vo Thanh Tung for some computer calculations in the initial stage of this project.
+This work was partially completed while the second and third authors were visiting the Vietnam Institute for Advanced Study in Mathematics (VIASM).
+The third author gratefully acknowledges support from Hung Vuong University’s Fundamental Research Program (Project No.
+03/2024).
+We would like to thank Vo Thanh Tung for providing some computer calculations during the initial stages of this project.
+We are also grateful to the anonymous referee for their careful reading and numerous suggestions, which significantly improved the exposition.
 2. Recollection
@@
 2.1. Upper triangular invariants and Dickson invariants
-The canonical projection $S\to{\mathcal{Q}}_{m}$ is clearly a $\operatorname{GL}$-map, hence for each composition $\alpha$ of $n$, there is an induced $\mathbb{F}_{q}$-algebra homomorphism
-$S^{G}\subset S^{P(\alpha)}\to{\mathcal{Q}}_{m}^{P(\alpha)},$
-making the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}$ into a module over $S^{G}$.
-Thanks to the work of Dickson [3] since 1911, the structure of the algebra $S^{G}$, which we now call the Dickson algebra, is well-known, and plays a significant role in our investigation.
+The canonical projection $S\to{\mathcal{Q}}_{m}$ is clearly a $G$-equivariant map. Thus, for each parabolic subgroup of $G$, there is an induced $\mathbb{F}_{q}$-algebra homomorphism
+$S^{G}\subset S^{P}\to{\mathcal{Q}}_{m}^{P},$
+which makes the invariant ring ${\mathcal{Q}}_{m}^{P}$ into a module over $S^{G}$.
+The structure of the invariant ring $S^{G}$, which is the Dickson algebra [3], is well-known and plays a significant role in our investigation.
 For each positive integer $k$, denote by $V_{k}$ the product
 $V_{k}(x_{1},\ldots,x_{k})=\prod_{\lambda_{i}\in\mathbb{F}_{q}}(x_{k}+\lambda_{1}x_{1}+\ldots+\lambda_{k-1}x_{k-1}).$
-It is well-known that the invariant ring $S^{B}$ under the Borel subgroup of upper triangular matrices is a polynomial algebra generated by $V_{i}^{q-1}$, $1\leq i\leq n$.
-Let $X$ be an indeterminate and recall the fundamental equation
+It is well-known that the invariant ring $S^{B}$ under the Borel subgroup of upper triangular matrices is a polynomial algebra generated by $V_{1}^{q-1},V_{2}^{q-1},\ldots,V_{n}^{q-1}$ (see for example, [10, Theorem 3.4]).
+Let $X$ be an indeterminate, and recall the fundamental equation
 $V_{n+1}(x_{1},\ldots,x_{n},X)=X^{q^{n}}+\sum_{i=0}^{n-1}(-1)^{n-i}Q_{n,i}X^{q^{i}}.$
-The polynomials $Q_{n,i}=Q_{n,i}(x_{1},\ldots,x_{n})$ are evidently $\operatorname{GL}_{n}$-invariant.
-They are called the Dickson invariants, because of the following celebrated theorem due to Dickson:
+The polynomials $Q_{n,i}=Q_{n,i}(x_{1},\ldots,x_{n})$ are evidently $G$-invariant, and are called Dickson invariants because of the following celebrated theorem due to Dickson:
 ${\mathcal{D}}_{n}=S^{G}=\mathbb{F}_{q}[Q_{n,0},\ldots,Q_{n,n-1}].$
-The Dickson invariants $Q_{n,i}$ can be described in a more explicit form as follows. For any $n$-tuple of non-negative integers $(r_{1},\dots,r_{n})$, put
-$[r_{1},\dots,r_{n}]:=\det(x_{i}^{q^{r_{j}}})_{1\leq i,j\leq n}$
-and let $L_{n}=L_{n}(x_{1},\ldots,x_{n})$ denote $[0,1\ldots,n-1]$. We then have for $1\leq i\leq n$
-$Q_{n,i}=[0,\ldots,\hat{i},\ldots,n]/L_{n},$
-In particular, $Q_{n,n}=1$ and $Q_{n,0}=L_{n}^{q-1}$. By convention, we set $Q_{n,i}=0$ if $i<0$ or $i>n$. The following relations are well-known (see for example [13]):
-$L_{n}=V_{1}\cdots V_{n},\quad Q_{n,i}=Q_{n-1,i-1}^{q}+Q_{n-1,i}V_{n}^{q-1}.$
+The Dickson invariants $Q_{n,i}$ can be described more explicitly as follows. For any $n$-tuple of non-negative integers $(r_{1},\dots,r_{n})$, put
+$[r_{1},\dots,r_{n}]=\det(x_{i}^{q^{r_{j}}})_{1\leq i,j\leq n}.$
+In particular, let $L_{n}=L_{n}(x_{1},\ldots,x_{n})$ denote $[0,1\ldots,n-1]$. We then have
+$Q_{n,i}=[0,\ldots,\hat{i},\ldots,n]/L_{n}.$
+In particular, $Q_{n,n}=1$ and $Q_{n,0}=L_{n}^{q-1}$. By convention, we set $Q_{n,i}=0$ if $i<0$ or $i>n$. The following relations are well-known (see for example, [15, Proposition 1.3]):
+Proposition 2.1.
+Let $n$ be a positive integer and $0\leq i\leq n$. The following relations hold:
+  • (i)
+    $L_{n}=V_{1}V_{2}\cdots V_{n}$.
+  • (ii)
+    $Q_{n,i}=V_{n}^{q-1}Q_{n-1,i}+Q_{n-1,i-1}^{q}$.
 Note that $L_{n}$ is, up to sign, the product of all lines in the $\mathbb{F}_{q}$-vector space spanned by $x_{1},\ldots,x_{n}$.
 2.2. The $\Delta$ subspace of the Dickson algebra
-Given a positive integer $s$ and a partition $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{s}$ with at most $s$ nonzero parts, we say that a Dickson monomial $Q_{s,s-1}^{e_{1}}\ldots Q_{s,0}^{e_{s}}$ in ${\mathcal{D}}_{s}$ is of type $(\lambda_{1},\ldots,\lambda_{s})$ (or q-compatible with $(\lambda_{1},\ldots,\lambda_{s})$, in the language of [15, Definition 5.4]) if for each $1\leq i\leq s$,
+Lewis, Reiner and Stanton [9, Subsection 7.4] speculated that a portion of the Dickson algebra studied in [12, Section 5] might form part of an $\mathbb{F}_{q}$-basis for ${\mathcal{Q}}_{m}(n)^{G}$.
+Our work confirms this intuition.
+In fact, this portion, which we call the Delta subspace, plays a central role.
+We now review its construction and some elementary properties.
+Given a positive integer $s$ and a partition $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{s}$ with at most $s$ nonzero parts, we say that a Dickson monomial $Q_{s,s-1}^{e_{1}}\ldots Q_{s,0}^{e_{s}}$ in ${\mathcal{D}}_{s}$ is of type $(\lambda_{1},\ldots,\lambda_{s})$ (or q-compatible with $(\lambda_{1},\ldots,\lambda_{s})$, in the language of [12, Definition 5.4]) if for each $1\leq i\leq s$,
 $e_{i}\in\bigg{[}\frac{q^{\lambda_{i}}-q^{\lambda_{i+1}}}{q-1},\frac{q^{\lambda_{i}+1}-q^{\lambda_{i+1}}}{q-1}\bigg{)}.$
-It is clear that each Dickson monomial is $q$-compatible with a unique partition $(\lambda_{1},\ldots,\lambda_{s})$ of length $\leq s$.
-Now if $s\leq m$ then we restrict our attention to partitions which fit inside a rectangular $s\times(m-s)$, that is partitions with at most $s$ nonzero parts and $m-s\geq\lambda_{1}$.
-Define $\Delta_{(\lambda_{1},\ldots,\lambda_{s})}$ as the subset of the Dickson algebra ${\mathcal{D}}_{s}$ consisting of Dickson monomials of type $(\lambda_{1},\ldots,\lambda_{s})$, and let $\Delta^{m}_{s}$ denote the disjoint union of all subsets over all such partitions fitting inside $s\times(m-s)$.
-By convention, we put $\Delta^{m}_{s}=\emptyset$ if $s>m$, $\Delta^{m}_{0}=\{1\}$.
+By convention, $\lambda_{s+1}=0$. Every Dickson monomial in ${\mathcal{D}}_{s}$ is $q$-compatible with a unique partition $(\lambda_{1},\ldots,\lambda_{s})$ of length $\leq s$.
+Now suppose $s\leq m$, and we will restrict our attention to partitions whose Ferrers diagrams fit inside a $s\times(m-s)$ rectangle, that is, partitions with at most $s$ nonzero parts and $m-s\geq\lambda_{1}$.
+Define $\Delta_{(\lambda_{1},\ldots,\lambda_{s})}$ as the subset of the Dickson algebra ${\mathcal{D}}_{s}$ consisting of Dickson monomials of type $(\lambda_{1},\ldots,\lambda_{s})$, and let $\Delta^{m}_{s}$ denote the disjoint union of all subsets over all such partitions fitting inside an $s\times(m-s)$ rectangle.
+By convention, we put $\Delta^{m}_{s}=\emptyset$ if $s>m$, and $\Delta^{m}_{0}=\{1\}$.
 The Dickson monomials in $\Delta^{m}_{s}$ for $s\leq\min(m,n)$ together with the delta operators recalled below are the building blocks for our basis.
-For this reason, we will say that a Dickson monomial in $\Delta^{m}_{s}$ an essential monomial.
-We will also work with the edge monomials, which is a Dickson monomial not in $\Delta^{m}_{s}$ but has the form $Q_{s,i}f$ where $f\in\Delta^{m}_{s}$, for some $10\leq i\leq s-$ 1.
-Clearly, the edge monomials are those such that $e_{i}\in[\frac{q^{\lambda_{i}}-q^{\lambda_{i+1}}}{q-1},\frac{q^{\lambda_{i}+1}-q^{\lambda_{i+1}}}{q-1})$ for all $i\neq j$ and $e_{j}=\frac{q^{\lambda_{j}+1}-q^{\lambda_{j+1}}}{q-1}$ for some $1\leq j\leq s$.
-It is easy to check that $\Phi\Delta^{m}_{s}\subset\Delta^{m}_{s+1}$.
-In fact, $\Phi$ sends a Dickson monomial in ${\mathcal{D}}_{s}$ of type $(\lambda_{1},\ldots,\lambda_{s})$ to a Dickson monomial in ${\mathcal{D}}_{s+1}$ of type $(\lambda_{1},\ldots,\lambda_{s},0)$.
+For this reason, we will say that Dickson monomials in $\Delta^{m}_{s}$ essential monomials.
+We will also work with the edge monomials, which are the Dickson monomials not in $\Delta^{m}_{s}$ that have the form $Q_{s,i}f$ where $f\in\Delta^{m}_{s}$, for some $10\leq i\leq s-$ 1.
+Clearly, the edge monomials are those such that $e_{j}=\frac{q^{\lambda_{j}+1}-q^{\lambda_{j+1}}}{q-1}$ for some $1\leq j\leq s$ and $e_{i}\in[\frac{q^{\lambda_{i}}-q^{\lambda_{i+1}}}{q-1},\frac{q^{\lambda_{i}+1}-q^{\lambda_{i+1}}}{q-1})$ for all $i\neq j$.
 2.3. The delta operator
@@
 $\delta_{a;b}(f)=\sum\delta_{a;b}(g)h(x_{a+1},\ldots,x_{c+1}).$
-In this paper, with the number $m$ fixed from the beginning, we will always write $\delta_{s}$ for $\delta_{s;m}$ and occasionally use the full notation $\delta_{a;b}$ when necessary.
-It was shown in greater generality in [6] that $\delta_{s}(f)$ is generally not a polynomial, but it is very often a polynomial and is a $\operatorname{GL}_{s}$-invariant.
-We give a precise statement and provide an elementary proof in the case for $\delta_{3}$ to make the paper self-contained.
+It was shown in greater generality in [6] that $\delta_{s}(f)$ is generally not a polynomial.
+Still, in the cases that we are interested in, it is a genuine polynomial.
+Moreover, $\delta_{s}(f)$ is rarely a Dickson polynomial but it is a $G_{s}$-invariant modulo $I_{m}$.
+We give a precise statement and provide an elementary proof for $\delta_{3}$ to make the paper self-contained.
 Proposition 2.3.
-If $f$ is a $\operatorname{GL}_{2}$-invariant polynomial, then $\delta_{3}(f)$ is a polynomial and is $\operatorname{GL}_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$.
+If $f$ is a $G_{2}$-invariant polynomial, then $\delta_{3}(f)$ is also a polynomial and is $G_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$.
 Proof.
 We first show that $\delta_{3}(f)$ is a polynomial.
-The denominator $L_{3}$ is a product of linear forms, so it suffices to prove that the numerator becomes zero whenever there exists a nontrivial linear relation $a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}=0$.
-This is reduced to the proof that
-$a_{1}x_{1}^{q^{m}}f(x_{2},x_{3})+a_{2}x_{2}^{q^{m}}f(x_{1},x_{3})+a_{3}x_{3}^{q^{m}}f(x_{1},x_{2})=0,$
-provided $a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}=0$ and not all scalars $a_{i}$ are zero.
-Without loss of generality, assume that $a_{3}=-1$.
-If both $a_{1},a_{2}\neq 0$, then since $f$ is a $\operatorname{GL}_{2}$-invariant, we have
+Recall the denominator $L_{3}=V_{1}V_{2}V_{3}$ is a product of nonzero linear forms in the variables $x_{1},x_{2},x_{3}$, so it suffices to prove that the numerator of $\delta_{3}(f)$ becomes zero whenever there exists a nontrivial linear relation among the $x_{i}$, which by symmetry, can be assumed to have the following form:
+(2.1) $x_{3}=a_{1}x_{1}+a_{2}x_{2},\quad(a_{1},a_{2})\in\operatorname{\mathbb{F}}_{q}\times\operatorname{\mathbb{F}}_{q}.$
+We claim that this implies a similar linear relation among the columns of the numerator of $\delta_{3}(f)$, that is,
+(2.2) $x_{3}^{q^{m}}f(x_{1},x_{2})=a_{1}x_{1}^{q^{m}}f(x_{2},x_{3})+a_{2}x_{2}^{q^{m}}f(x_{1},x_{3}).$
+Indeed, this is clear if $a_{1}=a_{2}=0$. If the scalars $a_{1},a_{2}$ are both nonzero, then there is an equality
 $f(x_{2},a_{1}x_{1}+a_{2}x_{2})=f(x_{1},a_{1}x_{1}+a_{2}x_{2})=f(x_{1},x_{2}),$
-and the above expression becomes
-$[a_{1}x_{1}^{q^{m}}+a_{2}x_{2}^{q^{m}}+a_{3}x_{3}^{q^{m}}]f(x_{1},x_{2})=[a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}]^{q^{m}}f(x_{1},x_{2})=0.$
-If one of the $a_{i}$ is zero, say $a_{1}=0$, then $f(x_{2},x_{3})=f(x_{2},a_{3}x_{2})=0$ and $f(x_{1},x_{3})=f(x_{1},x_{2})$, so that the sum is reduced to
-$a_{2}x_{2}^{q^{m}}f(x_{1},x_{3})+a_{3}x_{3}^{q^{m}}f(x_{1},x_{2})=(a_{2}x_{2}+a_{3}x_{3})^{q^{m}}f(x_{1},x_{2})=0.$
-If $a_{1}=a_{2}=0$ or $x_{3}=0$, this is also clear. We have proved that $\delta_{3}(f)$ is a polynomial.
-Next we show that $\delta_{3}(f)$ is $\operatorname{GL}_{3}$-invariant modulo $I_{m}$.
-Clearly, it is symmetric, so it suffices to prove that $\delta_{3}(f)$ does not change under the operation that sends $x_{1}\mapsto x_{1}+x_{2}$ and fixes $x_{2},x_{3}$.
+since $f$ is a $G_{2}$-invariant. So the right-hand side of (2.2) becomes
+$(a_{1}x_{1}^{q^{m}}+a_{2}x_{2}^{q^{m}})f(x_{1},x_{2})=(a_{1}x_{1}+a_{2}x_{2})^{q^{m}}f(x_{1},x_{2})=x_{3}^{q^{m}}f(x_{1},x_{2}).$
+Finally, if one of the $a_{i}$ is zero, say $a_{1}=0$, Equation (2.2) is reduced to
+$(a_{2}x_{2})^{q^{m}}f(x_{1},x_{2})=a_{2}x_{2}^{q^{m}}f(x_{1},a_{2}x_{2}),$
+which is again true because $f$ is $G_{2}$-invariant. We have finished the proof that $\delta_{3}(f)$ is a polynomial.
+The next step is to show that $\delta_{3}(f)$ is $G_{3}$-invariant modulo $I_{m}$.
+By symmetry, it suffices to prove that $\delta_{3}(f)$ does not change under the operation that sends $x_{1}$ to $x_{1}+x_{2}$ and fixes $x_{2},x_{3}$.
 The denominator $L_{3}$ certainly satisfies this condition.
-By direct inspection, $\delta_{3}(f)(x_{1}+x_{2},x_{2},x_{3})-\delta_{3}(f)(x_{1},x_{2},x_{3})$ equals
+By direct inspection, the difference $\delta_{3}(f)(x_{1}+x_{2},x_{2},x_{3})-\delta_{3}(f)(x_{1},x_{2},x_{3})$ equals
 $x_{2}^{q^{m}}\times\frac{\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{q}&x_{2}^{q}&x_{3}^{q}\\ f(x_{2},x_{3})-f(x_{1}+x_{2},x_{3})&f(x_{1}+x_{2},x_{3})-f(x_{1},x_{3})&0\end{vmatrix}}{L_{3}}.$
-For this expression to be zero in ${\mathcal{Q}}_{m}(3)$, we must prove that the second factor is a polynomial. The rest of the argument is similar to the first part of the proof. ∎
+We already know that this is a polynomial.
+Moreover, $x_{2}$ occurs as a simple factor in the denominator $L_{3}$, and the determinant in the numerator vanishes when $x_{2}=0$.
+It follows that this polynomial is divisible by $x_{2}^{q^{m}}$, thus trivial in ${\mathcal{Q}}_{m}(3)$.
+∎
 Example 2.4.
-The $\operatorname{GL}_{1}$-polynomials are of the form $f(x)=x^{s(q-1)}=Q_{1,0}^{s}$, $s\geq 0$.
-Then $\delta_{2}(x^{s(q-1)})$ is a $\operatorname{GL}_{2}$-invariant polynomial in ${\mathcal{Q}}_{m}(2)$.
+The $G_{1}$-polynomials are of the form $f(x)=x^{s(q-1)}=Q_{1,0}^{s}$, $s\geq 0$.
+Then $\delta_{2}(x^{s(q-1)})$ is a $G_{2}$-invariant polynomial in ${\mathcal{Q}}_{m}(2)$.
 For $s<[m]_{q}$, this is Goyal’s $y_{s}$ family in [5] where it was defined as
 $y_{s}=x_{1}^{q^{m}-q}x_{2}^{s(q-1)}+x_{1}^{q^{m}-q-(q-1)}x_{2}^{(s+1)(q-1)}+\ldots+x_{1}^{s(q-1)}x_{2}^{q^{m}-q}.$
-Our construction is also valid for $s\geq[m]_{q}$. However, it is trivial in ${\mathcal{Q}}_{m}$ except when $s=[m]_{q}+1$, in which case
+Our construction is also valid for $s\geq[m]_{q}$. However, it is trivial in ${\mathcal{Q}}_{m}(2)$ except when $s=[m]_{q}+1$, in which case
 $\delta_{2}(Q_{1,0}^{[m]_{q}+1})=-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}=-\delta_{1}^{2}(1).$
-It is easy to check that $y_{s}$ is a genuine $\operatorname{GL}_{2}$-invariant (that is, a Dickson polynomial) whenever $s=\frac{q^{i}-1}{q-1}$ for some non-negative integer $i$.
+It is easy to check that $y_{s}$ is a genuine $G_{2}$-invariant (that is, a Dickson polynomial) whenever $s=\frac{q^{i}-1}{q-1}$ for some non-negative integer $i$.
 Example 2.5.
-For $s\leq[m]_{q}$, $\delta_{2}^{2}(Q_{1,0}^{s})$ is a $\operatorname{GL}_{3}$-invariant of degree $2(q^{m}-q)+s(q-1)$ in ${\mathcal{Q}}_{m}$.
+For $s\leq[m]_{q}$, $\delta_{2}^{2}(Q_{1,0}^{s})$ is a $G_{3}$-invariant of degree $2(q^{m}-q)+s(q-1)$ in ${\mathcal{Q}}_{m}(3)$.
 This family of elements coincides with Goyal’s polynomials $a_{m,3,s}$ in [5, Corollary 4.3] where it was written in a different form:
 $a_{m,3,s}=\sum_{(i_{1},i_{2},i_{3})}x_{1}^{i_{1}(q-1)}x_{2}^{i_{2}(q-1)}x_{3}^{i_{3}(q-1)},$
-where the sum runs over all triple $(i_{1},i_{2},i_{3})$ of non-negative integers such that $i_{s}<[m]_{q}$ and $i_{1}+i_{2}+i_{3}=2[m]_{q}-2+s$. Again, our definition works also for $s>[m]_{q}$:
-$\delta_{2}^{2}(Q_{1,0}^{[m]_{q}+2})=\delta_{1}^{3}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$
+where the sum is taken over all triples $(i_{1},i_{2},i_{3})$ of non-negative integers such that $i_{s}<[m]_{q}$ and $i_{1}+i_{2}+i_{3}=2[m]_{q}-2+s$. Again, our definition also works for $s>[m]_{q}$:
+$\delta_{2}^{2}(Q_{1,0}^{[m]_{q}+2})=\delta_{1}^{3}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1},$
 and is trivial in all other cases.
+Warning 2.6.
+It is not true in general that if $f=g$ in ${\mathcal{Q}}_{m}(n)$ then $\delta_{s}(f)=\delta_{s}(g)$ in ${\mathcal{Q}}_{m}(n)$, even in the case where all expressions involved are polynomials.
+This is evident from Example 2.4 above for $f=Q_{1,0}^{[m]_{q}+1}$ and $g=0$.
+Thus when working with iterated delta operators, for example, $\delta_{2}^{2}(f)=\delta_{2}(\delta_{2}f)$, we cannot take $\delta_{2}f$ modulo $I_{m}$ before applying $\delta_{2}$ again.
 2.4. The delta operator and the Dickson algebra
-For each composition $\alpha$ of $n$, the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is a module over the Dickson algebra ${\mathcal{D}}_{n}$.
-The next result describes how the delta operator interacts with the Dickson algebra in low rank.
+For each composition $\alpha$ of $n$, the invariant ring ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$ is a module over the Dickson algebra ${\mathcal{D}}_{n}$.
+The next result describes how the delta operator interacts with the Dickson algebra in low ranks.
 Proposition 2.7.
@@
   1. (1)
-    $Q_{s,0}\delta_{s}(f)=0$ for all $f$.
+    $Q_{s,0}\delta_{s}(f)=0$ for all $f\in S$.
   2. (2)
@@
 Proof.
-We will prove only statements (3) for $i=1$ and the last two parts, as the proof of other statements use similar arguments and are even more straightforward.
+We will make use frequently of the relation 2.1 relating upper-triangular invariants $V_{k}$ and Dickson invariants $Q_{n,i}$.
+The first part is clear since $Q_{s,0}=L_{s}^{q-1}$. Note that the numerator of $\delta_{s}(f)$ always vanishes in ${\mathcal{Q}}_{m}$ by considering the Laplace expansion along the last row.
+Proof of (2). Since $f\in\mathcal{D}_{1}$, let us assume that $f=Q_{1,0}^{s}$ for some $s\geq 0$. Since
+$Q_{2,1}=V_{2}^{q-1}(x_{1},x_{2})+V_{1}^{q(q-1)}(x_{1})=V_{2}^{q-1}(x_{2},x_{1})+V_{1}^{q(q-1)}(x_{2}),$
+the numerator of the difference $Q_{2,1}\delta_{2}(f)-\delta_{2}(Q_{1,0}^{q}f)$ has the form
+$\begin{vmatrix}x_{1}&x_{2}\\ x_{1}^{q^{m}}V_{2}(x_{2},x_{1})^{q-1}x_{2}^{s(q-1)}&x_{2}^{q^{m}}V_{2}(x_{1},x_{2})^{q-1}x_{1}^{s(q-1)}\end{vmatrix},$
+Observe that $x_{1}V_{2}(x_{1},x_{2})=x_{2}V_{2}(x_{2},x_{1})=L_{2}(x_{1},x_{2})$, it follows that
+$Q_{2,1}\delta_{2}(f)-\delta_{2}(Q_{1,0}^{q}f)=x_{2}^{q^{m}}V_{2}(x_{1},x_{2})^{q-2}x_{1}^{s(q-1)}-x_{1}^{q^{m}}V_{2}(x_{2},x_{1})^{q-2}x_{2}^{s(q-1)}=0\quad\text{in ${\mathcal{Q}}_{m}(2)$}.$
 Proof of (3).
+We will treat the case $i=1$, as the situation for $i=2$ is similar.
 Consider the difference $Q_{3,1}\delta_{3}(f)-\delta_{3}(Q_{2,0}^{q}f)$.
 Using the identity $Q_{3,1}=V_{3}^{q-1}Q_{2,1}+Q_{2,0}^{q}$, the (3,1)-entry in the last row of the determinant in the numerator can be simplified as
-$x_{1}^{q^{m}}V_{3}^{q-1}(x_{2},x_{3},x_{1})Q_{2,1}(x_{2},x_{3}).$
-Since $V_{3}(x_{2},x_{3},x_{1})L_{2}(x_{2},x_{3})=L_{3}(x_{2},x_{3},x_{1})=L_{3}(x_{1},x_{2},x_{3})$, the Laplace expansion along the last row of the determinant in the numerator shows that the resulting polynomial belongs to $I_{m}$, which is zero in ${\mathcal{Q}}_{m}$.
-Proof of (4). Write $Q_{3,2}=V_{3}^{q-1}+Q_{2,1}^{q}$. From (1) and (2) we have in ${\mathcal{Q}}_{m}$
-$Q_{3,2}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,1}^{q}\delta_{2}f).$
-Using the formula $Q_{2,1}=V_{2}^{(q-1)}+Q_{1,0}^{q}$ and unravelling the definition, we see that the difference $\delta_{2}(Q_{2,1}^{q}\delta_{2}f)-\delta_{2}^{2}(Q_{1,0}^{q^{2}}f)$ can be written as the sum
-$x_{3}^{q^{m}}\frac{x_{1}^{q^{m}}x_{2}f(x_{2})V_{2}^{q^{2}-q-1}(x_{2},x_{3})-x_{2}^{q^{m}}x_{1}f(x_{1})V_{2}^{q^{2}-q-1}(x_{1},x_{3})}{L_{2}(x_{1},x_{2})}+\\ x_{1}^{q^{m}}x_{2}^{q^{m}}f(x_{3})\frac{x_{2}V_{2}^{q^{2}-q-1}(x_{3},x_{2})-x_{1}V_{2}^{q^{2}-q-1}(x_{3},x_{1})}{L_{2}(x_{1},x_{2})}.$
-It is clear that the first summand is a polynomial, and is a multiple of $x_{3}^{q^{m}}$. The second summand belongs to the ideal $(x_{1}^{q^{m}},x_{2}^{q^{m}})$.
-Proof of (5). We have in ${\mathcal{Q}}_{m}$
-$Q_{3,1}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,0}^{q}\delta_{2}f)$
-and
-$Q_{2,0}^{q}\delta_{2}(f)=L_{2}^{q^{2}-q-1}(x_{1},x_{2})x_{1}f(x_{1})x_{2}^{q^{m}}-L_{2}^{q^{2}-q-1}(x_{1},x_{2})x_{2}f(x_{2})x_{1}^{q^{m}}.$
-The rest of the argument is similar to (4) above. ∎
+$x_{1}^{q^{m}}V_{3}^{q-1}(x_{2},x_{3},x_{1})Q_{2,1}(x_{2},x_{3})f(x_{2},x_{3}).$
+There is a similar description for other entries in this row. Since
+$V_{3}(x_{2},x_{3},x_{1})L_{2}(x_{2},x_{3})=L_{3}(x_{2},x_{3},x_{1})=L_{3}(x_{1},x_{2},x_{3}),$
+the Laplace expansion along the last row of the determinant in the numerator shows that the resulting polynomial belongs to $I_{m}$, which is zero in ${\mathcal{Q}}_{m}(3)$.
+Proof of (4). From Proposition 2.1, we have $Q_{3,2}=V_{3}^{q-1}+Q_{2,1}^{q}$. We write $Q_{3,2}$ in two different ways:
+$Q_{3,2}=V_{3}^{q-1}(x_{2},x_{3},x_{1})+Q_{2,1}^{q}(x_{2},x_{3})=V_{3}^{q-1}(x_{1},x_{3},x_{2})+Q_{2,1}^{q}(x_{1},x_{3}).$
+It follows that $Q_{3,2}\delta_{2}(\delta_{2}f)-\delta_{2}(Q_{2,1}^{q}\delta_{2}f)$ equals
+$\frac{1}{L_{2}(x_{1},x_{2})}\begin{vmatrix}x_{1}&x_{2}\\ x_{1}^{q^{m}}V_{3}^{q-1}(x_{2},x_{3},x_{1})(\delta_{2}f)(x_{2},x_{3})&x_{2}^{q^{m}}V_{3}^{q-1}(x_{1},x_{3},x_{2})(\delta_{2}f)(x_{1},x_{3})\end{vmatrix}.$
+Since $x_{2}V_{3}(x_{2},x_{3},x_{1})$ and $x_{1}V_{3}(x_{1},x_{3},x_{2})$ are divisible by $L_{2}(x_{1},x_{2})$, the above expression is a polynomial in the ideal generated by $(x_{1}^{q^{m}},x_{2}^{q^{m}})$.
+Thus
+$Q_{3,2}\delta_{2}(\delta_{2}f)=\delta_{2}(Q_{2,1}^{q}\delta_{2}f)\quad\text{in${\mathcal{Q}}_{m}(3)$}.$
+At this point, as noted in Warning 2.6, we cannot apply immediately part (2), replacing $Q_{2,1}^{q}\delta_{2}f$ with $\delta_{2}(Q_{1,0}^{q^{2}}f)$ on the right-hand side.
+Rather, we make use of the above strategy one more time, writing
+$Q_{2,1}(x_{2},x_{3})=V_{2}^{q-1}(x_{3},x_{2})+V_{1}^{q^{2}(q-1)}(x_{3}),$
+and similarly for $Q_{2,1}(x_{1},x_{3})$. The difference $\delta_{2}(Q_{2,1}^{q}\delta_{2}f)-\delta_{2}^{2}(Q_{1,0}^{q^{2}}f)$ can be written as the sum
+$x_{3}^{q^{m}}\frac{x_{2}^{q^{m}}x_{1}f(x_{1})V_{2}^{q^{2}-q-1}(x_{1},x_{3})-x_{1}^{q^{m}}x_{2}f(x_{2})V_{2}^{q^{2}-q-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}-\\ x_{2}^{q^{m}}f(x_{3})\frac{x_{1}x_{1}^{q^{m}}V_{2}^{q^{2}-q-1}(x_{3},x_{1})-x_{2}x_{1}^{q^{m}}V_{2}^{q^{2}-q-1}(x_{3},x_{2})}{L_{2}(x_{1},x_{2})}.$
+Now the first summand is a polynomial (since the numerator is divisible by $x_{1}$ and vanishes whenever there is a linear relation $x_{2}=a_{1}x_{1}$) and is a multiple of $x_{3}^{q^{m}}$.
+Similarly, the second summand belongs to the ideal $(x_{2}^{q^{m}})$.
+Proof of (5). We first apply the same strategy as part (4), writing
+$Q_{3,1}=V_{3}^{q-1}(x_{1},x_{3},x_{2})Q_{2,1}(x_{1},x_{3})+Q_{2,0}^{q}(x_{3},x_{2})=V_{3}^{q-1}(x_{2},x_{3},x_{1})Q_{2,1}(x_{2},x_{3})+Q_{2,0}^{q}(x_{3},x_{1}),$
+and obtain an equality in ${\mathcal{Q}}_{m}(3)$:
+$Q_{3,1}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,0}^{q}\delta_{2}f).$
+The right-hand side, by direct inspection, can be written as
+$x_{3}^{q^{m}}\frac{x_{2}^{q^{m}}x_{1}^{2}f(x_{1})L_{2}^{q^{2}-q-1}(x_{1},x_{3})-x_{1}^{q^{m}}x_{2}^{2}f(x_{2})L_{2}^{q^{2}-q-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}+\\ x_{2}^{q^{m}}f(x_{3})x_{3}\frac{x_{1}x_{1}^{q^{m}}L_{2}^{q^{2}-q-1}(x_{1},x_{3})-x_{2}x_{1}^{q^{m}}L_{2}^{q^{2}-q-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}.$
+By the same argument as in part (4), the first summand is divisible by $x_{3}^{q^{m}}$ and the second by $x_{2}^{q^{m}}$. ∎
 2.5. The Borel subgroup
-In [6], we have constructed a basis for the invariant ring of ${\mathcal{Q}}_{m}^{B}$ under the action of the Borel subgroup for arbirary $m$ and $n$.
-In the case $n=3$, the $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{B}$ consists 4 families:
+In [6], we constructed a basis for the invariant ring of ${\mathcal{Q}}_{m}(n)$ under the action of the Borel subgroup for arbitrary $m$ and $n$.
+In the case $n=3$, the $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{B}$ consists of $4$ families:
   1. (1)
@@
 3. A lower bound on the total dimension of the invariant subspaces
-The original parabolic conjecture predicted that the Hilbert-Poincaré series for the space of invariants ${\mathcal{Q}}_{m}^{\alpha}(n)$, where $\alpha$ is a composition of $n$, is the (finite) polynomial $C_{\alpha,m}(t)$, whose explicit form is given in 1.1.
-In particular, the value of $C_{\alpha,m}(t)$ when $t=1$ should be the total dimension of the graded $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}^{\alpha}(n)$.
+The original Parabolic Conjecture predicted that the Hilbert series of the invariant space ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$, where $\alpha$ is a composition of $n$, is given by the (finite) polynomial $C_{\alpha,m}(t)$, whose explicit form is stated in Conjecture 1.1.
+In particular, the value of $C_{\alpha,m}(t)$ when $t=1$ is expected to equal the total dimension of the finite-dimensional graded $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$.
 In this section, we will show that:
 Proposition 3.1.
-For each $m,n\geq 1$ and any composition $\alpha$ of $n$, the total dimension of the graded vector space ${\mathcal{Q}}_{m}^{\alpha}(n)$ is at least $C_{\alpha,m}(1)$.
-This result is already implicit in the discussion in Section 6 of [10], and is valid for any $n$.
-This relatively simple remark has the following important consequence in practice: suppose we can construct a generating set for the invariant subspace ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ such that, by dimension counting, the Hilbert-Poincare series $C^{\prime}_{m,\alpha}(t)$ for ${\mathcal{Q}}_{m}^{\alpha}(n)$ is not greater than $C_{\alpha,m}(t)$ ($f(t)\leq g(t)$ iff $g-f$ is a polynomial with non-negative coefficients).
-Then since there is a reverse inequality $C^{\prime}_{\alpha,m}(1)\geq C_{\alpha,m}(1)$ when evaluated at $t=1$, we can conclude that these two series are identical since both are finite degree polynomials with non-negative coefficients.
-Hence our generating set is actually a basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$.
-Proof of Proposition 3.1: Consider the following ungraded quotient of the polynomial algebra
+For each $m,n\geq 1$ and any composition $\alpha$ of $n$, the total dimension of the graded vector space ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$ is at least $C_{\alpha,m}(1)$.
+This result is already implicit in [9, Section 6] and holds for arbitrary $n\geq 1$.
+This relatively simple observation has an important practical consequence: suppose we can construct a generating set for the invariant subspace ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$ such that, by dimension counting, the associated Hilbert series $C^{\prime}_{m,\alpha}(t)$ does not exceed $C_{\alpha,m}(t)$ (i.e.
+$f(t)\leq g(t)$ if and only if $g(t)-f(t)$ is a polynomial with non-negative coefficients).
+Then, since the reverse inequality $C^{\prime}_{\alpha,m}(1)\geq C_{\alpha,m}(1)$ holds evaluated at $t=1$, we may conclude that these two series must be identical since both are finite degree polynomials with non-negative coefficients.
+It follows that our generating set is, in fact, a basis for ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$.
+Proof of Proposition 3.1: Consider the following ungraded quotient of the polynomial algebra:
 $\mathcal{R}=\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{q^{m}}-x_{1},\ldots,x_{n}^{q^{m}}-x_{n}).$
-The action of $\operatorname{GL}_{n}$ on $\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ descends to this quotient. There is a natural filtration:
+The action of $G=GL_{n}$ on $\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ descends to this quotient. There is a natural filtration:
 $\{0\}=F_{-1}\subset F_{0}\subset F_{1}\subset\cdots\subset F_{n(q^{m}-1)}={\mathcal{R}},$
@@
 $\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]\to{\mathcal{R}}\to\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}},\quad x_{i}\mapsto\bar{x}_{i}\in F_{1}/F_{0},$
-then induces an isomorphism of $\operatorname{GL}_{n}$-equivariant $\operatorname{\mathbb{F}}_{q}$-algebras:
+induces an isomorphism of $G$-equivariant $\operatorname{\mathbb{F}}_{q}$-algebras:
 ${\mathcal{Q}}_{m}(n)\cong\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}}.$
-On the other hand, if we apply the functor taking $\operatorname{GL}_{n}$-invariant on each short exact sequence of $\operatorname{GL}_{n}$-modules
+On the other hand, if we apply the functor taking $G$-invariants on each short exact sequence of $G$-modules
 $0\to F_{i-1}\to F_{i}\to F_{i}/F_{i-1}\to 0,$
-and computing the dimension, we obtain an inequality
-$\dim(F_{i})^{\operatorname{GL}_{n}}\leq\dim(F_{i-1})^{\operatorname{GL}_{n}}+\dim(F_{i}/F_{i-1})^{\operatorname{GL}_{n}}$
+and compute the dimension, we obtain an inequality
+$\dim(F_{i})^{G}\leq\dim(F_{i-1})^{G}+\dim(F_{i}/F_{i-1})^{G}$
 and so
-$\dim{\mathcal{R}}^{\operatorname{GL}_{n}}\leq\dim(\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}})^{\operatorname{GL}_{n}}\cong\dim{\mathcal{Q}}_{m}^{\operatorname{GL}_{n}}(n).$
-To compute the $\operatorname{\mathbb{F}}_{q}$-dimension of ${\mathcal{R}}^{\operatorname{GL}_{n}}$, we extend the field $\operatorname{\mathbb{F}}_{q}$ to a field $\operatorname{\mathbb{F}}$ which contains $\operatorname{\mathbb{F}}_{q^{m}}$.
+$\dim{\mathcal{R}}^{G}\leq\dim(\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}})^{G}=\dim{\mathcal{Q}}_{m}(n)^{G}.$
+To compute the $\operatorname{\mathbb{F}}_{q}$-dimension of ${\mathcal{R}}^{G}$, we extend the field $\operatorname{\mathbb{F}}_{q}$ to a field $\operatorname{\mathbb{F}}$ which contains $\operatorname{\mathbb{F}}_{q^{m}}$.
 Then the evaluation map
@@
 $\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}\xrightarrow{\cong}\operatorname{Map}\big{(}\operatorname{\mathbb{F}}_{q^{m}}^{n},\operatorname{\mathbb{F}}\big{)}.$
-This isomorphism is also $\operatorname{GL}_{n}$-equivariant where $\operatorname{GL}_{n}$ acts on the right-hand side via the embedding $\operatorname{GL}_{n}=\operatorname{GL}_{n}(\operatorname{\mathbb{F}}_{q})\subset\operatorname{GL}_{n}(\operatorname{\mathbb{F}}_{q^{m}}).$ It follows that there is an isomorphism
-$(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{\operatorname{GL}_{n}}\cong\operatorname{Map}\big{(}\operatorname{\mathbb{F}}_{q^{m}}^{n}/\operatorname{GL}_{n},\operatorname{\mathbb{F}}\big{)}.$
-The $\operatorname{\mathbb{F}}$-dimension of $(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{\operatorname{GL}_{n}}$ is thus equal to the cardinality of the orbit set $(\operatorname{\mathbb{F}}_{q^{m}})^{n}/\operatorname{GL}_{n}$.
+This isomorphism is also $G$-equivariant where $G$ acts on the right-hand side via the embedding $G=GL_{n}(\operatorname{\mathbb{F}}_{q})\subset GL_{n}(\operatorname{\mathbb{F}}_{q^{m}}).$ It follows that there is an isomorphism
+$(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{G}\cong\operatorname{Map}\big{(}\operatorname{\mathbb{F}}_{q^{m}}^{n}/G,\operatorname{\mathbb{F}}\big{)}.$
+The $\operatorname{\mathbb{F}}_{q}$-dimension of $(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{G}$ is thus equal to the cardinality of the orbit set $\operatorname{\mathbb{F}}_{q^{m}}^{n}/G$.
 Finally this cardinality is equal to the sum $\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ where $\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ is the $q$-binomial coefficient which counts the number of $s$-dimensional subspaces of the $\operatorname{\mathbb{F}}_{q}$-vector space $\operatorname{\mathbb{F}}_{q^{m}}$.
 In summary, we have shown that
-$\dim_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{\operatorname{GL}_{n}}=\dim_{F}(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{\operatorname{GL}_{n}})=\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}=C_{m,n}(1)\leq\dim{\mathcal{Q}}_{m}^{\operatorname{GL}_{n}}(n).$
-The argument above clearly is not affected if one uses a parabolic subgroup $P_{\alpha}$ instead of $\operatorname{GL}_{n}$. The proposition is proved. ∎
+$\dim_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{G}=\dim_{F}(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{G})=\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}=C_{m,n}(1)\leq\dim{\mathcal{Q}}_{m}(n)^{G}.$
+The argument above clearly is not affected if one uses a parabolic subgroup $P_{\alpha}$ instead of $G$. The proposition is proved. ∎
 We end this section with the following observation about our proposed basis $\mathcal{B}_{m}(\alpha)$.
 Lemma 3.2.
-For each composition $\alpha$ of $n$, the Hilbert-Poincaré series for the $\mathbb{F}_{q}$-space spanned by $\mathcal{B}_{m}(\alpha)$ is not greater than $C_{\alpha,m}(t)$.
+For each composition $\alpha$ of $n$, the Hilbert series for the $\mathbb{F}_{q}$-space spanned by $\mathcal{B}_{m}(\alpha)$ is not greater than $C_{\alpha,m}(t)$.
 Proof.
-Recall that
+This is almost obvious from the definition of $C_{\alpha,m}(t)$ and the relationship between the subset $\Delta^{m}_{s}\subset{\mathcal{D}}_{s}$ and the (multinomial) $(q,t)$-coefficient. Recall that
 $C_{\alpha,m}(t)=\sum_{\beta\leq\alpha,|\beta|\leq m}t^{e(m,\alpha,\beta)}\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t},$
-where
-$e(m,\alpha,\beta)=\sum(\alpha_{i}-\beta_{i})(q^{m}-q^{B_{i}})=(\alpha_{1}-\beta_{1})(q^{m}-q^{\beta_{1}})+q^{\beta_{1}}\sum_{i\geq 2}(\alpha_{i}-\beta_{i})(q^{m-\beta_{1}}-q^{\beta_{2}+\ldots+\beta_{i}}).$
-Also, from [15, Section 7], we have the following identity for the $(q,t)$-multinomial coefficient:
-$\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t}=\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\;\Biggl{(}\begin{bmatrix}m-\beta_{1}\\ \beta_{2},\ldots,\beta_{\ell},m-|\beta|\end{bmatrix}_{q,t}\biggl{)}^{q^{\beta_{1}}}.$
-Thus if we write $\beta^{\prime}=(\beta_{2},\ldots,\beta_{\ell})$ and similarly for $\alpha^{\prime}$, then $C_{\alpha,m}(t)$ can be written as
-$\sum_{\beta_{1}\leq\min(\alpha_{1},m)}t^{(\alpha_{1}-\beta_{1})(q^{m}-q^{\beta_{1}})}\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\biggl{(}\sum_{\beta^{\prime}\leq\alpha^{\prime},|\beta^{\prime}|\leq m-\beta_{1}}t^{e(m-\beta_{1},\beta^{\prime},\alpha^{\prime})}\begin{bmatrix}m-\beta_{1}\\ \beta^{\prime},m-\beta_{1}-|\beta^{\prime}|\end{bmatrix}_{q,t}\biggl{)}^{q^{\beta_{1}}}.$
-Comparing this last formula with the recursive definition of $\mathcal{B}_{m}(\alpha)$, noting that $\delta_{s+1;m}^{\alpha_{1}-s}$ raises degree by $(\alpha_{1}-s)(q^{m}-q^{s})$, the Frobenius operator multiply the degree by $q$, and the Hilbert-Poincaré series for $\Delta^{m}_{s}$ is exactly $\begin{bmatrix}m\\ s\end{bmatrix}_{q,t}$, we see that the result follow easily by induction.
-∎
+where $e(m,\alpha,\beta)=\sum(\alpha_{i}-\beta_{i})(q^{m}-q^{B_{i}})$.
+Our conjectural basis explains precisely how the summands in the above expression occur.
+Indeed, for each composition $\beta$, with $\beta\leq\alpha$ and $|\beta|\leq m$, $e(m,\alpha,\beta)$ is exactly the total degree raised by the delta operators in a basis element
+$\delta_{B_{1}+1;m}^{\alpha_{1}-\beta_{1}}\bigg{(}f_{1}\delta_{B_{2}+1;m}^{\alpha_{2}-\beta_{2}}\big{(}\cdots f_{\ell-1}\delta_{B_{\ell}+1;m}^{\alpha_{\ell}-\beta_{\ell}}(f_{\ell})\big{)}\bigg{)},$
+and the $(q,t)$-multinomial coefficient $\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t}$ corresponds to the number of choices of $\ell$-tuple $(f_{1},\ldots,f_{\ell})$ where $f_{i}\in\Phi^{B_{i-1}}\Delta_{\beta_{i}}^{m-B_{i-1}}$.
+Note that according to [12, (7.1)], we can write
+$\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t}=\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\varphi^{\beta_{1}}\begin{bmatrix}m-B_{1}\\ \beta_{2}\end{bmatrix}_{q,t}\varphi^{\beta_{2}}\begin{bmatrix}m-B_{2}\\ \beta_{3}\end{bmatrix}_{q,t}\ldots$
+where $\varphi$ is the (genuine) Frobenius operator, which reflects how our operator $\Phi$ affects the degree. ∎
 The results in this section simplify our tasks considerably.
-Indeed, in order to show that $\mathcal{B}_{m}(\alpha)$ is a basis for ${\mathcal{Q}}_{m}^{P_{\alpha}}(n)$, it suffices to verify that it is a generating set.
+Indeed, in order to show that $\mathcal{B}_{m}(\alpha)$ is a basis for ${\mathcal{Q}}_{m}(n)^{P_{\alpha}}$, it suffices to verify that it is a generating set.
 4. Proof of Proposition 1.6
-As remarked in the previous section, it suffices to show that $\mathcal{B}_{m}(2)$ is a generating set for ${\mathcal{Q}}_{m}^{G}(2)$.
-The proof of the Proposition 1.6 is done in two steps.
+As remarked in the previous section, it suffices to show that $\mathcal{B}_{m}(2)$ is a generating set for ${\mathcal{Q}}_{m}(2)^{G}$.
+The proof of Proposition 1.6 is done in two steps.
 Firstly, we use the relative transfer from the known invariant ring ${\mathcal{Q}}^{B}$ of the Borel subgroup $B$ to show that a bigger set $\mathcal{B}^{\prime}$, consisting of 3 families:
   1. (1)
-    $\delta^{2}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$,
+    $\delta_{1}^{2}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$,
   2. (2)
@@
     ${\mathcal{D}}_{2}$,
-is a generating set for ${\mathcal{Q}}_{m}^{G}(2)$.
+is a generating set for ${\mathcal{Q}}_{m}(2)^{G}$.
 Then we show that to generate the invariant ring, it is possible to replace the Dickson algebra ${\mathcal{D}}_{2}$ in the third family by the smaller subspace $\Delta^{m}_{2}$.
 Lemma 4.1.
-The set $\mathcal{B}^{\prime}$ spans ${\mathcal{Q}}_{m}^{G}(2)$.
+The set $\mathcal{B}^{\prime}$ consisting of 3 families above spans ${\mathcal{Q}}_{m}(2)^{G}$.
 Proof.
-In [6], we showed that an $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(1,1)$ for the ring of invariants under the Borel subgroup consisting of the following 2 families
+In [6], we described an $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(1,1)$ for the ring of invariants under the Borel subgroup consisting of the following 2 families:
   1. (i)
@@
     $Q_{1,0}^{i_{1}}Q_{2,1}^{i_{2}}$, $i_{1}<[m]_{q}$, $i_{2}\leq[m-1]_{q}$.
-It is a standard fact about the relative transfer $\operatorname{Tr}\colon{\mathcal{Q}}_{m}^{B}(2)\to{\mathcal{Q}}_{m}^{G}(2)$ that it is onto, so that a generating set for ${\mathcal{Q}}^{G}(2)$ can be obtained from the image of $\operatorname{Tr}$.
-By direct inspection, we have the following identity in ${\mathcal{Q}}_{m}$ for any $0\leq s<[m]_{q}$:
+It is a standard fact about the relative transfer $\operatorname{Tr}\colon{\mathcal{Q}}_{m}(2)^{B}\to{\mathcal{Q}}_{m}(2)^{G}$ that it is onto since the index of $B$ in $G$ is relatively prime to $p$ (see [14, Section 2.2]), so a generating set for ${\mathcal{Q}}_{m}(2)^{G}$ can be obtained from the image of $\operatorname{Tr}$.
+In what follows, we adopt Goyal’s notation $y_{s}$ for $\delta_{2}(Q_{1,0}^{s})$ for the sake of brevity.
+By direct inspection, we have the following identity in ${\mathcal{Q}}_{m}(2)$ for any $0\leq s<[m]_{q}$:
 $x_{1}^{q^{m}-1}x_{2}^{s(q-1)}=y_{0}x_{1}^{(s+1)(q-1)}-y_{s+1}.$
-Since $y_{s}$ is already a $\operatorname{GL}_{2}$-invariant, this formula allows us to compute the transfer of the first family:
+Since $y_{s}$ is already a $G_{2}$-invariant, this formula allows us to compute the transfer of the first family:
 $\operatorname{Tr}(x_{1}^{q^{m}-1}x_{2}^{s(q-1)})=y_{0}\operatorname{Tr}(x_{1}^{(s+1)(q-1)})-y_{s+1}.$
-Clearly, $\operatorname{Tr}(x_{1}^{(s+1)(q-1)})$ is again a Dickson polynomial in ${\mathcal{D}}_{2}$.
-By Proposition 2.5, the product of $y_{0}$ with a Dickson polynomial is either zero or another element in the $y$ family.
-Thus $\operatorname{Tr}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})$ is either $-y_{s+1}$ or zero.
-The transfer of the second family $Q_{1,0}^{i_{1}}Q_{2,1}^{i_{2}}$ is clearly a Dickson polynomial. The lemma follows. ∎
+Note that $\operatorname{Tr}(x_{1}^{(s+1)(q-1)})$ is a Dickson polynomial divisible by $Q_{2,0}$ because $x_{1}^{(s+1)(q-1)}$ is a genuine $B$-invariant which vanishes when setting $x_{1}=0$.
+By the first two parts of Proposition 2.7, the product of $y_{0}$ with $Q_{2,0}$ is trivial in ${\mathcal{Q}}_{m}(2)$.
+Thus
+$\operatorname{Tr}(x_{1}^{q^{m}-1}x_{2}^{s(q-1)})=-y_{s+1}.$
+For the second family, we argue similarly, noting that
+$\operatorname{Tr}(Q_{1,0}^{i_{1}}Q_{2,1}^{i_{2}})=\operatorname{Tr}(Q_{1,0}^{i_{1}})Q_{2,1}^{i_{2}}\in{\mathcal{D}}_{2}.$
+The lemma follows. ∎
 The next step is to show that one only needs part of the Dickson algebra.
@@
 Proposition 4.2.
-For each $10\leq i\leq m-$ 1, we have the following decomposition in ${\mathcal{Q}}_{m}(2)$:
+For each $10\leq i\leq m-$ 1, we have the following decomposition in $S$:
 $Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}=\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})+\text{essential monomials divisible by $Q_{2,0}^{\frac{q^{i+1}-1}{q-1}}$}.$
@@
 It is easy to see, by determinant manipulation, that $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})=y_{\frac{q^{i}-1}{q-1}}$ is a genuine Dickson polynomial.
-When $i=0$, the Dickson monomial $y_{0}$ must contains $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ as a montrivial summand since both are reduced to $x_{1}^{q^{m}-q}$ when setting $x_{2}=0$.
+When $i=0$, the Dickson monomial $y_{0}$ must contain $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ as a nontrivial summand since both are reduced to $x_{1}^{q^{m}-q}$ when setting $x_{2}=0$.
 Furthermore, $y_{0}-Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ is divisible by $Q_{2,0}$.
@@
 Remark 4.3.
-It is easy to check that if $i>[m-1]_{q}$, then $Q_{2,0}^{i}=L_{2}^{i(q-1)}=0$ in ${\mathcal{Q}}_{m}(2)$. Also, by induction on $m$, we have another amusing formula:
+It is easy to check that if $i>[m-1]_{q}$, then $Q_{2,0}^{i}=L_{2}^{i(q-1)\geq q^{m-1}}=0$ in ${\mathcal{Q}}_{m}(2)$. Also, by induction on $m$, we have another interesting formula:
 $Q_{2,1}^{\frac{2(q^{m-1}-1)}{q-1}}=y_{\frac{q^{m}-q}{q-1}}=x_{1}^{q^{m}-q}x_{2}^{q^{m}-q}\quad\text{in ${\mathcal{Q}}_{m}(2)$}.$
@@
 It is straightforward to check that all elements in the six families listed in Proposition 1.7 are $P(2,1)$-invariants.
-It suffices to show that $\mathcal{B}_{m}(2,1)$ is a generating set for ${\mathcal{Q}}_{m}^{P(2,1)}(3)$.
-Suppose $f$ is a nonzero polynomials in ${\mathcal{Q}}_{m}^{P(2,1)}(3)$.
-If $f$ is considered as a polynomial in $x_{3}$, then its coefficients are $\operatorname{GL}_{2}$-invariant in the $2$ variables $x_{1},x_{2}$, which are classified in Section 4.
-If one of these coefficients has the form $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$, then the corresponding term belongs to the last 2 families in the list.
-By subtracting it if necessary, let us assume that this term does not occur in $f$.
-Now suppose that the $x_{1}$ (and hence $x_{2}$ by symmetry) degree of $f$ is strictly less than $q^{m}-1$.
-Then $f$ is a linear combination of polynomials of the last two types in the list of Borel invariants in 2.5.
-In particular, the $x_{3}$-degree of $f$ must be divisible by $q^{2}-q$.
-Write $f$ as
-$f=x_{3}^{aq(q-1)}f_{aq}+x_{3}^{(aq-1)(q-1)}f_{aq-1}+\ldots,$
-where $f_{i}=f_{i}(x_{1},x_{2})\in{\mathcal{Q}}_{m}^{G}(2)$.
-If $f_{aq}$ contains $y_{t}$ for some $t\geq 0$, then we can subtract from $f$ an appropriate polynomial in the third or fourth family in the list 1.7.
-So we can assume that $f_{aq}$ belongs to $\Delta^{m}_{2}$.
-But this implies that this top degree monomial is a non-trivial summand of a $B$-invariant polynomial of the form $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$.
-In particular, this shows that $a$ is divisible by $q$.
-Write $a=bq$.
-Then by subtracting from $f$ an appropriate linear combination of first two families of $\mathcal{B}_{m}(2,1)$, we will obtain a $G(2,1)$-invariant polynomial of strictly smaller $x_{3}$-degree.
-The result follows by induction on $x_{3}$-degree.
+By results in Section 3, it suffices to show that $\mathcal{B}_{m}(2,1)$ is a generating set for ${\mathcal{Q}}_{m}(3)^{P(2,1)}$.
+Suppose $f$ is a nonzero polynomial in ${\mathcal{Q}}_{m}(3)^{P(2,1)}$, and we write $f$ as a polynomial in $x_{3}$:
+$f=x_{3}^{a(q-1)}f_{a}+x_{3}^{(a-1)(q-1)}f_{a-1}+\ldots,$
+We will prove by downward induction on the highest $x_{3}$-degree $a(q-1)$.
+Observe that the coefficients $f_{i}$ belong to the space of invariants ${\mathcal{Q}}_{m}(2)^{G}$ in the variables $x_{1}$ and $x_{2}$.
+If $f_{a}$ is a scalar multiple of $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$, then since $x_{3}^{a(q-1)}x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$ is already $P(2,1)$-invariant (which is a member of the 5th and 6th family in $\mathcal{B}_{m}(2,1)$), we can subtract it from $f$ to obtain a new one with strictly smaller $x_{3}$-degree.
+Now let us assume that $f_{a}$, and hence all $f_{i}$, for degree reasons, are generated by $\delta_{2}(\Delta^{m}_{1})$ and $\Delta^{m}_{2}$.
+In particular, its $x_{1}$-degree (and hence $x_{2}$-degree as well, by symmetry) is strictly less than $q^{m}-1$.
+We claim that $a$ must be a multiple of $q$.
+Indeed, since $f$ is a $B$-invariant, from the list of generators given in 2.5, we see that $f$ must be a linear combination of polynomials of the form $x_{1}^{j_{1}(q-1)}\delta_{2}(Q_{2,1}^{j_{2}})$ or $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$.
+In both cases, the highest $x_{3}$-degrees are divisible by $q$.
+If $f_{a}$ contains $\delta_{2}(Q_{1,0}^{i_{1}})$ for some $i_{1}<[m]_{q}$, then we can subtract from $f$ an appropriate polynomial from the third or fourth family in the list 1.7.
+So we can assume that $f_{a}\in\Delta^{m}_{2}$.
+In this case, it follows that $x_{3}^{a(q-1}f_{a}$ is a nontrivial summand of a $B$-invariant polynomial of the form $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$.
+In particular, $a$ must be divisible by $q^{2}$.
+Finally, when $a(q-1)=iq^{2}(q-1)$ for some $i<[m-2]_{q}$ or $a(q-1)=q^{m}-q^{2}$, by subtracting from $f$ an appropriate linear combination of the first two families of $\mathcal{B}_{m}(2,1)$, we will obtain a $P(2,1)$-invariant polynomial of strictly smaller $x_{3}$-degree.
+The result follows by induction on the highest $x_{3}$-degree.
 ∎
@@
 Consider these elements of $\mathcal{B}_{m}(2,1)$ as polynomials in $x_{3}$.
-The table below records the coefficient of the highest $x_{3}$-exponent.
-$\mathcal{B}_{m}(2,1)$ $x_{3}$-degree Coefficient
+The table below records the coefficient of the highest $x_{3}$-degree.
+$\mathcal{B}_{m}(2,1)$ Highest $x_{3}$-degree Coefficient
 (1) $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}Q_{3,2}^{i}$ $i(q^{3}-q^{2})<q^{m}-q^{2}$ $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$, $(i_{1},i_{2})\in\Delta^{m}_{2}$
@@
 6. Proof of Proposition 1.8
-It suffices to show that $\mathcal{B}_{m}(1,2)$ is a generating set, and the proof is by downward induction on the smallest $x_{1}$-degree. Write
+It suffices to show that $\mathcal{B}_{m}(1,2)$ is a generating set, and the proof is by downward induction on the lowest $x_{1}$-degree. Write
 $f=x_{1}^{i(q-1)}g_{i}(x_{2},x_{3})+x_{1}^{(i+1)(q-1)}g_{i+1}(x_{2},x_{3})+\ldots$
-Note that all coefficient polynomials $g_{j}$ are $\operatorname{GL}_{2}$-invariant in $x_{2}$ and $x_{3}$.
-If $i(q-1)=q^{m}-1$, then $x_{1}^{q^{m}-1}g(x_{2},x_{3})$ is a linear combination of polynomials from the last 3 families of $\mathcal{B}_{m}(1,2)$.
+Note that all coefficient polynomials $g_{j}$ are $G_{2}$-invariant in $x_{2}$ and $x_{3}$.
+If $i(q-1)=q^{m}-1$, then $x_{1}^{q^{m}-1}g(x_{2},x_{3})$ is a linear combination of polynomials from the last three families of $\mathcal{B}_{m}(1,2)$.
 We can then subtract from $f$ this polynomial and can now assume that $i(q-1)\leq q^{m}-q$.
 Since $f$ is a $B$-invariant, it is a linear combination of polynomials of the form $x_{1}^{j_{1}(q-1)}\delta_{2}(Q_{2,1}^{j_{2}})$ or $Q_{1,0}^{j_{1}}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$.
-By direct inspection, the lowest $x_{1}$-degree of these polynomials are $x_{1}^{j_{1}(q-1)}x_{2}^{q^{m}-q}x_{3}^{j_{2}q(q-1)}$ and $x_{1}^{j_{1}(q-1)}x_{2}^{j_{2}q(q-1)}Q_{2,1}^{qj_{3}}(x_{2},x_{3})$ respectively.
-In both cases, the exponent of $x_{2}$ and $x_{3}$ in the coefficient of the lowest $x_{1}$-degree is divisible by $q(q-1)$.
-We conclude that $g_{i}(x_{2},x_{3})$ is a polynomial in which all exponents of $x_{2}$ and $x_{3}$ are divisible by $q(q-1)$.
-Since $g_{i}\in{\mathcal{Q}}_{m}^{G}(x_{2},x_{3})$, it must be a $q$-th power of a polynomial in ${\mathcal{Q}}_{m-1}(2)$ in the variables $x_{2},x_{3}$.
-On the other hand, the lowest $x_{1}$-degree of the first 3 families of $\mathcal{B}_{m}(1,2)$ are
+Their lowest $x_{1}$-degree summands are $x_{1}^{j_{1}(q-1)}x_{2}^{q^{m}-q}x_{3}^{j_{2}q(q-1)}$ and $x_{1}^{j_{1}(q-1)}x_{2}^{j_{2}q(q-1)}Q_{2,1}^{qj_{3}}(x_{2},x_{3})$ respectively.
+In both cases, the exponents of $x_{2}$ and $x_{3}$ in these summands are divisible by $q(q-1)$.
+So the same is true for $g_{i}(x_{2},x_{3})$.
+Since $g_{i}\in{\mathcal{Q}}_{m}(x_{2},x_{3})^{G}$, it must be a $q$-th power of a polynomial in ${\mathcal{Q}}_{m-1}(2)$ in the variables $x_{2},x_{3}$.
+On the other hand, the lowest $x_{1}$-degree summand of the first three families of $\mathcal{B}_{m}(1,2)$ are
   1. (1)
-    $x_{1}^{j_{1}(q-1)}[Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})]^{q}$, $(i_{1},i_{2})\in\Delta^{m-1}_{2}$;
+    $x_{1}^{j_{1}(q-1)}\big{(}Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})\big{)}^{q}$, $(i_{1},i_{2})\in\Delta^{m-1}_{2}$;
   2. (2)
-    $x_{1}^{j_{1}(q-1)}[\delta_{2;m-1}(Q_{1,0}^{j_{2}})]^{q}$, $j_{2}<[m-1]_{q}$;
+    $x_{1}^{j_{1}(q-1)}\big{(}\delta_{2;m-1}(Q_{1,0}^{j_{2}})\big{)}^{q}$, $j_{2}<[m-1]_{q}$;
   3. (3)
-    $x_{1}^{j_{1}(q-1)}[x_{2}^{q^{m-1}-1}x_{3}^{q^{m-1}-1}]^{q}$.
-Hence $x_{1}^{i(q-1)}g_{i}(x_{2},x_{3})$ is the lowest $x_{1}$-degree of a polynomial in the span of $\mathcal{B}_{m}(1,2)$.
-So by subtracting from $f$ an appropriate linear combination of polynomials in $\mathcal{B}_{m}(1,2)$, we obtain another $P(1,2)$-invariant polynomial whose smallest $x_{1}$-degree is strictly greater than that of $f$.
+    $x_{1}^{j_{1}(q-1)}\big{(}x_{2}^{q^{m-1}-1}x_{3}^{q^{m-1}-1}\big{)}^{q}$.
+Hence $x_{1}^{i(q-1)}g_{i}(x_{2},x_{3})$ is the lowest $x_{1}$-degree part of a polynomial in the span of $\mathcal{B}_{m}(1,2)$.
+By subtracting from $f$ an appropriate linear combination of polynomials in $\mathcal{B}_{m}(1,2)$, we obtain another $P(1,2)$-invariant polynomial whose lowest $x_{1}$-degree is strictly greater than that of $f$.
 By induction, the required result follows.
@@
 Remark 6.1.
-We can also prove directly $\mathcal{B}_{m}(1,2)$ is linearly independent.
+We can also prove directly that $\mathcal{B}_{m}(1,2)$ is linearly independent.
 In fact, the generators can be distinguished by looking at the lowest $x_{1}$-degrees and the corresponding coefficients as given in the table below.
-$\mathcal{B}_{m}(1,2)$ $x_{1}$-degree Coefficient
+$\mathcal{B}_{m}(1,2)$ Lowest $x_{1}$-degree Coefficient
 (1) $Q_{1,0}^{j_{1}}Q_{3,2}^{i_{1}}Q_{3,1}^{i_{2}}$ $j_{1}(q-1)<q^{m}-1$ $Q_{2,1}^{qi_{1}}(x_{2},x_{3})Q_{2,0}^{qi_{2}}(x_{2},x_{3})$, $(i_{1},i_{2})\in\Delta^{m-1}_{2}$
@@
 7. Proof of Proposition 1.9
-We need to prove that $\mathcal{B}_{m}(3)$ is a generating set for ${\mathcal{Q}}_{m}^{G}(3)$.
-The proof’s strategy is the same as in the rank $2$ case, but the details are much more involved.
+We need to prove that $\mathcal{B}_{m}(3)$ is a generating set for ${\mathcal{Q}}_{m}(3)^{G}$.
+Our strategy is the same as in the rank $2$ case, but the details are much more involved.
 The proof is divided into three steps.
@@
 This is proved by first showing that all the edge monomials of $\Delta^{m}_{3}$ are in this span.
-Thus the subspace spanned by $\mathcal{B}_{m}(3)$ is a ${\mathcal{D}}_{3}$-submodule of ${\mathcal{Q}}_{m}^{G}(3)$.
-Since it contains $\Delta^{m}_{3}$, it will contain the entire ${\mathcal{D}}_{3}$.
+Thus the subspace spanned by $\mathcal{B}_{m}(3)$ is a ${\mathcal{D}}_{3}$-submodule of ${\mathcal{Q}}_{m}(3)^{G}$.
+Since it contains $\Delta^{m}_{3}$, it will contain the entirety of ${\mathcal{D}}_{3}$.
 Lemma 7.1.
-The set $\mathcal{B}^{\prime}$ consisting of 4 families below is a generating set for ${\mathcal{Q}}_{m}^{G}(3)$.
+The set $\mathcal{B}^{\prime}$ consisting of 4 families below is a generating set for ${\mathcal{Q}}_{m}(3)^{G}$.
   1. (1)
@@
 Proof.
-We use the relative transfer from $P(1,2)$ to $G$.
-Among the six families in $\mathcal{B}_{m}(1,2)$, the last family is already a $G$-invariant.
-So are $\delta_{3}(Q_{2,1}^{j_{2}})$ and $\delta_{2}^{2}(1)$.
-On the other hand, $\operatorname{Tr}(x_{1}^{j_{1}(q-1)})$ is clearly an element of ${\mathcal{D}}_{3}$ and is a multiple of $Q_{3,0}$.
-It is known that
-$Q_{3,0}\delta_{3}(Q_{2,1}^{j_{2}})=Q_{3,0}\delta_{2}^{2}(1)=0\quad\text{in ${\mathcal{Q}}_{m}$}.$
-It follows that the transfer of the second and third family in $\mathcal{B}_{m}(1,2)$ contribute $G$-invariants of the form $\delta_{3}(Q_{2,1}^{j_{2}})$, $j_{2}<[m-1]_{q}$ and $\delta_{2}^{2}(1)$.
-The transfer of the 4th family clearly belongs to ${\mathcal{D}}_{3}$, so it remains to consider the 5th family.
-We claim that if $j<[m]_{q}$, then
-$\operatorname{Tr}_{P(1,2)}^{G}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=a_{m,3,j+1}.$
-Indeed, consider $x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3})$ as a $B_{3}$-invariant and taking the transfer from $B_{3}$ to $P(2,1)$. In Section 4, we have already proved that
-$\operatorname{Tr}_{B_{2}}^{\operatorname{GL}_{2}}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})=y_{a+1}(x_{1},x_{2}).$
-Hence
-$\operatorname{Tr}_{B_{3}}^{P(2,1)}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=\sum_{a=j}^{(q^{m}-q)/(q-1)}y_{a+1}(x_{1},x_{2})x_{3}^{q^{m}-q+j(a-1)-a(q-1)}.$
-The last sum is exactly $a_{m,3,j+1}=\delta_{2}^{2}(Q_{1,0}^{j+1})$, which is already a $\operatorname{GL}_{3}$-invariant.
-By the transitivity property of the transfer, we have the required claim.
-The proof is finished.
-∎
-The next step is to show that in the third family, one need not use the entire ${\mathcal{D}}_{2}$. The following technical lemma will be used repeatedly.
+A generating set for ${\mathcal{Q}}_{m}(3)^{G}$ can be obtained by taking the image of the relative transfer
+$\operatorname{Tr}_{P(2,1)}^{G}\colon{\mathcal{Q}}_{m}^{P(1,2)}\longrightarrow{\mathcal{Q}}_{m}^{G}$
+from the parabolic subgroup $P(1,2)$ to $G$, applied to the basis of ${\mathcal{Q}}_{m}^{P(1,2)}$ given in Proposition 1.8.
+First, observe that $Q_{1,0}^{j_{1}}Q_{3,2}^{i_{1}}Q_{3,1}^{i_{2}}$ is a genuine $P(1,2)$-invariant, and hence its transfer is a genuine $G$-invariant, i.e., a Dickson polynomial.
+Next, consider the second family. Since $\delta_{3}(Q_{2,1}^{j_{2}})$ is already a $G$-invariant in ${\mathcal{Q}}_{m}(3)$, we have
+$\operatorname{Tr}_{P(2,1)}^{G}(Q_{1,0}^{j_{1}}\delta_{3}(Q_{2,1}^{j_{2}}))=(\operatorname{Tr}_{P(2,1)}^{G}Q_{1,0}^{j_{1}})\delta_{3}(Q_{2,1}^{j_{2}}).$
+On the other hand, the transfer $\operatorname{Tr}_{P(2,1)}^{G}(Q_{1,0}^{j_{1}})$ is a genuine Dickson polynomial divisible by $Q_{3,0}$ unless $j_{1}=0$.
+Therefore, if $j_{1}$ is positive, the whole expression vanishes in ${\mathcal{Q}}_{m}(3)$:
+$\operatorname{Tr}_{P(2,1)}^{G}(Q_{1,0}^{j_{1}}\delta_{3}(Q_{2,1}^{j_{2}}))=0\quad\text{in ${\mathcal{Q}}_{m}$}.$
+A similar argument applies to the third family.
+Since $\delta_{2}^{2}(1)$ is already a $G$-invariant and is annihilated by $Q_{3,0}$, its transfer also vanishes.
+Hence from the second and third families in $\mathcal{B}_{m}(1,2)$, only elements $\delta_{3}(Q_{2,1}^{j_{2}})$ (for $j_{2}<[m-1]_{q}$) and $\delta_{2}^{2}(1)$ contribute nontrivially.
+Now we turn to the fourth family. Recall that
+$\delta_{1}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})=x_{1}^{q^{m}-1}Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3}).$
+We claim that
+$\operatorname{Tr}_{P(1,2)}^{G}(\delta_{1}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}))=\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}+1}).$
+Indeed, by Proposition 2.3, the right-hand side is $G$-invariant. Moreover, we use the identity:
+$V_{3}(x_{2},x_{1},x_{3})=V_{2}(V_{2}(x_{2},x_{1}),V_{2}(x_{2},x_{3}))=V_{2}(x_{2},x_{3})^{q}-V_{2}^{q-1}(x_{2},x_{1})V_{2}(x_{2},x_{3}),$
+from which it follows that
+$L_{2}^{q}(x_{2},x_{3})=x_{2}^{q}V_{2}(x_{2},x_{3})^{q}=x_{2}^{q}V_{3}(x_{2},x_{1},x_{3})+Q_{2,0}(x_{1},x_{2})L_{2}(x_{2},x_{3}).$
+Applying the Laplace expansion along the last row of the numerator, we obtain:
+$\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}+1})=Q_{2,0}\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})-\delta_{2}(Q_{1,0}Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}).$
+Now consider $x_{1}^{q^{m}-1}Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})$ as a $B_{3}$-invariant and take the transfer from $B_{3}$ to $P(2,1)$. From Section 4, we have
+$\operatorname{Tr}_{B_{2}}^{G_{2}}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})=-y_{a+1}(x_{1},x_{2})=-\delta_{2}(Q_{1,0}^{a+1}).$
+Write
+$Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})=\sum_{(a,b)}\lambda(a,b)x_{2}^{a(q-1)}x_{3}^{b(q-1)},\quad\lambda(a,b)\in\mathbb{F}_{q},$
+then
+$\operatorname{Tr}_{B_{3}}^{P(2,1)}\big{(}x_{1}^{q^{m}-1}Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})\big{)}=$ $-\sum_{(a,b)}\lambda(a,b)\delta_{2}(Q_{1,0}^{a+1})x_{3}^{b(q-1)}$
+$=$ $-\delta_{2}\big{(}\sum_{(a,b)}\lambda(a,b)x_{1}^{(a+1)(q-1)}x_{2}^{b(q-1)}\big{)}$
+$=$ $-\delta_{2}(Q_{1,0}Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$
+$=$ $\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}+1})-Q_{2,0}\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}).$
+Finally, we observe that
+$\operatorname{Tr}_{P(2,1)}^{G}\big{(}Q_{2,0}\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})\big{)}=\operatorname{Tr}_{P(2,1)}^{G}(Q_{2,0})\delta_{3}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})=0,$
+which proves the claim for the fourth family.
+Now we consider the fifth family (note that the 6th family already consists of $G$-invariants). A similar argument shows that if $j<[m]_{q}$, then
+$\operatorname{Tr}_{P(1,2)}^{G}(x_{1}^{q^{m}-1}\delta_{2}(Q_{1,0}^{j})(x_{2},x_{3}))=-\delta_{2}^{2}(Q_{1,0}^{j+1}).$
+Indeed, recall
+$\operatorname{Tr}_{B_{3}}^{P(2,1)}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=-\sum_{a=j}^{(q^{m}-q)/(q-1)}y_{a+1}(x_{1},x_{2})x_{3}^{q^{m}-q+j(a-1)-a(q-1)}.$
+This sum is exactly $-a_{m,3,j+1}=-\delta_{2}^{2}(Q_{1,0}^{j+1})$, which is a $G_{3}$-invariant. The result follows by the transitivity of the transfer. The proof is complete. ∎
+The next step is to show that in the third family, one does not need to use all ${\mathcal{D}}_{2}$. The following technical lemma will be used repeatedly.
 Lemma 7.2.
-Let $s,t$ be non-negative integers and $0\leq i<[m]_{q}$. Then the following equality holds in ${\mathcal{Q}}_{m}(3)$:
+Let $s,t,i$ be non-negative integers. Then the following equality holds in ${\mathcal{Q}}_{m}(3)$:
 $\delta_{2}\big{(}V_{1}^{s(q-1)}V_{2}^{t(q-1)}\delta_{2}(Q_{1,0}^{i})\big{)}=\begin{cases}0&\text{if $t\geq 1$ and $s>1$};\\ x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-1)(q-1)}&\text{if $t\geq 1$ and $s=1$};\\ (t+1)x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-2)(q-1)}&\text{if $t\geq 1$ and $s=0$};\\ \delta_{2}^{2}(Q_{1,0}^{i+s})&\text{if $t=0,s\leq 1$};\\ \delta_{2}^{2}(Q_{1,0}^{i+s})-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(s+i-2)(q-1)}&\text{if $t=0,s\geq 2$}.\end{cases}$
 Proof.
-Recall that $\delta_{s+1}(f)$ is a polynomial if $f$ is a polynomial (possibly with higher than $s$ variables), which is $\operatorname{GL}_{s}$-invariant.
-Hence the expression under consideration is a polynomial.
-If $t\geq 1$, then by definition, $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{t(q-1)}\delta_{2;m}(Q_{1,0}^{i}))$ equals
+Recall that $\delta_{2}(f)$ is a polynomial if $f$ is a polynomial (possibly with more than one variable), which is $G_{1}$-invariant.
+Hence, the expression under consideration is a polynomial.
+If $t\geq 1$, unravelling the definition, we see that $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{t(q-1)}\delta_{2}(Q_{1,0}^{i}))$ equals
 $\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{2}^{q^{m}}x_{1}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{1},x_{3})\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{1}^{i(q-1)}\end{vmatrix}-\\ x_{1}^{q^{m}}x_{2}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{2},x_{3})\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{2}^{i(q-1)}\end{vmatrix}\bigg{]}\\ =x_{3}^{q^{m}}\;\frac{x_{2}^{q^{m}}x_{1}^{(s+i)(q-1)+1}V_{2}^{t(q-1)-1}(x_{1},x_{3})-x_{1}^{q^{m}}x_{2}^{(s+i)(q-1)+1}V_{2}^{t(q-1)-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}-\\ x_{1}^{q^{m}}x_{2}^{q^{m}}x_{3}^{i(q-1)+1}\;\frac{x_{1}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{1},x_{3})-x_{2}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}.$
-The first summand is clearly trivial in ${\mathcal{Q}}_{m}$ because the quotient involved is a polynomial.
-For the second summand, it is also trivial if $s>1$, and is $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-1)(q-1)}$ if $s=1$.
-If $s=0$, the result is
-$-(t(q-1)-1)x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-2)(q-1)}.$
-Now suppose $t=0$. Then we have
-$\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{2}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{1}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\big{]}\\ =\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{2}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{1}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\big{]}+\\ \frac{x_{1}^{q^{m}}x_{2}^{q^{m}}x_{3}^{i(q-1)}}{L_{2}(x_{1},x_{2})}\Bigg{[}\frac{x_{2}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{2}^{q-1}-x_{3}^{q-1}}-\frac{x_{1}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{1}^{q-1}-x_{3}^{q-1}}\Bigg{]}$
-The first summand is $\delta_{2}^{2}(Q_{1,0}^{s+i})$.
-In ${\mathcal{Q}}_{m}(3)$, the second summand is trivial when $s\leq 1$ and equals $-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(s-2+i)(q-1)}$ if $s\geq 2$.
-∎
-We are now ready to show that
+The first summand is trivial in ${\mathcal{Q}}_{m}(3)$ because the quotient involved is a polynomial.
+Indeed, in this quotient, the numerator is divisible by $x_{1}$ and vanishes whenever there is a linear relation $x_{2}=a_{1}x_{1}$.
+By a similar argument, the second summand is also trivial if $s>1$, and equals $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-1)(q-1)}$ if $s=1$. If $s=0$, it equals
+$(t(q-1)-1)x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-2)(q-1)}.$
+Now suppose $t=0$, we get
+$\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{2}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{1}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\big{]}\\ =\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{2}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{1}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\big{]}+\\ \frac{x_{1}^{q^{m}}x_{2}^{q^{m}}x_{3}^{i(q-1)}}{L_{2}(x_{1},x_{2})}\Bigg{[}\frac{x_{2}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{2}^{q-1}-x_{3}^{q-1}}-\frac{x_{1}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{1}^{q-1}-x_{3}^{q-1}}\Bigg{]}.$
+The first summand is $\delta_{2}^{2}(Q_{1,0}^{s+i})$. The second summand is trivial when $s\leq 1$ and equals $-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(s-2+i)(q-1)}$ if $s\geq 2$. ∎
+Corollary 7.3.
+If $g\in{\mathcal{D}}_{2}$ and $i\geq 0$, then $\delta_{3}(g\delta_{2}(Q_{1,0}^{i}))$ is a linear combination of polynomials from the following families:
+  • $\delta_{2}^{2}(Q_{1,0}^{s})$, $0\leq s<[m]_{q}$.
+  • $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{s(q-1)}$, $0\leq s\leq[m]_{q}$.
+Proof.
+We first observe that by direct inspection, $\delta_{3}(g\delta_{2}(Q_{1,0}^{i}))$ can be written as a quotient of determinants:
+$\delta_{3}(g\delta_{2}(Q_{1,0}^{i}))=\frac{\begin{vmatrix}x_{1}^{i(q-1)+1}&x_{2}^{i(q-1)+1}&x_{3}^{i(q-1)+1}\\ x_{1}^{q^{m}}&x_{2}^{q^{m}}&x_{3}^{q^{m}}\\ x_{1}^{q^{m}}g(x_{2},x_{3})&x_{2}^{q^{m}}g(x_{1},x_{3})&x_{3}^{q^{m}}g(x_{1},x_{2})\end{vmatrix}}{L_{3}(x_{1},x_{2},x_{3})}.$
+Since
+$L_{3}(x_{1},x_{2},x_{3})=$ $V_{3}(x_{1},x_{2},x_{3})L_{2}(x_{1},x_{2})$
+$=$ $[V_{2}(x_{2},x_{3})^{q-1}-V_{2}(x_{2},x_{1})^{q-1}]V_{2}(x_{2},x_{3})L_{2}(x_{1},x_{2})$
+$=$ $[V_{2}(x_{1},x_{3})^{q-1}-V_{2}(x_{1},x_{2})^{q-1}]V_{2}(x_{1},x_{3})L_{2}(x_{1},x_{2}),$
+the quotient above can be rewritten in the form:
+$\varphi(h):=$ $\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}h(x_{1},x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{i(q-1)+1}&x_{3}^{i(q-1)+1}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}h(x_{2},x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{i(q-1)+1}&x_{3}^{i(q-1)+1}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\bigg{]}$
+$=$ $\frac{1}{L_{2}(x_{1},x_{2})}(x_{1}^{q^{m}}x_{2}h(x_{1},x_{2},x_{3})\delta_{2}(Q_{1,0}^{i})(x_{2},x_{3})-x_{2}^{q^{m}}x_{1}h(x_{2},x_{1},x_{3})\delta_{2}(Q_{1,0}^{i})(x_{1},x_{3})),$
+where
+$h(x_{1},x_{2},x_{3})=\frac{g(x_{2},x_{3})-g(x_{2},x_{1})}{V_{2}(x_{2},x_{3})^{q-1}-V_{2}(x_{2},x_{1})^{q-1}}.$
+Since $g\in{\mathcal{D}}_{2}$, it can be written in terms of upper triangular invariants
+$g=\sum_{s,t}V_{1}^{s(q-1)}V_{2}^{t(q-1)},$
+where the sum is taken over certain subset $A\subset\mathbb{N}^{2}$. Then
+$h(x_{1},x_{2},x_{3})=\sum_{(s,t)\in A,t\geq 1}x_{2}^{s(q-1)}\bigg{(}\sum_{j=0}^{t-1}V_{2}(x_{2},x_{3})^{(t-1-j)(q-1)}V_{2}(x_{2},x_{1})^{j(q-1)}\bigg{)}.$
+Observe that if $j>0$, then both $x_{1}^{q^{m}}x_{2}V_{2}(x_{2},x_{1})^{j(q-1)}/L_{2}$ and $x_{2}^{q^{m}}x_{1}V_{2}(x_{1},x_{2})^{j(q-1)}/L_{2}$ are zero in ${\mathcal{Q}}_{m}$, so the terms in $h$ corresponding to those $j$ do not contribute to $\varphi(h)$.
+It follows that $\varphi(h)$ is reduced to the sum
+$\sum_{(s,t)\in A,t\geq 1}\varphi(x_{2}^{s(q-1)}V_{2}(x_{2},x_{3})^{(t-1)(q-1)}).$
+On the other hand, we have an equality which is essentially by definition:
+$\varphi(x_{2}^{s(q-1)}V_{2}(x_{2},x_{3})^{(t-1)(q-1)})=\delta_{2}(V_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}\delta_{2}(Q_{1,0}^{i})).$
+We can now use the computation in Lemma 7.2 of the righ-hand side to finish the proof. ∎
+We are now ready to replace ${\mathcal{D}}_{2}$ by $\Delta^{m}_{2}$ in the generating set $\mathcal{B}^{\prime}$ of ${\mathcal{Q}}_{m}(3)^{G}$.
 Lemma 7.4.
-The subspace of ${\mathcal{Q}}_{m}^{G}(3)$ spanned by the following three families of elements
+The subspace of ${\mathcal{Q}}_{m}(3)^{G}$ spanned by the following three families of elements
   1. (1)
@@
 for some $10\leq i\leq m-$ 1, $0\leq a,b$. Recall from 4.2 that
-$Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}=\delta_{2}(x^{q^{i}-1})+\text{essential monomials divisible by $Q_{2,0}^{\frac{q^{i+1}-1}{q-1}}$}.$
-By repeating this process if necessary, we can write $f$ as a sum of essential monomials and Dickson polynomials of the form
-$\delta_{2}(x^{q^{j}-1})Q_{2,1}^{a_{j}}Q_{2,0}^{b_{j}},\quad j\geq i.$
-Thus it suffices to prove the lemma for $f=g\delta_{2}(x^{q^{i}-1})$, where $g\in{\mathcal{D}}_{2}$ and $10\leq i\leq m-$ 1.
-To evaluate $\delta_{3}(g\delta_{2}(x^{q^{i}-1}))$, we first observe that by direct inspection, it can be simplified as a determinant:
-$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))=\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{2}^{q^{m}}&x_{3}^{q^{m}}\\ x_{1}^{q^{m}}g(x_{2},x_{3})&x_{2}^{q^{m}}g(x_{1},x_{3})&x_{3}^{q^{m}}g(x_{1},x_{2})\end{vmatrix}}{L_{3}(x_{1},x_{2},x_{3})}.$
-Since
-$L_{3}(x_{1},x_{2},x_{3})=$ $V_{3}(x_{1},x_{2},x_{3})L_{2}(x_{1},x_{2})$
-$=$ $[V_{2}(x_{2},x_{3})^{q-1}-V_{2}(x_{2},x_{1})^{q-1}]V_{2}(x_{2},x_{3})L_{2}(x_{1},x_{2})$
-$=$ $[V_{2}(x_{1},x_{3})^{q-1}-V_{2}(x_{1},x_{2})^{q-1}]V_{2}(x_{1},x_{3})L_{2}(x_{1},x_{2}),$
-this determinant can be rewritten in the form:
-$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}h(x_{1},x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}h(x_{2},x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\bigg{]}.$
-where
-$h(x_{1},x_{2},x_{3})=\frac{g(x_{2},x_{3})-g(x_{2},x_{1})}{V_{2}(x_{2},x_{3})^{q-1}-V_{2}(x_{2},x_{1})^{q-1}}.$
-This quotient is a symmetric polynomial in $V_{2}^{q-1}(x_{2},x_{3})$ and $V_{2}^{q-1}(x_{2},x_{1})$ with coefficient in $\mathbb{F}_{q}[x_{2}^{q-1}]$.
-Moreover, if one summand of this quotient is divisible by $V_{2}(x_{2},x_{1})^{q-1}$, then the common denominator $L_{2}(x_{1},x_{2})$ will be canceled out, and the result is zero in ${\mathcal{Q}}_{m}$.
-Thus if we write $g$ in terms of the upper triangular invariants, $g=\sum_{s,t}V_{1}^{s(q-1)}V_{2}^{t(q-1)}$, then the determinant above equals in ${\mathcal{Q}}_{m}$ to a sum over all such pair $(s,t)$, $t\geq 1$ of
-$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}V_{2}^{(t-1)(q-1)}(x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}(x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\bigg{]},$
-which is $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$. We can now use the computation in Lemma 7.2 to see that
-$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))\in\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{2}(Q_{1,0}^{a}),a\geq 0)\}.$
-On the other hand, $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})$ is a genuine Dickon polynomial, hence $\delta_{3}(g\delta_{2}(x^{q^{i}-1}))$ is a $\operatorname{GL}_{3}$-invariant in ${\mathcal{Q}}_{m}$.
-This shows that
-$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))\in\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{3}(1)\}.$
+$Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}=\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})+\text{essential monomials divisible by $Q_{2,0}^{\frac{q^{i+1}-1}{q-1}}$}.$
+Repeating this process if necessary, we can write $f$ as a sum of essential monomials and Dickson polynomials of the form
+$\delta_{2}(Q_{1,0}^{\frac{q^{j}-1}{q-1}})Q_{2,1}^{a_{j}}Q_{2,0}^{b_{j}},\quad j\geq i.$
+Thus it suffices to consider $\delta_{3}(f)$ for $f=g\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})$, where $g\in{\mathcal{D}}_{2}$ and $10\leq i\leq m-$ 1.
+We know from Corollary 7.3 that $\delta_{3}(g\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$ is a linear combination of polynomials in the following families:
+  • $\delta_{2}^{2}(\Delta^{m}_{1})$,
+  • $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{s(q-1)}$, $0\leq s\leq[m]_{q}$.
+Since $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})$ is a genuine Dickson polynomial, $\delta_{3}(g\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$ is a $G_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$.
+We conclude that $\delta_{3}(g\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$ must belong to the space
+$\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{3}(1)\},$
+because any monomial of the form $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{i(q-1)}$, $0\leq i<[m]_{q}$, is not $G_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$. ∎
 The third, and final step is to replace the full Dickson algebra ${\mathcal{D}}_{3}$ by $\Delta^{m}_{3}$.
 We first need some preliminary calculations.
-In the following, we write $\delta_{3;a}(i_{1},i_{2})$ for $\delta_{3;a}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$ to simplify the notation.
-Again $\delta_{3}$ will stand for $\delta_{3;m}$.
+In the following, $\delta_{3}$ will again stand for $\delta_{3;m}$.
 Lemma 7.5.
-The following equalities hold:
+The following equalities hold in $S$:
   1. (1)
-    $\delta_{3}(0,\frac{q^{i}-1}{q-1})=\delta_{3;m-i}^{q^{i}}(1)Q_{3,0}^{\frac{q^{i}-1}{q-1}}$ for all $i\geq 0$. In particular,
-    $\delta_{3}(0,{\frac{q^{m-2}-1}{q-1}})=Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}\text{ and }\delta_{3}(0,\frac{q^{m-3}-1}{q-1})=Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}.$
+    $\delta_{3}(Q_{2,0}^{\frac{q^{i}-1}{q-1}})=\big{(}\delta_{3;m-i}(1)\big{)}^{q^{i}}Q_{3,0}^{\frac{q^{i}-1}{q-1}}$ for all $i\geq 0$. In particular,
+    $\delta_{3}(Q_{2,0}^{\frac{q^{m-2}-1}{q-1}})=Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}\text{\quad and \quad}\delta_{3}(Q_{2,0}^{\frac{q^{m-3}-1}{q-1}})=Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}.$
   2. (2)
@@
   3. (3)
-    $\delta_{3}(q^{m-3},\frac{q^{m-3}-1}{q-1})=Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$.
+    $\delta_{3}(Q_{2,1}^{q^{m-3}}Q_{2,0}^{\frac{q^{m-3}-1}{q-1}})=Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$.
   4. (4)
     In general, for $\lambda_{2}\geq\lambda_{3}\geq 0$, then
-    $\delta_{3}\bigg{(}\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1}\bigg{)}=\delta_{3;m-\lambda_{3}}^{q^{\lambda_{3}}}\bigg{(}\frac{q^{\lambda_{2}-\lambda_{3}}-1}{q-1},0\bigg{)}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}.$
+    $\delta_{3}\big{(}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{2,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}\big{)}=\bigg{(}\delta_{3;m-\lambda_{3}}\big{(}Q_{2,1}^{\frac{q^{\lambda_{2}-\lambda_{3}}-1}{q-1}}\big{)}\bigg{)}^{q^{\lambda_{3}}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}.$
 Proof.
-The proof of these statements are quite similar, and essentially makes use of the Laplace expansion.
+The proofs of these statements are quite similar and essentially make use of the Laplace expansion.
 We will prove the last statement only.
-Apply the Laplace expansion along the last row of the determinant in the numerator for $\delta_{3}(\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1})$, we obtain
+Apply the Laplace expansion along the last row of the determinant in the numerator of $\delta_{3}\big{(}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{2,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}\big{)}$, we obtain
 $x_{1}^{q^{m}}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}(x_{2},x_{3})L_{2}^{q^{\lambda_{3}}}(x_{2},x_{3})-x_{2}^{q^{m}}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}(x_{1},x_{3})L_{2}^{q^{\lambda_{3}}}(x_{1},x_{3})+x_{3}^{q^{m}}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}(x_{1},x_{2})L_{2}^{q^{\lambda_{3}}}(x_{1},x_{3}).$
-The result follows easily after writing this expression as a $q^{\lambda_{3}}$ power. ∎
-Our second technical result plays a similar role as Proposition 4.2 in the case of 3 variables.
+After writing this expression as a $q^{\lambda_{3}}$ power, the result follows easily. ∎
+Our second technical result is similar to Proposition 4.2 in the case of 3 variables.
 Proposition 7.6.
-Suppose $20\leq\lambda\leq m-$ 2, then in ${\mathcal{Q}}_{m}$, $\delta_{3}(\frac{q^{\lambda}-1}{q-1},0)$ equals a Dickson polynomial in which the only non-essential summand which is not divisible by $Q_{3,0}$ is $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,1}^{\frac{q^{\lambda}-1}{q-1}}$.
+Let $m\geq 2$ be an integer.
+  1. (i)
+    The following equality holds in $S$:
+    $\frac{[0,1,m-1]}{[0,1,2]}\;\delta_{3;m}(Q_{2,1})-\frac{[0,2,m-1]}{[0,1,2]}\;\delta_{3;m}(1)=Q_{3,1}^{\frac{q^{m-2}-1}{q-1}}+\text{other terms},$
+    where the other terms are in the ideal $(Q_{3,0})$ of ${\mathcal{D}}_{3}$ generated by $Q_{3,0}$.
+  2. (ii)
+    For $20\leq\ell\leq m-$ 2, the following equality holds in $S$:
+    $\frac{[0,1,\ell+1]}{[0,1,2]}\;\delta_{3;m}(Q_{2,1})-\frac{[0,2,\ell+1]}{[0,1,2]}\;\delta_{3;m}(1)=Q_{3,2}^{\frac{q^{m-2}-q^{\ell}}{q-1}}Q_{3,1}^{\frac{q^{\ell}-1}{q-1}}+\text{other terms},$
+    where the other terms are in the ideal $(Q_{3,1}^{\frac{q^{\ell+1}-1}{q-1}},Q_{3,0})$ of ${\mathcal{D}}_{3}$.
 Proof.
-If $\lambda=0$, then $\delta_{3}(1)$ is a genuine Dickson polynomial in degree $q^{m}-q^{2}$.
-In this degree, there is only one Dickson monomial not in $\Delta^{m}_{3}$, namely $Q_{3,2}^{\frac{q^{m-2}-1}{q-1}}$.
-If we set $x_{2}=x_{3}=0$, then both $Q_{3,1}$ and $Q_{3,0}$ becomes zero, $\delta_{3}(1)$ becomes $x_{1}^{q^{m}-q^{2}}$.
-Thus in the expansion of $\delta_{3}(1)$ into a sum of Dickson monomials, $Q_{3,2}^{\frac{q^{m-2}-1}{q-1}}$ must occur with non-zero coefficient.
-The lemma is proved for $\lambda=0$.
-We now assume that $\lambda\geq 1$. Note that $\delta_{3}(Q_{2,1})=\frac{[0,2,m]}{[0,1,2]}$ is a genuine Dickson polynomial, hence in ${\mathcal{Q}}_{m}$,
-$\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})=Q_{3,2}^{\frac{q^{\lambda-1}-1}{q-1}}\delta_{3}(Q_{2,1})=Q_{3,2}^{\frac{q^{\lambda-1}-1}{q-1}}\frac{[0,2,m]}{[0,1,2]}.$
-So $\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ equals in ${\mathcal{Q}}_{m}$ to a Dickson polynomial and we are going to use this representation.
-Setting $x_{3}=0$, $\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ is reduced to
-$\frac{[\lambda+1,m]}{[1,2]}=y^{q^{\lambda+1}}_{m-\lambda-1,0}Q_{2,0}^{\frac{q^{\lambda+1}-q}{q-1}}=Q_{2,1}^{\frac{q^{m-1}-q^{\lambda+1}}{q-1}}Q_{2,0}^{\frac{q^{\lambda+1}-q}{q-1}}+\text{others}.$
-It follows that in this representation in ${\mathcal{Q}}_{m}$ of $\delta_{3;m}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ as a Dickson polynomial, $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,1}^{\frac{q^{\lambda-1}-1}{q-1}}$ must occur with nonzero coefficient.
-Moreover, any other monomials which involve $Q_{3,1}$ and $Q_{3,2}$ only must be divisible by $Q_{3,1}^{\frac{q^{\lambda}-1}{q-1}}$, and by degree reason, these monomials must be essential.
-∎
+The key observation is that if two Dickson polynomials coincide after setting one of the variables, say $x_{3}$, to zero, then their difference must be a multiple of $Q_{3,0}$.
+Note that in Dickson’s notation, we can write
+$\delta_{3;m}(Q_{2,1})=\frac{[0,2,m]}{[0,1,2]},\quad\delta_{3;m}(1)=\frac{[0,1,m]}{[0,1,2]},$
+so that all the terms in the above equations are genuine Dickson polynomials.
+Proof of (i): When $x_{3}=0$, the equation under consideration becomes
+$\frac{[1,m-1]}{[1,2]}\cdot\frac{[2,m]}{[1,2]}-\frac{[2,m-1]}{[1,2]}\cdot\frac{[1,m]}{[1,2]}=Q_{2,0}^{\frac{q^{m-1}-q}{q-1}}.$
+It is now a simple computation that the two sides are the same.
+Proof of (ii): Again, we consider the reduction of the left-hand side when setting $x_{3}=0$. The result, after simplification, is
+$\delta_{2;m-\ell-1}(1)^{q^{\ell+1}}Q_{2,0}^{\frac{q^{\ell+1}-q}{q-1}}.$
+Now setting $x_{2}=0$, $\delta_{2;m-\ell-1}(1)$ is reduced to $Q_{1,0}^{\frac{q^{m-\ell-1}-q}{q-1}}$, which implies that
+$\delta_{2;m-\ell-1}(1)=Q_{2,1}^{\frac{q^{m-\ell-2}-1}{q-1}}+\text{other terms divisible by $Q_{2,0}$}.$
+It follows that
+$\delta_{2;m-\ell-1}(1)^{q^{\ell+1}}Q_{2,0}^{\frac{q^{\ell+1}-q}{q-1}}=Q_{2,1}^{\frac{q^{m-1}-q^{\ell+1}}{q-1}}Q_{2,0}^{\frac{q^{\ell+1}-q}{q-1}}+\text{otherterms divisible by $Q_{2,0}^{\frac{q^{\ell+2}-q}{q-1}}$.}$
+Thus, we can conclude that the left-hand side equals
+$Q_{3,2}^{\frac{q^{m-2}-q^{\ell}}{q-1}}Q_{3,1}^{\frac{q^{\ell}-1}{q-1}}+\text{other terms},$
+where the other terms consist of Dickson monomials which are divisible by $Q_{3,1}^{\frac{q^{\ell+1}-1}{q-1}}$ or by $Q_{3,0}$. ∎
 The following corollary will be useful for induction.
 Corollary 7.7.
-Suppose $m-3\leq\lambda_{2}\leq\lambda_{3}\leq 0$ and $\lambda\leq m-2$.
+Suppose $20\leq\ell\leq m-$ 2 and $m-3\geq\lambda_{2}\geq\lambda_{3}\geq 0$. The following equalities hold in $S$:
   1. (1)
-    $\delta_{3}(\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1})=Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}+\text{other terms},$
-    where the other terms are either essential or non-essential and divisible by $Q_{3,0}^{\frac{q^{\nu}-1}{q-1}}$, where $\nu>\lambda_{3}$.
+    $\delta_{3;m-\ell-1}(1)^{q^{\ell}}\delta_{3;m}(Q_{2,1}^{q^{\ell}}Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})-\delta_{3;m-\ell-1}(Q_{2,1})^{q^{\ell}}\delta_{3;m}(Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})=Q_{3,1}^{\frac{q^{m-2}-q^{\ell}}{q-1}}Q_{3,0}^{\frac{q^{\ell}-1}{q-1}}+\text{other terms},$
+    where the other terms are in the ideal $(Q_{3,0}^{\frac{q^{\ell+1}-1}{q-1}})$.
   2. (2)
-    $\delta_{3}(\frac{q^{m-2}-q^{\lambda}}{q-1},\frac{q^{\lambda}-1}{q-1})=Q_{3,1}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,0}^{\frac{q^{\lambda}-1}{q-1}}+\text{other terms},$
-    where the other terms are either essential or non-essential and divisible by $Q_{3,0}^{\frac{q^{\nu}-1}{q-1}}$ where $\nu>\lambda$.
+    More generally,
+    $\delta_{3;\lambda_{2}-\lambda_{3}+1}(1)^{q^{\lambda_{3}}}\delta_{3;m}(Q_{2,1}^{q^{\lambda_{3}}}Q_{2,0}^{\frac{q^{\lambda_{3}}-1}{q-1}})-\delta_{3;\lambda_{2}-\lambda_{3}+1}(Q_{2,1})^{q^{\lambda_{3}}}\delta_{3;m}(Q_{2,0}^{\frac{q^{\lambda_{3}}-1}{q-1}})=\\ Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}+\text{other terms},$
+    where the other terms are in the ideal $(Q_{3,0}^{\frac{q^{\lambda_{3}+1}-1}{q-1}},Q_{3,1}^{\frac{q^{\lambda_{2}+1}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}-1}}{q-1}})$.
 Proof.
-This is just combining Lemma 7.4 and Proposition 7.5. ∎
+This is just combining Lemma 7.5 and Proposition 7.6.
+For the first part, note that
+$\delta_{3;m}(Q_{2,1}^{q^{\ell}}Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})=\delta_{3;m}(Q_{2,1}^{\frac{q^{\ell+1}-q^{\ell}}{q-1}}Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})=\delta_{3;m-\ell}(Q_{2,1})^{q^{\ell}}Q_{3,0}^{\frac{q^{\ell}-1}{q-1}},$
+and similarly
+$\delta_{3;m}(Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})=\delta_{3;m-\ell}(1)^{q^{\ell}}Q_{3,0}^{\frac{q^{\ell}-1}{q-1}}.$
+The left-hand side of the required formula becomes
+$\bigg{(}\delta_{3;m-\ell-1}(1)\delta_{3;m-\ell}(Q_{2,1})-\delta_{3;m-\ell-1}(Q_{2,1})\delta_{3;m-\ell}(1)\bigg{)}^{q^{\ell}}Q_{3,0}^{\frac{q^{\ell}-1}{q-1}}.$
+We can then apply 7.6 part (i) to finish the proof.
+The second part is proved similarly. ∎
 Corollary 7.8.
@@
 Proof.
-We have already seen from part (1) of Lemma 7.4 that $Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}$ belongs to $\Delta^{m}_{3}$. For the other 2 families, we will prove by downward induction on $\lambda_{3}$.
-If $\lambda_{3}=m-3$, then $\lambda_{2}=\lambda_{3}=m-3$.
-The edge monomials $Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$ and $Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$ belong to the span of $\mathcal{B}_{m}(3)$ by the last 2 parts of Lemma 7.4.
-We now suppose that the results have been verified for all edge elements corresponding to $\lambda_{3}+1$ and above.
-Consider the edge elements corresponding to $\lambda_{3}$:
+We have already seen from part (1) of Lemma 7.5 that the monomial $Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}$ belongs to $\delta_{3}(\Delta^{m}_{2})$.
+For the remaining $2$ families, we will show that they belong to the span of $\mathcal{B}_{m}(3)$ by downward induction on $\lambda_{3}$.
+When $\lambda_{3}=m-3$, it follows that $\lambda_{2}=\lambda_{3}=m-3$. The edge monomials
+$Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}\quad\text{and}\quad Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$
+belong to the span of $\mathcal{B}_{m}(3)$ by parts (1) and (3) of Lemma 7.5.
+In addition, by Proposition 2.7, any Dickson monomial divisible by these monomials is also in the image of $\delta_{3}$, and thus belongs to the span of $\mathcal{B}_{m}(3)$.
+Now suppose, as induction hypothesis, that all edge monomials with $Q_{3,0}$-exponent $\frac{q^{\lambda_{3}+1}-1}{q-1}$ as well as their multiples are already in the span of $\mathcal{B}_{m}(3)$.
+We consider the edge monomials corresponding to $\lambda_{3}$:
   1. (1)
@@
     $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$.
-Corollary 7.6 shows that these are in the image of $\delta_{3}(\Delta^{m}_{2})$ modulo those non-essential elements with strictly higher $\lambda_{3}$.
-By induction, these elements are already in the span of $\mathcal{B}_{m}(3)$.
+Part (i) of Corollary 7.7 shows that $Q_{3,1}^{\frac{q^{m-2}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$ lies in the image of $\delta_{3}(\Delta^{m}_{2})$ modulo terms with strictly higher $\lambda_{3}$.
+By the induction hypothesis, these higher terms are already in the span of $\mathcal{B}_{m}(3)$.
+For the second monomial, observe that if there exists a monomial with the same $Q_{3,0}$-exponent $\frac{q^{\lambda_{3}}-1}{q-1}$ but a higher exponent of $Q_{3,1}$, such that it becomes non-essential, then this monomial must be a multiple of
+$Q_{3,1}^{\frac{q^{m-2}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}},$
+which by the previous step, is already in the span of $\mathcal{B}_{m}(3)$.
+Alternatively, if the $Q_{3,0}$-exponent is strictly greater than $\frac{q^{\lambda_{3}}-1}{q-1}$, the induction hypothesis can be applied directly.
 ∎
-Now we are ready to finish the proof.
-Corollary 7.7 shows that the subspace of ${\mathcal{Q}}_{m}^{G}(3)$ spanned by $\mathcal{B}_{m}(3)$ is closed under the action of the Dickson algebra ${\mathcal{D}}_{3}$.
-Indeed, we have already computed the action of ${\mathcal{D}}_{3}$ on $\delta_{s+1}^{3-s}(\Delta^{m}_{s})$ for $0\leq s\leq 2$ in 2.5, and the product of $Q_{3,i}$ with a monomial in $\Delta^{m}_{3}$ is clearly an edge monomial, which is in the span of $\mathcal{B}_{m}(3)$ by the above demonstration.
+Now we are ready to finish the proof of the claim that $\mathcal{B}_{m}(3)$ is a generating set for ${\mathcal{Q}}_{m}(3)^{G}$.
+Corollary 7.8 shows that the subspace of ${\mathcal{Q}}_{m}(3)^{G}$ spanned by $\mathcal{B}_{m}(3)$ is closed under the action of the Dickson algebra ${\mathcal{D}}_{3}$.
+Indeed, we have already computed the action of ${\mathcal{D}}_{3}$ on $\delta_{s+1}^{3-s}(\Delta^{m}_{s})$ for $0\leq s\leq 2$ in 2.7, and the product of $Q_{3,i}$ with a monomial in $\Delta^{m}_{3}$ is clearly an edge monomial, which is in the span of $\mathcal{B}_{m}(3)$ by Corollary 7.8.
 The proof is finished.
-8. A filtration of submodules
-For each positive integer $n$, $0\leq k\leq n$, let $\mathcal{F}_{n,k}$ denote the $\mathbb{F}_{q}$-subspace of ${\mathcal{Q}}^{G}_{m}(n)$:
-$\mathcal{F}_{n,k}=\operatorname{Span}(\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq\min(m,k)).$
-For $n=3$, we have shown that this increasing filtration of subspaces is exhaustive.
-The calculation in previous sections shows that our basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{G}(3)$ is natural in the sense that the resulting filtration is a filtration of ${\mathcal{D}}_{3}$-submodules.
-Moreover, the action of the Dickson invariants is very similar to the pattern discussed in [10] about the structure of the cofixed space $S_{G}$ as a $S^{G}$-module.
-Indeed, we can easily check that $Q_{3,0}$ annihilates $\mathcal{F}_{3,2}$; $Q_{3,0}$ and $Q_{3,1}$ annihilate $\mathcal{F}_{3,1}$; and $Q_{3,0},Q_{3,1},Q_{3,2}$ annihilate $\mathcal{F}_{3,0}$.
-A similar pattern also occurs in the case $n=2$.
+8. A filtration of submodules over the Steenrod algebra
+In this final section, we investigate the filtration $\mathcal{F}_{n,k}$ of ${\mathcal{Q}}_{m}(n)^{G}$ as defined in the Introduction 1.10.
+Recall that for each positive integer $n$, $0\leq k\leq\min(m,n)$, $\mathcal{F}_{n,k}$ is the $\mathbb{F}_{q}$-subspace of ${\mathcal{Q}}_{m}(n)^{G}$:
+$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq\min(m,k)\}.$
 The aim of this final section is to demonstrate Theorem 1.12.
-We will show the filtration $\mathcal{F}_{n,*}$ also behaves well with respect to the action of the mod $q$ Steenrod algebra $\mathcal{A}$, at least in the case $n\leq 3$.
-For background on the mod $q$ Steenrod algebra and its role in modular invariant theory, we refer to Larry Smith’s article [16], which treats the Steenrod algebra completely algebraically over a general finite field.
-Recall that for each $i\geq 0$, the $i$th Steenrod reduced power operation $\mathcal{P}^{i}$ is a $\mathbb{F}_{q}$-linear map satisfying the following conditions:
+For $n\leq 3$, we have shown that this increasing filtration of subspaces is exhaustive.
+We begin with the following observation which allows more flexibility in working with $\mathcal{F}_{n,k}$:
+Lemma 8.1.
+Let $n\leq 3$. For each $0\leq k\leq\min(m,n)$, we have
+$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in{\mathcal{D}}_{s},0\leq s\leq\min(m,k)\}.$
+The point is that this filtration does not depend on the subspace $\Delta^{m}_{s}$.
+Proof.
+For $n=2$ this is essentially pointed out in Example 2.4 and the proof of Proposition 1.6. For $n=3$, this is a combination of Example 2.5 and the proof of Proposition 1.9. ∎
+Corollary 8.2.
+For $n\leq 3$, $\mathcal{F}_{n,*}$ is a filtration of $\mathcal{D}_{n}$-submodules of the invariant ring ${\mathcal{Q}}_{m}(n)^{G}$.
+Moreover, for $0\leq k<\min(m,n)$, $\mathcal{F}_{n,k}$ is annihilated by $Q_{n,0},\ldots,Q_{n,n-k-1}$.
+Proof.
+This is immediate from the above Lemma and Proposition 2.7. ∎
+The calculations in the previous sections demonstrate that our basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}(3)^{G}$ is natural in the sense that the resulting filtration is a filtration by ${\mathcal{D}}_{3}$-submodules.
+Moreover, the action of the Dickson algebra follows a pattern similar to the one discussed in [9], where the structure of the cofixed space $S_{G}$ is described as an $S^{G}$-module.
+A similar pattern also emerges in the case $n=2$.
+We now turn our attention to the Steenrod algebra.
+For background on the mod $q$ Steenrod algebra $\mathcal{A}$ and its role in modular invariant theory, we refer to Larry Smith’s article [13], which provides a purely algebraic treatment of the Steenrod algebra over a general finite field.
+Recall that for each $i\geq 0$, the $i$th Steenrod reduced power operation $\mathcal{P}^{i}$ is an $\mathbb{F}_{q}$-linear map satisfying the following conditions:
   1. (1)
@@
 Moreover, $\mathcal{P}^{0}$ acts as the identity.
-Over the polynomial ring $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, in which $x_{i}$ are given degree $1$, the action of the Steenrod algebra is completely determined by the unstable condition and the Cartan formulae.
+Over the polynomial ring $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, in which each $x_{i}$ is given degree $1$, the action of the Steenrod algebra is completely determined by the unstable condition and the Cartan formulae.
 More concretely, we have
@@
 Thus, if $j$ is a $q$-power, then $\mathcal{P}^{i}(v^{j})\neq 0$ iff $i=0$ (in which case $\mathcal{P}^{0}$ is the identity operator) or $i=j$ (in which case $\mathcal{P}^{i}(v^{i})=v^{qi}$ is the Frobenius operator).
-The Steenrod algebra action and the $G$ action on $S$ commutes.
-Furthermore, it is clear that $I_{m}$ is stable under the action of the Steenrod operations, so is ${\mathcal{Q}}_{m}=S/I_{m}$.
-It follows that a Steenrod reduced power yields a map from ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ to itself.
-In other words, ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is an $\mathcal{A}$-module.
-The proof of Theorem 1.11 uses several lemmas. We begin with a general formula describing how the Steenrod operation interacts with the delta operator.
+It is convenient to make the convention that $\mathcal{P}^{k}=0$ if $k<0$.
+The Steenrod algebra action and the group $G$ action on $S$ commute.
+Moreover, it is evident that the Frobenius ideal $I_{m}$ is stable under the action of the Steenrod operations.
+Consequently, the quotient ${\mathcal{Q}}_{m}=S/I_{m}$ inherits the structure of an $\mathcal{A}$-module.
+It follows that each Steenrod reduced power $\mathcal{P}^{i}$ induces a well-defined endomorphism of ${\mathcal{Q}}_{m}^{G}$.
+In other words, the subspace of invariants ${\mathcal{Q}}_{m}^{G}$ is closed under the action of the Steenrod algebra, and therefore forms an $\mathcal{A}$-module.
+The proof of Theorem 1.12 uses several lemmas. We begin with a general formula describing how the Steenrod operations interact with the delta operator.
 Lemma 8.3.
-If $f$ is a polynomial which is $\operatorname{GL}_{1}$-invariant in the first variable and $1\leq k\leq\deg(\delta_{2}(f))=q^{m}-q+\deg(f)$, then
+If $f$ is a polynomial which is $G_{1}$-invariant in the first variable, then for any $k\geq 0$, the following holds in $S$:
 $\mathcal{P}^{k}(\delta_{2}f)+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}f)+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}f)=\\ \delta_{2}(Q_{1,0}\mathcal{P}^{k-1}f)+\delta_{2}(\mathcal{P}^{k}f)+\delta_{2;m+1}(Q_{1,0}\mathcal{P}^{k-1-q^{m}}f)+\delta_{2;m+1}(\mathcal{P}^{k-q^{m}}f).$
-Moreover, the last two summands on the right-hand side vanish in ${\mathcal{Q}}_{m}$.
 Proof.
@@
 The equation above is then obtained by applying the Cartan formula for the product $L_{2}\delta_{2}(f)$ and then dividing both sides by $L_{2}$.
-Note that $\mathcal{P}^{q}L_{2}=Q_{2,1}L_{2}$, $\mathcal{P}^{q+1}L_{2}=Q_{2,0}L_{2}$ and $\mathcal{P}^{k}L_{2}=0$ for all $k\geq 1$.
+Note that $\mathcal{P}^{q}L_{2}=Q_{2,1}L_{2}$, $\mathcal{P}^{q+1}L_{2}=Q_{2,0}L_{2}$ and $\mathcal{P}^{k}L_{2}=0$ for all other $k\geq 1$.
 ∎
 Corollary 8.4.
-For each $0\leq k\leq 2$, $\mathcal{F}_{2,k}$ is a submodule of ${\mathcal{Q}}_{m}^{G}(2)$ over the Steenrod algebra.
+$\mathcal{F}_{2,1}$ is a submodule of ${\mathcal{Q}}_{m}(2)^{G}$ over the Steenrod algebra.
 Proof.
-We need to compute the action of the Steenrod operation on a typical element $\delta_{2}(f)$ where $f=Q_{1,0}^{a}$, $a<[m]_{q}$.
-In the rank $2$ case, the last two summands on the right-hand side of the formula given in Lemma 8.1 vanish in ${\mathcal{Q}}_{m}$.
-Indeed, one check by direct inspection that the degree of $\delta_{2;m+1}(Q_{1,0}\mathcal{P}^{k-1}f)$ is strictly greater than $2(q^{m}-1)$ for all $q$ and $m$.
-For $\delta_{2;m+1}(\mathcal{P}^{k-q^{m}}f)$, its degree is not greater than $2(q^{m}-1)$ only when $q=2$, $k\geq q^{m}$ and $\deg f=0$, but in this case $k>q^{m}-q+\deg(f)$.
-Now we can proceed by induction on $k$ using the above formula and the calculation in Proposition 2.5. ∎
-Corollary 8.3.
-$\mathcal{F}_{3,1}$ is a submodule of ${\mathcal{Q}}_{m}^{G}(3)$ over the Steenrod algebra.
+We must show that for each $0\leq s<[m]_{q}$, $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{s})\in\mathcal{F}_{2,1}$ for all $k>0$.
+If $k\geq q^{m}$, the degree of $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{s})$ is at least
+$q^{m}(q-1)+q^{m}-q+s(q-1)\geq 2(q^{m}-1).$
+So we can assume that $k<q^{m}$. In this case, when we apply the formula of Lemma 8.3 for $f=Q_{1,0}^{a}$, the last two summands on the right-hand side vanish. Thus we have an equality in $S$:
+(8.1) $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{s})+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}Q_{1,0}^{s})+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}Q_{1,0}^{s})=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}Q_{1,0}^{s})+\delta_{2}(\mathcal{P}^{k}Q_{1,0}^{s}).$
+Both summands on the right-hand side are evidently of the form $\delta_{2}(Q_{1,0}^{s})$.
+Now we can proceed by induction on $k$ using the above formula and the calculation in Proposition 2.7 when working in ${\mathcal{Q}}_{m}(2)$.
+∎
+Note that $\mathcal{F}_{2,0}$ is spanned by the top class $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$, and $\mathcal{F}_{2,2}={\mathcal{Q}}_{m}(2)^{G}$.
+Both are evidently $\mathcal{A}$-submodules of ${\mathcal{Q}}_{m}(2)^{G}$.
+Proposition 8.5.
+$\mathcal{F}_{3,1}$ is a submodule of ${\mathcal{Q}}_{m}(3)^{G}$ over the Steenrod algebra.
 Proof.
-We need to show that for each $k\geq 1$ and $a<[m]_{q}$, $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ is in the span of $\delta_{1}^{3}(1)$ and $\delta_{2}^{2}(\Delta^{m}_{1})$.
-Apply the formula provided in Lemma 8.1 for $f=\delta_{2}Q_{1,0}^{a}$.
-Again, the last two summands are trivial in ${\mathcal{Q}}_{m}$ because of the instability condition and degree calculation.
-We then obtain a recursive equation in ${\mathcal{Q}}_{m}$:
-$\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}Q_{1,0}^{a})+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}Q_{1,0}^{a})=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}\delta_{2}Q_{1,0}^{a})+\delta_{2}(\mathcal{P}^{k}\delta_{2}Q_{1,0}^{a}).$
-We now proceed by double induction on $k$ that $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ belongs to $\mathcal{F}_{3,1}$.
-We already know from Corolarry 8.2 that $\mathcal{P}^{k-1}(\delta_{2}Q_{1,0}^{a})$ and $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{a})$ belongs to the span of $\delta_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}(1)$.
-By Lemma 7.2, the two summands on the right-hand side in the equation above belong to the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$.
-Also by induction, we see that $Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}Q_{1,0}^{a})=0$ in ${\mathcal{Q}}_{m}$ and $Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}Q_{1,0}^{a})$ is also in the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$.
-We conclude that $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ is also in the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$, but since it is a $\operatorname{GL}_{3}$-invariant, it is spanned by $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{3}(1)$.
-∎
+We need to show that for each $k\geq 1$ and $0\leq a<[m]_{q}$, $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{s})$ is in the span of $\delta_{1}^{3}(1)$ and $\delta_{2}^{2}(\Delta^{m}_{1})$.
+Note that $\delta_{1}^{3}(1)$ spans the one-dimensional space of ${\mathcal{Q}}_{m}(3)^{G}$ in the top degree $3(q^{m}-1)$.
+First of all, if $k\geq q^{m}$, then the degree of $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{s})$, which is $q^{m}(q-1)+2(q^{m}-q)+s(q-1)$ is at least $3(q^{m}-1)$ unless $q=2$ and $k=2^{m}$, $s=0$.
+In this range, ${\mathcal{Q}}_{m}(3)$ is either spanned by $\delta_{3}^{3}(1)$ or zero.
+If $q=2$, $k=2^{m}$, $s=0$, then $\mathcal{P}^{2^{m}}(\delta_{2}^{2}(1))$ is in degree $13(2^{m}-1)-$ 1.
+We claim that there is no $G$-invariant polynomial in this degree.
+Indeed, the only non-zero symmetric polynomial in this degree is
+$x_{1}^{2^{m}-1}x_{2}^{2^{m}-1}x_{3}^{2^{m}-2}+x_{1}^{2^{m}-1}x_{2}^{2^{m}-2}x_{3}^{2^{m}-1}+x_{1}^{2^{m}-2}x_{2}^{2^{m}-1}x_{3}^{2^{m}-2}.$
+By direct inspection, this polynomial is not invariant under the action of the operation mapping $x_{1}\mapsto x_{1}+x_{2}$ and fixes $x_{2},x_{3}$.
+Now let us consider the case $k<q^{m}$. By Lemma 8.3, we have the following equality in $S$:
+(8.2) $\mathcal{P}^{k}(\delta_{2}f)+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}f)+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}f)=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}f)+\delta_{2}(\mathcal{P}^{k}f)$
+for all polynomial $f$ which is $GL_{1}$-invariant in the first variable.
+Applying this for $f=Q_{1,0}^{s}$, it follows by induction on $k$ that for $0\leq k<q^{m}$, $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{s})$ is a linear combination of polynomials of the form
+$Q_{2,1}^{a}Q_{2,0}^{b}\delta_{2}(Q_{1,0}^{c}).$
+Now applying (8.2) again but for $f=\delta_{2}(Q_{1,0}^{s})$, we get
+(8.3) $\mathcal{P}^{k}(\delta_{2}^{2}(Q_{1,0}^{s}))+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}(Q_{1,0}^{s}))+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}(Q_{1,0}^{s}))=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}\delta_{2}(Q_{1,0}^{s}))+\delta_{2}(\mathcal{P}^{k}\delta_{2}(Q_{1,0}^{s})).$
+The right-hand side is then a sum of polynomials of the following forms
+  • $\delta_{2}(Q_{1,0}Q_{2,1}^{a}Q_{2,0}^{b}\delta_{2}(Q_{1,0}^{c}))$, or
+  • $\delta_{2}(Q_{2,1}^{a}Q_{2,0}^{b}\delta_{2}(Q_{1,0}^{c}))$.
+Up to this point, we have been working in $S$.
+Now we consider the situation in ${\mathcal{Q}}_{m}(3)$.
+From Lemma 7.2, we see that this sum belongs to the subspace of ${\mathcal{Q}}_{m}(3)$ spanned by polynomials of the forms
+  • $\delta_{2}^{2}(Q_{1,0}^{a})$, or
+  • $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{a(q-1)}$.
+Both sides of Equation 8.3 now are $P(2,1)$-invariant. Take the transfer from $P(2,1)$ to $G$. We have $\mathcal{P}^{i}(\delta_{2}^{2}(Q_{1,0}^{s}))$ is $G$-invariant in ${\mathcal{Q}}_{m}(3)$ and
+$\operatorname{Tr}_{P(2,1)}^{G}(Q_{2,1})=\operatorname{Tr}_{P(2,1)}^{G}(Q_{2,0})=\operatorname{Tr}_{P(2,1)}^{G}(x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{a(q-1))})=0.$
+The first two terms vanish in $S$ for degree reasons: there are no genuine $G$-invariants in degrees $q^{2}-q$ and $q^{2}-1$.
+The third vanishes in ${\mathcal{Q}}_{m}(3)$ because if $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{a(q-1)}$ is a nontrivial summand of a $G$-invariant polynomial in ${\mathcal{Q}}_{m}(3)$, then it is a $P(1,2)$-invariant as well, but from Remark 6.1, this can only happen at the top degree.
+It follows that $\mathcal{P}^{k}(\delta_{2}^{2}(Q_{1,0}^{s}))$ must be a scalar multiple of some $\delta_{2}^{2}(Q_{1,0}^{a})$, and thus belongs to $\mathcal{F}_{3,1}$. ∎
 We now consider the interaction between the Steenrod powers and the operator $\delta_{3}$.
 Lemma 8.6.
-If $f\in{\mathcal{D}}_{2}$ and $k\geq 1$, then
-$\mathcal{P}^{k}(\delta_{3}f)+\sum_{i<k}\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)=\frac{\mathcal{P}^{k}\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{q}&x_{2}^{q}&x_{3}^{q}\\ x_{1}^{q^{m}}f(x_{2},x_{3})&x_{2}^{q^{m}}f(x_{1},x_{3})&x_{3}^{q^{m}}f(x_{1},x_{2})\end{vmatrix}}{L_{3}}.$
-Moreover, the right-hand side equals $\delta_{3;m}(F)$ or $\delta_{3;m+1}(F)$ where $F$ is a certain linear combination of polynomials of the form $g\mathcal{P}^{i}f$ where $g\in{\mathcal{D}}_{2}$ and $0\leq i\leq k$.
+If $f$ is a polynomial which is $G_{2}$-invariant in the first two variables, then for any $k\geq 0$, the following holds in $S$:
+$\mathcal{P}^{k}(\delta_{3}f)+Q_{3,2}\mathcal{P}^{k-q^{2}}(\delta_{3}f)+Q_{3,1}\mathcal{P}^{k-q^{2}-q}(\delta_{3}f)+Q_{3,0}\mathcal{P}^{k-q^{2}-q-1}(\delta_{3}f)=\\ \delta_{3}\bigg{(}\mathcal{P}^{k}(f)+Q_{2,1}\mathcal{P}^{k-q}(f)+Q_{2,0}\mathcal{P}^{k-q-1}(f)\bigg{)}+\delta_{3;m+1}\bigg{(}\mathcal{P}^{k-q^{m}}(f)+Q_{2,1}\mathcal{P}^{k-q^{m}-q}(f)+Q_{2,0}\mathcal{P}^{k-q^{m}-q-1}(f)\bigg{)}.$
 Proof.
-The above equality is a simple application of the Cartan formula for $\mathcal{P}^{k}(L_{2}\delta_{2;m}(f))$ (or $\mathcal{P}^{k}(L_{3}\delta_{3}f)$).
-For the right-hand side, note that because of the instability condition, we have $\mathcal{P}^{i}(x_{1})$ is either $x_{1}$ or $x_{1}^{q}$, depending on whether $i=0$ or $i=1$, and there is a similar formula for other entries in the determinant.
-In the first case, $\delta_{2;m+1}(g\mathcal{P}^{i}f)$ is trivial in ${\mathcal{Q}}_{m}$ unless $q=2$ and $g\mathcal{P}^{i}f=1$.
-In this case, $\delta_{2;m+1}(1)=x_{1}^{2^{m}-1}x_{2}^{2^{m}-1}$ because of degree reason.
-We also have $g=f=1$, and $k=2^{m}$, but then $\delta_{2}f$ has degree $2^{m}-2<2^{m}$.
-Finally, observe that any determinant of the form
-$\begin{vmatrix}x_{1}^{q^{a}}&x_{2}^{q^{a}}&x_{3}^{q^{a}}\\ x_{1}^{q^{b}}&x_{2}^{q^{b}}&x_{3}^{q^{b}}\\ x_{1}^{q^{c}}\mathcal{P}^{d}f(x_{2},x_{3})&x_{2}^{q^{c}}\mathcal{P}^{d}f(x_{1},x_{3})&x_{3}^{q^{c}}\mathcal{P}^{d}f(x_{1},x_{2})\end{vmatrix}$
-can be written as
-$L_{3}\delta_{3;c}\bigg{(}\frac{[a,b]}{[0,1]}\mathcal{P}^{d}f\bigg{)}.$
-where $\frac{[a,b]}{[0,1]}\in\mathcal{D}_{2}$. Note that $c$ is either $m$ or $m+1$. ∎
-It is also clear that $\mathcal{P}^{i}L_{3}$ is again a matrix whose entries are $q$-powers, so it is always a multiple of $L_{3}$ and the quotient $\frac{\mathcal{P}^{i}L_{3}}{L_{3}}$ is a Dickson polynomial in ${\mathcal{D}}_{3}$.
-The next lemma discusses how to deal with the case when $c=m+1$.
+The above equality is again an application of the Cartan formula to $\mathcal{P}^{k}(L_{3}\delta_{3}f)$, using the formulae for the action of the Steenrod powers on $L_{2}$ given in the proof of Lemma 8.3, and similar formulae for these actions on $L_{3}$: $\mathcal{P}^{q^{2}}L_{3}=Q_{3,2}L_{3}$, $\mathcal{P}^{q^{2}+q}L_{3}=Q_{3,1}L_{3}$, $\mathcal{P}^{q^{2}+q+1}L_{3}=Q_{3,0}L_{3}$ and $\mathcal{P}^{k}L_{3}=0$ for all other $k\geq 1$.
+∎
+The next lemma discusses how to deal with the last term in the above formula.
 Lemma 8.7.
@@
 Proof.
-The degree of $\delta_{3;m+1}(h)$ equals $q^{m+1}-q^{2}+\deg(h)$. If $q>3$, then this is strictly greater than $3(q^{m}-1)$, and we immediately have $\delta_{3;m+1}(h)\equiv 0$ in this case.
-From the fundamental equation, we know that $X^{q^{3}}-Q_{3,2}X^{q^{2}}+Q_{3,1}X^{q}-Q_{3,0}X=0$ if $X$ is in the span of $x_{1},x_{2},x_{3}$. It follows that
-$x_{i}^{q^{m+1}}-Q_{3,2}^{q^{m-2}}x_{i}^{q^{m}}+Q_{3,1}^{q^{m-2}}x_{i}^{q^{m-1}}-Q_{3,0}^{q^{m-2}}x_{i}^{q^{m-2}}=0,$
-and so
+Recall from the fundamental equation that $X^{q^{3}}-Q_{3,2}X^{q^{2}}+Q_{3,1}X^{q}-Q_{3,0}X=0$ if $X$ is in the span of $x_{1},x_{2},x_{3}$. It follows that
+$x_{i}^{q^{m+1}}-Q_{3,2}^{q^{m-2}}x_{i}^{q^{m}}+Q_{3,1}^{q^{m-2}}x_{i}^{q^{m-1}}-Q_{3,0}^{q^{m-2}}x_{i}^{q^{m-2}}=0,\quad 1\leq i\leq 3,$
+and we obtain the following identity in $S$ relating $\delta_{3}$ for various $m$:
 $\delta_{3;m+1}(h)-Q_{3,2}^{q^{m-2}}\delta_{3;m}(h)+Q_{3,1}^{q^{m-2}}\delta_{3;m-1}(h)-Q_{3,0}^{q^{m-2}}\delta_{3;m-2}(h)=0.$
-If $q=3$, then $q^{m+1}-q^{2}+\deg(h)\leq 3(q^{m}-1)$ implies $\deg(h)\leq q^{2}-3=6$. Since $h\in{\mathcal{D}}_{2}$, it follows that up to a scalar, $h=1$ or $h=Q_{2,1}$.
-In the case where $h=Q_{2,1}$, $\delta_{3;m+1}(Q_{2,1})$ has degree $33^{m+1}-$ 3 so that it is either zero or the top element $x_{1}^{3^{m}-1}x_{2}^{3^{m}-1}x_{3}^{3^{m}-1}$.
-We claim that $\delta_{3;m+1}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m}$.
-Indeed, when $q=3$, then by direct computation, we have $Q_{3,0}=Q_{3,1}=0$ in ${\mathcal{Q}}_{2}$ so that $Q_{3,1}^{q^{m-2}}=Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}$.
-It follows that in ${\mathcal{Q}}_{m}$, we have
-$\delta_{3;m+1}(Q_{2,1})=Q_{3,2}^{q^{m-2}}\delta_{3;m}(Q_{2,1}).$
-On the other hand, $Q_{3,2}=x_{1}^{6}x_{2}^{6}x_{3}^{6}$ in ${\mathcal{Q}}_{2}$ so that $Q_{3,2}^{q^{m-2}}=x_{1}^{2.3^{m-1}}x_{2}^{2.3^{m-1}}x_{3}^{2.3^{m-1}}$ in ${\mathcal{Q}}_{m}$.
-So if we can show that $\delta_{3;m}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m-1}$, then it follows immediately that $\delta_{3;m+1}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m}$.
-The claim can now be easily verified by checking that $\delta_{3;3}(Q_{2,1})=Q_{3,1}=0$ in ${\mathcal{Q}}_{2}$.
-Suppose now $h=1$. In this case, we have
-$\delta_{3;m+1}(1)=Q_{3,2}^{3^{m-2}}\delta_{3;m}(1)=\delta_{3;m}(Q_{2,1}^{3^{m-1}}),$
-as required. Here the second equality follows from 2.5.
-Now consider the case $q=2$. We still have $Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}(3)$, but now
-$Q_{3,1}^{2^{m-2}}=x_{1}^{2^{m-1}}x_{1}^{2^{m-1}}x_{1}^{2^{m-1}}.$
-By direct inspection, we have $\delta_{3;m+1}(h)=Q_{3,2}^{2^{m-2}}\delta_{3;m}(h)-Q_{3,1}^{2^{m-2}}\delta_{3;m-1}(h)$ equals in ${\mathcal{Q}}_{m}$ to
+Note that the degree of $\delta_{3;m+1}(h)$ equals $q^{m+1}-q^{2}+\deg(h)$.
+If $q>3$, this is strictly greater than $3(q^{m}-1)$, and we can conclude immediately that $\delta_{3;m+1}(h)=0$ in $\mathcal{Q}_{m}(3)$.
+Consider now the case $q=3$.
+Observe that, by direct computation, $Q_{3,0}=Q_{3,1}=0$, $Q_{3,2}=(x_{1}x_{2}x_{3})^{3^{2}-3^{1}}$ in ${\mathcal{Q}}_{2}(3)$.
+Hence $Q_{3,1}^{q^{m-2}}=Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}(3)$ and the above equation becomes
+(8.4) $\delta_{3;m+1}(h)=Q_{3,2}^{q^{m-2}}\delta_{3;m}(h)=(x_{1}x_{2}x_{3})^{3^{m}-3^{m-1}}\delta_{3;m}(h)\quad\text{in ${\mathcal{Q}}_{m}(3)$}.$
+On the other hand, for degree reasons, we must have $\deg(h)\leq 3(3^{m}-1)-(3^{m+1}-3^{2})=6$.
+In this range, there are (up to a scalar) only two Dickson polynomials, namely $h=1$ or $h=Q_{2,1}$.
+If $h=1$, then from Proposition 2.7, we have immediately
+$\delta_{3;m+1}(1)=Q_{3,2}^{3^{m-2}}\delta_{3;m}(1)=\delta_{3;m}(Q_{2,1}^{3^{m-1}}).$
+If $h=Q_{2,1}$, we prove by induction on $m\geq 2$ that
+$\delta_{3;m+1}(Q_{2,1})=0\quad\text{in ${\mathcal{Q}}_{m}(3)$}.$
+Indeed, the base case can be checked directly since $\delta_{3;3}(Q_{2,1})=Q_{3,1}$ in $S$, and the induction step follows immediately from Equation (8.4).
+The next and final case is $q=2$. Using the identity
+$x_{1}^{2^{m+1}}=x_{1}^{2^{m}}(x_{1}^{2^{m}}+x_{2}^{2^{m}}+x_{3}^{2^{m}})-x_{1}^{2^{m}}(x_{2}^{2^{m}}+x_{3}^{2^{m}}),$
+and similarly for $x_{2}^{2^{m+1}}$ and $x_{3}^{2^{m+1}}$, we see that $\delta_{3;m+1}(h)$ equals in ${\mathcal{Q}}_{m}(3)$ to
 $\frac{1}{L_{3}}\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{2}&x_{2}^{2}&x_{3}^{2}\\ x_{1}^{2^{m}}h(x_{2},x_{3})(x_{2}^{2^{m}}+x_{3}^{2^{m}})&x_{2}^{2^{m}}h(x_{1},x_{3})(x_{1}^{2^{m}}+x_{3}^{2^{m}})&x_{3}^{2^{m}}h(x_{1},x_{2})(x_{1}^{2^{m}}+x_{2}^{2^{m}})\end{vmatrix}.$
-It is straightforward to verify that
+It is straightforward to verify the following equality in $S$:
 $x_{1}^{2^{m}}+x_{2}^{2^{m}}=Q_{2,1}\delta_{2;m}(1)-\delta_{2;m}(Q_{1,0}^{2}),$
-so the expression above becomes
-$\delta_{3;m}(Q_{2,1}h\delta_{2;m}(1))-\delta_{3;m}(h\delta_{2;m}(Q_{1,0}^{2})).$
-Now we can use the argument as in the proof of Lemma 7.3 to conclude. ∎
+We then obtain an equality in ${\mathcal{Q}}_{m}(3)$:
+$\delta_{3;m+1}(h)=\delta_{3;m}(Q_{2,1}h\delta_{2;m}(1))-\delta_{3;m}(h\delta_{2;m}(Q_{1,0}^{2})).$
+The conclusion follows from Corollary 7.3. ∎
 Remark 8.8.
-With a little more work, using 7.2, it can be showed that when $q=2$, then for $h\in{\mathcal{D}}_{2}$, we have an explicit formula:
-$\delta_{3;m+1}(h)=\begin{cases}0&\text{if $h$ is a multiple of $Q_{2,0}$,}\\ \delta_{2;m}^{2}(Q_{1,0}^{2s})&\text{if $h=Q_{2,1}^{s}$}.\end{cases}$
+With a little more work, using 7.2, it can be shown that when $q=2$, then for $h\in{\mathcal{D}}_{2}$, we have the following explicit formula for $\delta_{3;m+1}(h)$ in ${\mathcal{Q}}_{m}(3)$:
+$\delta_{3;m+1}(h)=\begin{cases}0&\text{if $h$ is a multiple of $Q_{2,0}$;}\\ \delta_{2;m}^{2}(Q_{1,0}^{2s})&\text{if $h=Q_{2,1}^{s}$, $s\geq 0$}.\end{cases}$
 We leave the details for the interested reader.
 Now we can finish the proof of Theorem 1.12.
-It remains to show that $\mathcal{F}_{3,1}$ is an $\mathcal{A}$-submodule of ${\mathcal{Q}}_{m}^{G}(3)$.
-From Lemma 8.5 we see that if $f\in{\mathcal{D}}_{2}$, then
-$\mathcal{P}^{k}(\delta_{3}f)+\sum_{i<k}\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)$
-belongs to $\mathcal{F}_{3,1}$.
-By induction, so is $\mathcal{P}^{i}(\delta_{3}f)$ for each $i<k$ and from Proposition 2.5, $\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)$ also belongs to $\mathcal{F}_{3,1}$.
+It remains to show that $\mathcal{F}_{3,2}$ is an $\mathcal{A}$-submodule of ${\mathcal{Q}}_{m}(3)^{G}$.
+From Lemma 8.7, we see that if $f\in{\mathcal{D}}_{2}$, then
+$\mathcal{P}^{k}(\delta_{3}f)+Q_{3,2}\mathcal{P}^{k-q^{2}}(\delta_{3}f)+Q_{3,1}\mathcal{P}^{k-q^{2}-q}(\delta_{3}f)+Q_{3,0}\mathcal{P}^{k-q^{2}-q-1}(\delta_{3}f)\in\mathcal{F}_{3,2}.$
+We can proceed by induction.
+If $\mathcal{P}^{i}(\delta_{3}f)\in\mathcal{F}_{3,2}$ for all $i<k$, then $Q_{3,j}\mathcal{P}^{i}(\delta_{3}f)$ also belongs to $\mathcal{F}_{3,2}$ by Proposition 2.7.
+It follows that $\mathcal{P}^{k}(\delta_{3}f)$ also belongs to $\mathcal{F}_{3,2}$ and the proof is complete.